### v1

# ProFL: Performative Robust Optimal Federated Learning

Abstract
Performative prediction (PP) is a framework that captures distribution shifts that occur during the training of machine learning models due to their deployment. As the trained model is used, its generated data could cause the model to evolve, leading to deviations from the original data distribution. The impact of such model-induced distribution shifts in the federated learning (FL) setup remains unexplored, despite being increasingly likely to transpire in real-life use cases. Although Jin et al., (2024) recently extended PP to FL in a straightforward manner, the resulting model only converges to a performative stable point, which may be far from optimal. The methods in Izzo et al., (2021); Miller et al., (2021) can find a performative optimal point in centralized settings, but they require the performative risk to be convex and the training data to be noiseless—assumptions often violated in realistic FL systems. This paper overcomes all of these shortcomings and proposes Performative robust optimal Federated Learning (ProFL), an algorithm that finds performative optimal points in FL from noisy and contaminated data. We present the convergence analysis under the Polyak-Lojasiewicz condition, which applies to non-convex objectives. Extensive experiments on multiple datasets validate our proposed algorithms’ efficiency.
Keywords Federated learning  $\cdot$ Performative prediction $\cdot$ Robust
1 Introduction
Federated learning (FL) is a distributed learning paradigm that enables efficient learning from heterogeneous clients. Conventional FL (e.g., collaborative model training) often relies on the unrealistic assumption that each client’s data distribution remains static throughout the training process or between the training and testing phases. Recent work addresses this weakness by training models on time-evolving data (Guo et al.,, 2023; Ma et al.,, 2022) or training models on static data but deploying them on different data distributions (Nguyen et al.,, 2022; Jiang and Lin,, 2022). However, these works assume that the changes in data distribution are exogenous (i.e., are independent of the FL system). Examples of such exogenous distribution shifts include variations in data quality, lighting conditions, dynamic climate data, etc.
In many real-world applications, data often undergoes performative endogenous shifts induced by the model’s deployment. For instance, content recommended by machine learning algorithms on digital platforms can reshape consumer preferences, steering them toward modes of consumption that are easier to predict and monetize (Dean and Morgenstern,, 2022). Similarly, navigation systems can influence traffic patterns, while individuals may adapt their behavior to game decision-making systems, such as those used for loan approvals or job applications (Hardt et al.,, 2016; Zhang et al.,, 2022). These types of shifts have been studied in centralized settings under the framework of performative prediction (PP), which seeks performative stable (Perdomo et al.,, 2020) or optimal (Izzo et al.,, 2021; Miller et al.,, 2021) solutions. Addressing endogenous shifts in federated learning (FL) is not a simple extension of these centralized efforts and remains a significant challenge. Failing to do so can lead to suboptimal performance or even non-convergence.
Another challenge in practical FL systems, beyond model-induced endogenous shifts, is data contamination. Client data may be noisy, corrupted, or maliciously manipulated. Unlike centralized systems, in FL, such data cannot be removed before training, as it remains with the clients. Developing FL systems that are robust to endogenous shifts and data contamination remains a significant challenge.
The only work to date that addresses the endogenous shifts in FL is Jin et al., (2024). Their approach assumes noiseless client data and only finds a performative stable point, which may be suboptimal. While Izzo et al., (2021); Miller et al., (2021) achieve a performative optimal point, their methods are centralized, assume noiseless data, and require convex performative risk. As we will show, the method that directly extends Izzo et al., (2021) to federated settings is highly vulnerable to data contamination and performs poorly when distribution shifts are complex.
In this paper, we solve all these challenges simultaneously and propose Performative robust optimal Federated Learning (ProFL), an FL algorithm that finds a performative optimal solution under endogenous distribution shifts with contaminated data. We provide a theoretical analysis of data contamination and the convergence of ProFL. Compared to the federated extension of Izzo et al., (2021), ProFL employs simple yet effective strategies and converges to the performative optimal solution for a larger class of (non-convex) objective functions. In Appendix B, we provide more related work and highlight differences.
Summary of contributions. Our key contributions compared to prior works are:
  1. 1.
    Finding the optimal solution. Unlike Jin et al., (2024), which only guarantees convergence to the stable point, ProFL finds the optimal point.
  2. 2.
    Convergence analysis for a larger class of objectives. Unlike prior works (Izzo et al.,, 2021; Jin et al.,, 2024) that show the convergence of the proposed algorithms for the convex performative risk, we show the convergence of ProFL under Polyak-Lojasiewicz (PL) condition, which is non-convex.
  3. 3.
    Robustness to data contamination. We analyze the impact of contaminated data on the FL system and show that ProFL effectively tackles complex distribution shifts and is more resilient to estimation error and data contamination.
2 Problem Formulation
Consider a set of clients $\mathcal{V}=\{1,\ldots,N\}$ collaboratively training an FL model $\theta\in\Theta\subset\mathbb{R}^{d}$ without sharing their local data. Here, $\Theta$ is a compact set (closed and bounded), and $d$ represents the model’s dimension. Let $\mathcal{Z}$ denote the sample space. Throughout training [^1] One can also interpret training as retraining or fine-tuning of models after deployment., each client interacts with the local model, causing its data to evolve based on the perceived model. Such model-induced data distributions, studied in PP (Perdomo et al.,, 2020), are characterized by a performative distribution map $D_{i}:\Theta\to\mathcal{P}(\mathcal{Z})$. Specifically, deploying model $\theta$ at client $i$ results in data following distribution $D_{i}(\theta)$. The goal is to train a global model that minimizes the total loss across all clients:
${{\theta}}^{\text{PO}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}(\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$ (1)
where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\ell(Z_{i};\theta)]$ is the local performative risk of client $i$ for a given loss function $\ell(Z_{i};\theta)$ on local data $Z_{i}\sim D_{i}(\theta)$. This is called performative risk because, instead of measuring the risk over a fixed distribution, it measures the expected loss of model $\theta$ on the data distribution $D_{i}(\theta)$ it induces. $\mathcal{L}(\theta)$ is the performative risk (PR) of all clients. Here, $\alpha_{i}\geq 0$ represents the fraction of client $i$’s data relative to the total, with $\sum_{i\in\mathcal{V}}\alpha_{i}=1$. The solution ${\theta}^{\text{PO}}$ that minimizes performative risk is called performative optimal (PO) point (Perdomo et al.,, 2020).
In realistic FL systems, each client’s data distribution $D_{i}(\theta)$ is unknown, and the training data may be noisy, corrupted, or maliciously manipulated. To account for this, we assume each client $i$ has a fraction $\epsilon_{i}\in[0,1]$ of data samples drawn from another fixed but arbitrary distribution $Q_{i}$. Thus, the data acquired from client $i$ follows a mixture distribution ${Z}_{i}\sim P_{i}(\theta)=(1-\epsilon_{i})D_{i}(\theta)+\epsilon_{i}Q_{i}$. We term $Q_{i}$ data “contamination”.
Our goal is to (i) design FL algorithms that converge to the performative optimal ${\theta}^{\text{PO}}$; (ii) study the impact of data contamination $Q_{i}$ and develop robust solutions. While data contamination has been studied in FL with static data, its effect in performative settings is unexplored. A moment’s thought reveals that when client data depends on the FL system, contamination in one client may cause a cascading effect across the network, making it crucial to investigate how this disrupts FL convergence. We will explore this in the paper. We next present the metrics in PP.
Performative stable point. Since the data distribution $D_{i}(\theta)$ depends on $\theta$, which is the variable being optimized, finding the performative optimal point is often challenging. Thus, existing works (Perdomo et al.,, 2020; Mendler-Dünner et al.,, 2020; Zhao,, 2022) have mostly focused on finding the so-called performative stable (PS) point. In an FL system, the PS point ${{\theta}}^{\text{PS}}$ is defined as follows:
${{\theta}}^{\text{PS}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}({{\theta}}^{\text{PS}};\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}({{\theta}}^{\text{PS}};\theta)$ (2)
where $\mathcal{L}_{i}({{\theta}}^{{}^{\prime}};\theta)=\mathbb{E}_{Z_{i}\sim D_{i}({{\theta}}^{{}^{\prime}})}\left[\ell(Z_{i};\theta)\right]$ is called the decoupled performative risk (DPR) of client $i$. Unlike performative risk (PR), where the distribution $D_{i}(\theta)$ depends on $\theta$, the DPR decouples the two; i.e., the data distribution is induced by ${\theta}^{\text{PS}}$ while the variable being optimized is $\theta$. Since ${\theta}^{\text{PS}}$ is the fixed point of Eq. (2), it can be found via repeated risk minimization (RRM) or repeated gradient descent (RGD) (Perdomo et al.,, 2020). RRM recursively finds a model $\theta^{t+1}$ that minimizes DPR on the distribution $D(\theta^{t})$ induced by the previous model, whereas RGD recursively finds a model $\theta^{t+1}$ through gradient descent, and $\theta^{t+1}$ is not necessarily the minimizer of the DPR on $D(\theta^{t})$.
Following the idea of RGD, Jin et al., (2024) proposed the performative federated learning (PFL) algorithm to find the PS point ${\theta}^{\text{PS}}$ in FL. In PFL, each client $i$ updates its local model to reduce local DPR based on the previous model’s induced distribution: $\theta_{i}^{t+1}\leftarrow\theta_{i}^{t}-\eta_{t}\nabla\ell(Z_{i}^{t+1};\theta_{i}^{t})$, where $Z_{i}^{t+1}\sim D_{i}(\theta_{i}^{t})$ is drawn from the uncontaminated distribution. These updates are repeated for several epochs before being sent to the server for global aggregation. Jin et al., (2024) showed that PFL converges to ${\theta}^{\text{PS}}$ under clean data. Although the PS point is easier to find than the PO point, prior works (Izzo et al.,, 2021; Miller et al.,, 2021) show that the PS point may be far from optimal or may not even exist. Thus, algorithms like PFL, which target the PS point, are not ideal.
[…]
$\nabla\mathcal{L}_{i}(\theta)=\nabla\mathcal{L}_{i,1}(\theta)+\nabla\mathcal{L}_{i,2}(\theta)$ (3)
where
$\nabla\mathcal{L}_{i,1}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\mathbf{\theta})}[\nabla\ell(Z_{i};\mathbf{\theta})]$
and
$\nabla\mathcal{L}_{i,2}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}\left[\ell(Z_{i};\theta)\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}\right].$
Note that PFL proposed by Jin et al., (2024) only updates the model with $\nabla\mathcal{L}_{i,1}(\theta)$ while completely neglecting $\nabla\mathcal{L}_{i,2}(\theta)$. This error can accumulate, and eventually, PFL can converge to a PS point far from the optimal. Thus, computing the PG for a more accurate estimate is crucial, but estimating $\nabla\mathcal{L}_{i,2}(\theta)$ is challenging. Although Izzo et al., (2021) proposed a method to estimate the PG, its performance is only evaluated in a centralized setting under a specific performative distribution map $D_{i}(\theta)$ (i.e., when $f_{i}(\theta)$ is a linear function of $\theta$) without considering data contamination. Moreover, it assumes that PR is convex, which is often unrealistic.
In summary, an FL algorithm that yields the PO point, even under a clean distribution and with rather limiting assumptions, has not been developed previously. This paper fills that gap by proposing an FL algorithm that provably converges to the PO point, even with non-convex PR, and handles data contamination.
Remark 2.1.
FL finds the solution by alternating between global model aggregation at the central server and local model updates at the clients. In our proposed algorithms, we focus on the local updates, which work with different aggregation rules imposed by the server, such as differing weighted averages.
3 A Naive Extension from CS to FL
Consider first directly extending Izzo et al., (2021) to the FL setting. We shall call this naive algortihm Performative optimal Federated Learning (PoFL). The pseudocode is provided in Appendix A. In this algorithm, each active client $i$ in set $\mathcal{I}_{t}\subseteq\mathcal{V}$, after synchronization with the current global model $\overline{\theta}^{t}$, uses its local data to update the model from $\theta_{i}^{t}$ to $\theta_{i}^{t+R}$. During these updates, the data distribution changes, and consequently, client $i$ would use samples from the contaminated distribution $P_{i}(\theta_{i}^{t})$ to update $\theta_{i}^{t}$. To find the optimal ${\theta}^{\text{PO}}=\arg\min_{\theta}\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$, each client $i$ should estimate its local performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$ to update the local model $\theta_{i}^{t}$, i.e.,
$\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\nabla\mathcal{L}_{i}(\theta^{t}_{i})\right).$ (4)
By Eq. (3), $\nabla\mathcal{L}_{i}(\theta_{i}^{t})=\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ consists of two terms. PoFL estimates $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$ by simply averaging $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$. We denote the empirical risk on data drawn from distribution $Q_{i}$ or $D_{i}(\mathbf{\theta}_{i}^{t})$ as ^$\widehat{\nabla}\mathcal{L}_{i,1}(Q_{i},\theta_{i}^{t})$ and ^$\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$, respectively. Then the empirical risk on $P_{i}(\mathbf{\theta}_{i}^{t})$ is ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t}).$ Let ^$\widehat{\mathcal{L}}_{i,1}(\theta_{i}^{t})$ denotes the average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$. We denote ^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$, ^$\widehat{\nabla}\mathcal{L}_{i,2}(Q_{i},\theta_{i}^{t})$, and ^$\widehat{\nabla}\mathcal{L}_{i,2}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$ similarly. ^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ is estimated by a weighted average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$ with weight vector
$w(z_{j};\theta_{i}^{t}):=\left.\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p(z_{j};f_{i}(\theta))}{\partial f_{i}(\theta)}\right|_{\theta=\mathbf{\theta}_{i}^{t}}\in\mathbb{R}^{d}$ (5)
Because the distribution map $D_{i}(\theta)=p(z;f_{i}(\theta))$ and $f_{i}(\theta)$ are unknown, we need to estimate $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}.$ Following Izzo et al., (2021), we adopt finite difference approximation $\left.\frac{df_{i}(\theta)}{d\theta}\right|_{\theta=\theta^{t}_{i}}\approx\Delta f_{i}(\Delta\theta_{i})^{\dagger}$. Specifically, at iteration $t$, each client $i$ collects the $H$ most recent models (excluding $\theta^{t}_{i}$) and forms a matrix $\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}$, then
$\Delta\theta_{i}$ $=\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}-\theta^{t}_{i}\textbf{1}^{T}_{H};$ (6)
$\Delta f_{i}$ $=\begin{bmatrix}f_{i}(\theta^{t-H}_{i})&\cdots&f_{i}(\theta^{t-1}_{i})\end{bmatrix}-f_{i}(\theta^{t}_{i})\textbf{1}^{T}_{H},$
Where $\textbf{1}_{H}$ is an $H$-dimensional all-ones vector, both $\Delta f_{i}$ and $\Delta\theta_{i}$ are $d\times H$ matrices, and $(\Delta\theta_{i})^{\dagger}$ is the pseudo-inverse of $\Delta\theta_{i}$. For each iteration of each client, we know $\theta_{i}^{t}$, but do not know the exact value of $f_{i}(\theta_{i}^{t})$. Denote the estimator as $\widehat{f}_{i}(\mathcal{S}_{i})$, meaning we use the dataset $\mathcal{S}_{i}$ and the function $\widehat{f}_{i}(.)$ to estimate the value of $f_{i}(\theta_{i}^{t})$. This estimator is then used to compute an estimate of $\frac{df_{i}}{d\theta}$. The estimator of $\frac{df_{i}}{d\theta}$ is denoted as $\frac{\widehat{df_{i}}}{d\theta}$ and can be computed as $\frac{\widehat{df_{i}}}{d\theta}\approx\widehat{f}_{i}(\Delta\theta_{i})^{\dagger}$. Finally, $\frac{\widehat{df_{i}}}{d\theta}$ will be used to calculate ^$\widehat{\nabla}\mathcal{L}_{i,2}$. By combining this with ^$\widehat{\nabla}\mathcal{L}_{i,1}$, we obtain the performative gradient, allowing each client to update the local model $\theta_{i}^{t}$ using Eq. (4).
Concerns with POFL. Although Izzo et al., (2021) showed that the above method for estimating $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}$ was effective in centralized settings, it was evaluated under specific scenarios where $f_{i}(\theta)$ is linear in $\theta$ and data samples perfectly represent $D_{i}(\theta)$ without contamination. In practical FL settings with model-induced distribution shift, this method will perform poorly due to the following reasons:
  1. 1.
    Non-linearity of $f_{i}(\theta)$. When $f_{i}(\theta)$ is non-linear, ignoring the higher-order terms in the Taylor expansion of $\frac{\widehat{df_{i}}}{d\theta}$ may be biased. While Izzo et al., (2021) demonstrated that this method works well for linear $f_{i}(\theta)$, the impact of non-linearity on performance remains unclear.
  2. 2.
    Amplified sampling error. Since $\frac{df_{i}(\theta)}{d\theta}$ is estimated from limited samples, the sampling error of the estimation $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$ is inherited when computing PG, $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$. As shown in Lemmas 5.7, 5.8 and  5.9 in Section 5, the sampling error in estimating $\frac{df_{i}(\theta)}{d\theta}$ is amplified when computing $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$, making the performative gradient and algorithm more sensitive to sampling errors.
  3. 3.
    Contaminated data. Due to contamination $Q_{i}$, the data used for estimating $\frac{df(\theta)}{d\theta}$ and $f_{i}(\theta)$ does not follow the actual $D_{i}(\theta)$. Even a small error in a single model update for one client can cascade, affecting subsequent data and the entire network.
4 The Proposed Algorithm: PROFL
The previous section summarized the drawbacks of a straightforward extension of the centralized solution to the federated setting. To address all these weaknesses, we propose several simple yet effective strategies that lead to a robust new FL algorithm. We shall call it Performative robust optimal Federated Learning (ProFL). The pseudocode is provided in Algorithm 1. Below are its key features.
Algorithm 1 ProFL
1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$, error bound $\Phi_{i}$ for all $i\in\mathcal{V}$.
2:  for $t=0$ to $T-1$ do
3:     If $t\mod R=0$, server sends $\overline{\theta}^{t}$ to a selected set $\mathcal{I}_{t}\subseteq\mathcal{V}$ of clients and client $i\in\mathcal{I}_{t}$ sets the local model $\theta^{t}_{i}=\overline{\theta}^{t}$; otherwise $\mathcal{I}_{t}=\mathcal{I}_{t-1}.$
4:     for $i\in\mathcal{I}_{t}$ in parallel do
5:        Deploy $\theta^{t}_{i}$ and local environment changes and draw $n_{i}$ samples $\mathcal{S}_{i}:=(z_{j})^{n_{i}}_{j=1}\mathop{\sim}\limits^{iid}P_{i}(\theta^{t}_{i})$.
6:        if $t\leq H$ then
7:           ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right).$
8:           Compute estimator $\widehat{f}\left(\mathcal{S^{\prime}}_{i}\right)$ of $f^{t}_{i}$.
9:           Update $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})\right)$
10:        else
11:           ^$\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right)$
12:           Compute estimator $\widehat{f}\left(\mathcal{S^{\prime}}_{i}\right)$ of $f^{t}_{i}$.
13:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
14:           $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\left(\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})+\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})\right)\right)$.
15:           Input={^$=\{\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});H;d;\eta;\varphi_{i};\Delta f_{i}(\Delta\theta_{i})^{\dagger};\Phi_{i}\}$
16:           $n_{i}\leftarrow\textsc{Adaptive Sample Size}\left(\texttt{Input}\right)$.
17:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
18:  Output: $\overline{\theta}^{T}$
  1. 1.
    Handling complex distribution shifts with non-linear $f_{i}(\theta)$. ProFL employs the following strategies: (i) reducing the learning rate $\eta$ and the estimation window $H$; (ii) estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server. Reducing $\eta$ and $H$ helps reduce the estimation error resulting from non-linear $f_{i}(\theta)$ and stabilizes the algorithm. When the local sample size $n_{i}$ is small, $\widehat{f}_{i}(\mathcal{S}_{i})$ may significantly deviate from $f_{i}(\theta)$, leading to poor estimation of $\frac{df_{i}}{d\theta}$. Estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server mitigates the negative impact of sampling error.
  2. 2.
    Balancing the impact of sampling error and computational costs. ProFL adaptively selects the sample size $n_{i}$ based on the error bounds of the PG and local loss. While a larger $n_{i}$ generally leads to a more accurate estimate, it may not always be feasible in FL due to clients’ limited computational capabilities. By using predefined error bounds and relevant computed values—such as model variance, loss function, gradient information, and learning rates—we can determine the lower bound for each client’s $n_{i}$, balancing sampling error and computational costs. As these values change during updates, $n_{i}$ should be selected adaptively. More details are provided in Section 5.3.
  3. 3.
    Handling data contamination. ProFL mitigates contamination $Q_{i}$ by reducing $\epsilon_{i}$ through local outlier identification and removal. Since the samples $\mathcal{S}_{i}\sim P_{i}(\theta_{i}^{t})$are noisy, and the raw data is not visible to the central server, contamination must be identified and cleaned locally before estimating the performative gradient $\nabla\mathcal{L}_{i}=\nabla\mathcal{L}_{i,1}+\nabla\mathcal{L}_{i,2}$. Given the contaminated dataset $\mathcal{S}_{i}$, each client in ProFL uses algorithm Robust gradient to identify “clean” samples $\mathcal{S}^{\prime}_{i}=\{z_{j}\in\mathcal{S}_{i}:z_{j}\sim D_{i}(\theta_{i}^{t})\}$ and compute robust gradient $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$. $\mathcal{S}^{\prime}_{i}$ are then used to estimate $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$. The outlier identification and removal mechanism reduces $\epsilon_{i}$ (the fraction of contaminated data) and can be selected based on prior knowledge of $\epsilon_{i},Q_{i}$ (if there are any). In Appendix A, we provide an example of Robust Gradient for a general case with unknown $\epsilon_{i}$ and arbitrary $Q_{i}$. It uses singular value decomposition (SVD) to detect gradients that meet two key properties: 1) having a significant impact on the model and potentially disrupting training (i.e., gradients with large magnitudes that systematically point in a specific direction); and 2) being far from the average in the vector space (i.e., “long tail” data). We iteratively remove gradients whose distance from the average exceeds a threshold. More details on the Robust Gradient mechanism and its comparison to prior work are provided in Appendix  A.
In section 5, we analyze both algorithms’ convergence and the impact of the factors above on performance. Compared to PFL Jin et al., (2024) and methods in Izzo et al., (2021); Miller et al., (2021), we relax assumptions and show convergence under weaker conditions.
5 Performance Analysis
To facilitate analysis, we consider scenarios where $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}(\theta)$ represents a mixture of $K$ group distributions. $\nu_{i,k}(\theta)$ is the proportion of samples from group $k$ and $\sum_{k=1}^{K}\nu_{i,k}(\theta)=1$. Let $f_{i,k}(.)$ denote the function for each group’s distribution.
[…]
Assumption 5.1.
Let $W_{1}(D,D^{\prime})$ measure the Wasserstein-1 distance between two distributions $D$ and $D^{\prime}$. Then $\forall i\in\mathcal{V}$, there exists $\gamma_{i}>0$ such that $\forall\theta,\theta^{\prime}\in\Theta:W_{1}\left(D_{i}(\theta),D_{i}(\theta^{\prime})\right)\leq\gamma_{i}{\|\theta-\theta^{\prime}\|}.$
Assumption 5.2.
$\ell(z;\mathbf{\theta})$ is continuously differentiable and $L$-smooth, i.e., for all $\theta\in\Theta\text{ and }z,z^{\prime}\in\mathcal{Z}:\|\nabla\ell(z;\mathbf{\theta})-\nabla\ell(z^{\prime};\mathbf{\theta^{\prime}})\|\leq L(\|\mathbf{\theta}-\mathbf{\theta^{\prime}}\|+\|z-z^{\prime}\|).$
Assumption 5.3.
$f_{i}(\theta)$ is twice continuously differentiable for all $\theta\in\Theta,$ i.e., the first and second derivatives $\frac{df_{i}(\theta)}{d\theta}$ and $\frac{d^{2}f_{i}(\theta)}{d\theta^{2}}$ exist and are continuous.
Assumption 5.4 (PL Condition Polyak, (1964)).
The local performative risk $\mathcal{L}_{i}(\theta)$ of client $i$ satisfies Polyak- Lojasiewicz (PL) condition, i.e., for all $\theta\in\Theta$, the following holds for some $\rho>0$:
$\frac{1}{2}\|\nabla\mathcal{L}_{i}(\theta)\|^{2}\geq\rho\left(\mathcal{L}_{i}(\theta)-\mathcal{L}_{i}(\theta_{i}^{\text{PO}})\right).$
Remark 5.5.
Unlike most works that require convex performative risk, we demonstrate the convergence of our algorithm under a weaker PL condition, which permits the performative risk to be non-convex.
5.1 Error of the performative gradient
As discussed in Section 3, the key to converging to $\theta^{\text{PO}}$ is estimating the performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})=\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ in Eq. (3) accurately. Thus, we analyze the estimation errors of $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$ and $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ and explore how these errors are influenced by the non-linearity of $f_{i}(\theta)$, sampling error, and contaminated data $Q_{i}$. The results of this analysis will then be used to assess the convergence of ProFL (and PoFL) in Section 5.2.
Lemma 5.6.
Under Assumption 5.3, there exist $F,M<\infty$ such that $\left\|\frac{df_{i}(\theta)}{d\theta}\right\|\leq F$ and $\left\|\frac{d^{2}f_{i}(\theta)}{d\theta^{2}}\right\|\leq M$ hold for all $i\in\mathcal{V}.$ Under Assumption 5.2, there exist $G,\ell_{\max}<\infty$ such that $\|\nabla\ell(z;\mathbf{\theta})\|\leq G$ and $\ell(z;\mathbf{\theta})\leq\ell_{\max}$ hold for all $z\in\mathcal{Z}$.
Lemma 5.6 is proved using Weierstrass’s theorem. We will use $F,M,G,\ell_{\max}$ to denote the upper bounds of these quantities in the rest of the paper.
Lemma 5.7 (Estimation error of $\frac{\widehat{df_{i}}}{d\theta}$).
Under Assumption 5.3, let $\lambda_{i,\min}$ denote the minimal singular value of $\Delta\theta_{i}$ defined in Eq. (6) during all iterations. With probability at least $1-\varphi_{i}$, the estimation error of $\frac{\widehat{df_{i}}}{d\theta}$ is bounded by $\boldsymbol{\omega_{F}}$ specified as follows:
$\mathbb{E}\left[\left\|\frac{\widehat{df_{i}}}{d\theta}-\frac{df_{i}}{d\theta}\right\|^{2}\right]\leq\boldsymbol{\omega_{F}}:=2\epsilon_{i}^{2}F^{2}+(1-\epsilon_{i})^{2}\left(\frac{M^{2}\eta^{4}G^{4}H^{6}}{2}+\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}\right)\lambda^{-2}_{i,\min}.$ (7)
Notably, $f_{\varphi}(\varphi_{i};n_{i})$ is a function that depends on the properties of the distribution. For instance, $f_{\varphi}(\varphi_{i};n_{i})=\mathcal{O}(-\log\varphi_{i})$ in distribution dynamics, and $f_{\varphi}(\varphi_{i};n_{i})=\mathcal{O}(1/\varphi_{i})$ in contribution dynamics.
Lemma 5.8 (Estimation error of $\nabla\mathcal{L}_{i,1}$).
Under Assumption 5.3, with probability at least $1-\varphi_{i}$, we have:
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]\leq\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right).$
Lemma 5.9 (Estimation error of $\nabla\mathcal{L}_{i,2}$).
Under Assumptions 5.1, 5.2 and 5.3, the following holds with probability at least $1-\varphi_{i}$:
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]\leq(\boldsymbol{\omega_{D}})^{2}:=\mathcal{O}\left(\ell^{2}_{\max}d\left(\boldsymbol{\omega_{F}}+\frac{f_{\varphi}(\varphi_{i};n_{i})F^{2}}{n_{i}}\right)\right).$
$\mathbb{E}\left[\left\|{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]\leq(\boldsymbol{\omega_{Q}})^{2}:=\mathcal{O}\left(2\ell^{2}_{\max}d\left(\boldsymbol{\omega_{F}}+F^{2}\right)\right).$
Additionally, the notations $\boldsymbol{\omega_{D}}$ and $\boldsymbol{\omega_{Q}}$ represent upper bounds. From Lemmas 5.8 and 5.9, we find that since the estimation error of performative gradient is primarily dominated by the estimation error of $\nabla\mathcal{L}_{i,2}$, we have:
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]+\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]\leq(\boldsymbol{\omega_{D}})^{2}.$ (8)
5.2 Convergence analysis
Given Lemmas 5.7, 5.8, and 5.9, we are now ready to analyze the convergence. Denote $\overline{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}\gamma_{i}$, $\sigma^{2}_{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}(\overline{\gamma}-\gamma_{i})^{2}$, $\overline{\epsilon}=\sum_{i\in\mathcal{V}}\alpha_{i}\epsilon_{i}$, and $\varphi=\sum_{i\in\mathcal{V}}\varphi_{i}.$ Throughout the training process, let $W_{1}(D,Q)_{\max}$ represent the maximum Wasserstein-1 distance $W_{1}\left(D(\theta_{i}^{t}),Q_{i}\right)$ between actual data distribution and contamination $\forall i\in\mathcal{V}$. Let $\overline{\boldsymbol{\omega_{Q}}},\overline{\boldsymbol{\omega_{D}}}$ be the maximum of $\boldsymbol{\omega_{Q}}$ and $\boldsymbol{\omega_{D}}$ respectively in Lemma 5.9 $\forall i\in\mathcal{V}$. Let $g^{2}_{\min}$ be the minimal value of $\mathbb{E}[\|g^{t}\|^{2}]$ with $g^{t}=\sum_{i\in\mathcal{V}}\alpha_{i}\sum_{k=1}^{K}\widehat{\nabla}\mathcal{L}_{{i,k}}.$
Theorem 5.10 (Convergence rate).
Let $\{\overline{\theta}^{t}\}_{t\geq 0}$ be a sequence of global models generated by ProFL (and PoFL). Under Assumptions 5.1, 5.2, 5.3, and 5.4, with probability at least $1-T\varphi$, we have:
$\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{T}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]\leq$ $\left(1-\rho\eta\right)^{T}\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{0}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}-\frac{g^{2}_{\min}}{2}\right)\eta$
$\quad+\left(L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
$\quad+\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\eta^{3}.$
where $\overline{\boldsymbol{\omega_{D,Q}}}=\overline{\boldsymbol{\omega_{D}}}+LW_{1}(D,Q)_{\max}+\overline{\boldsymbol{\omega_{Q}}}.$
Theorem 5.10, together with Lemmas 5.7, 5.8 and 5.9 in Section 5.1, highlight the impact of sampling error, data contamination, client heterogeneity, and the non-linearity of $f_{i}(\theta)$ on convergence, as discussed below.
When there is no data contamination and $\epsilon_{i}=0$, $\boldsymbol{\omega_{F}}$ is reduced to $\left(\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}+\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}\right)\lambda_{i,\min}^{-2}$, where the first term is due to the non-linearity of $f_{i}(\theta)$ and becomes zero if $f_{i}(\theta)$ is linear with $M=0$; the second term, related to sampling error, approaches zero as the sample size $n_{i}\to\infty$. With sufficient samples, the error term $\frac{f_{\varphi}(\varphi_{i};n_{i})F^{2}}{n_{i}}$ in $\boldsymbol{\omega_{D}}$ also approaches zero.
When there is contamination or $f_{i}(\theta)$ is non-linear, the error always exists even with a sufficiently large sample size. The term $\overline{\epsilon}(\overline{\boldsymbol{\omega_{D}}}+LW_{1}(D,Q)_{\max}+\overline{\boldsymbol{\omega_{Q}}})^{2}$ reflects the impact of contaminated data, increasing as the fraction of contamination $\epsilon_{i}$ and/or the distance between actual data distribution $D_{i}(\theta)$ and contamination $Q_{i}$ increase.
The term $L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$ captures the effect of client heterogeneity. Recall that $R$ is the number of local updates. In the special case where $R=1$ , meaning clients update and aggregate models at each iteration, this term becomes zero.
Corollary 5.11.
If $\overline{\epsilon}=0$ (without contamination), the bound in Theorem 5.10 is reduced to the following:
$\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq$ $(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right]+\left(\overline{\boldsymbol{\omega_{D}}}^{2}-\frac{g^{2}_{\min}}{2}\right)\eta+L(1+\overline{\gamma})\left(\overline{\boldsymbol{\omega_{D}}}^{2}+G^{2}\right)\eta^{2}$
$\quad+\left(\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)\eta^{3}.$
Because $g^{2}_{\min}\geq 0,$ we consider two cases:
  1. 1.
    If $g^{2}_{\min}>0$, we can always adjust learning rate $\eta$, estimation window $H$, number of local updates $R$, and sample size $n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
    $\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right],$
    i.e., $\overline{\theta}^{t}$ converges to ${\theta}^{\text{PO}}$ at linear convergence rate.
  2. 2.
    If $g^{2}_{\min}=0$, this means there is at least one iteration $t$ that has $\|g^{t}\|=0$ and $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$. Except for the iterations where $\|g^{t}\|=0$, all the other iterations have positive $\|g^{t}\|>0$. We can choose $\eta,H,n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
    $\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T-b}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right],$
    where $b$ is the number of iterations with $\|g^{t}\|=0$.
All above theoretical results apply to both ProFL and PoFL. The robust strategies in ProFL reduce the upper bound in Theorem 5.10.
5.3 Discussion
Section 3 introduced simple yet effective strategies in ProFL to enhance PoFL. We now explain why these strategies are effective.
[…]
Balancing Sampling Error and Computational Costs. We know that $f_{\varphi}(\varphi_{i};n_{i})$ depends on the distribution. Once this information, along with the error bound $\Phi_{i}$, is known, we can calculate the lower bound of $n_{i}$. For example, when $d=1$, the estimator $\widehat{f}_{i}(\theta)$ in distribution dynamics estimates the mean of $D_{i}(\theta)$. In this case, $f_{\varphi}(\varphi_{i};n_{i})=\sigma^{2}_{i}\log(2/\varphi_{i})$, where $\sigma^{2}_{i}$ represents the data variance of client $i$. With the error bound $\Phi_{i}$, ProFL can compute $n_{i}$ as follows:
$n_{i}\geq\frac{2\left(2\ell^{2}_{\max}H\|\Delta\theta\|^{-2}+F^{2}\right)\sigma^{2}_{i}\log(2/\varphi_{i})}{2\Phi_{i}-M^{2}\eta^{2}G^{4}H^{6}\|\Delta\theta\|^{2}\ell^{2}_{\max}}.$
Here, $H$, $\varphi_{i}$, $\|\Delta\theta\|$, and $\eta$ are known, while $M$, $G$, and $F$ can be estimated from previous updates. For $\ell_{\max}$, we use the loss from the last iteration, as the loss typically decreases during training. The variance $\sigma^{2}_{i}$ is estimated from $S_{i}$ in the last iteration.
Handling contaminated data. Theorem 5.10 shows that contaminated data affects convergence mainly through the term $\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D}}}+LW_{1}(D,Q)_{\max}+\overline{\boldsymbol{\omega_{Q}}}\right)^{2}$, which is difficult to mitigate by tuning parameters. A more effective approach is to reduce $\overline{\epsilon}$ by removing contamination. For instance, if all contaminated data is removed ($\overline{\epsilon}=0$), ProFL converges to the global PO point, as stated in Corollary 5.11.
6 Experiments
Experimental Setup. We evaluate our algorithm in five case studies: 1) pricing with dynamic demands, 2) pricing with dynamic contribution, 3) binary strategic classification, 4) house pricing regression, and 5) regression with dynamic contribution.
Pricing with Dynamic Demands. Consider a company that aims to find the proper prices for various goods. Let $\theta\in\mathbb{R}^{d}$ be the prices associated with the set of $d$ goods, and $Z_{i}\in\mathbb{R}^{d}$ be the retailer $i$’s demands for these goods. We assume $Z_{i}\sim\mathcal{N}(\mu(\theta),\sigma^{2})$ is model dependent Gaussian distribution, i.e., mean demand for each good decreases linearly as the price increases. $\mu(\theta)=\mu_{0}-\gamma_{i}\theta$, here $\gamma_{i}\geq 0$ measures the price sensitivity and varies across different retailers. The goal is to maximizes total revenue over all retailers $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\theta^{T}Z_{i}]$.
Pricing with Dynamic Contribution. Consider a similar setting as above where a company aims to find prices for various goods that maximize the total revenue. Suppose the retailers it faces have consumers from $K$ groups (e.g., old/young) and these consumers have fixed demands but they have options to purchase the goods from other companies. Here we assume the demands of retailer $i$’s consumers from group $k\in[K]$ have fixed distribution $D_{i,k}=\mathcal{N}(\mu_{i,k},\sigma_{i}^{2})$, but the fraction of groups change dynamically based on prices $\theta$. Let $r_{i,k}(\theta)=B-\mathbb{E}_{Z_{i}\sim D_{i,k}}[\theta^{T}Z_{i}]$ be the expected remaining budget of group $k$ after purchasing the goods (with initial budget $B$) and assume the fraction of group $k$ is $\nu_{i,k}(\theta)=\frac{r_{i,k}(\theta)}{\sum_{k=1}^{K}r_{i,k}(\theta)}$ (i.e., groups with more remaining budgets have more incentives to stay in the system). Then the retailer $i$’s total demands would follow the mixture of distributions $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}$.
Binary Strategic Classification. Consider an FL system is used to make binary decisions about human agents (e.g., loan application, hiring, college admission), where client $i$ with data $Z_{i}=(X_{i},Y_{i})$ may manipulate their features strategically to increasing their chances of receiving favorable outcomes at the lowest costs. Suppose the clients are assigned positive decisions whenever $21/\left(1+\exp(-X^{T}\theta)\right)\geq 1/$ 2 and the cost it takes to manipulate feature from $X_{i}$ to $X_{i}^{\prime}$ is $\frac{1}{2\gamma_{i}}||X^{\prime}-X||$, then strategic clients would react to the decision model $\theta$ by manipulating their features to $\arg\max_{X_{i}^{\prime}}\left(-\langle X^{\prime}_{i},\theta\rangle-\frac{1}{2\gamma_{i}}\|X_{i}^{\prime}-X_{i}\|_{2}^{2}\right)=X_{i}-\gamma_{i}\theta$. For the initial data, we acquire it from a synthetic data where $X|Y=y\sim\mathcal{N}(\mu_{y},\sigma_{y}^{2})$ for $y\in\{0,1\}$ and two realistic Adult and Give Me Some Credit data. The details of these data are in Appendix E.1. The goal is to minimizes prediction loss over all clients.
House Pricing Regression. Consider a regression task that aims to find the proper listing prices for houses based on features such as size, age, number of bedrooms. Suppose the features of houses in district $i$ follow $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. Given predictive model $\theta$, let the listing price be $h(X_{i})=\theta^{T}X_{i}$. Since the houses with higher listing prices tend to lower the demand, the actual selling price $Y$ depends on $\theta$ and we assume $Y_{i}=(a_{i}-\gamma_{i}\theta)^{T}X_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{n})$ is Gaussian noise. The goal is to minimizes prediction error
Regression with dynamic contribution. Consider an example of retail inventory management, where the regression model $\theta$ predicts product sales considering features like seasonal trends and advertising. Distributor $i$ has retailers from two groups $k\in\{1,2\}$, and each retailer’s data distribution is fixed. The fraction of retailers from each group is $\nu_{i,k}(\theta)=\frac{\ell_{i,-k}(\theta)+c}{\sum_{k^{\prime}\in\{1,2\}}\left(\ell_{i,k^{\prime}}(\theta)+c\right)}$ where $-k=\{1,2\}\setminus k$ because the retailers are more likely to stop purchasing if they experience higher prediction error $\ell_{i,k}(\theta)$. The goal is to find $\theta$ that minimizes expected prediction error $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{(X_{i},Y_{i})\sim D_{i}(\theta)}\left[(\theta^{T}X_{i}-Y_{i})^{2}\right]$, where $Y_{i}$ is a weighted combination of fixed distributions.
For binary classification, we use both synthetic and real data (Adult and Give Me Some Credit datasets), while the other cases use synthetic data. All dynamics are synthetic. The parameter choices is in Table 3 of Appendix E.1. Information on the real datasets is in Appendix E.4. Each method is run 10 times under the same setup, and averages and variance are reported.
Baselines. To date, only Jin et al., (2024) has addressed model-dependent distribution shifts in FL. Thus, we compare our algorithm with this Performative Federated Learning (PFL). Unlike our algorithm, theirs converges to a performative stable solution ${\theta}^{\text{PS}}$ under stricter assumptions. We also compare the Performative Gradient (PG) algorithm from Izzo et al., (2021), in the centralized setting.
Table 1: Performative Loss of PG and ProFL
$\alpha$ 0 0.25 0.5
PG 5.56 $\pm$ 0.00 5.67 $\pm$ 0.02 6.32 $\pm$ 0.06
PROFL 5.56 $\pm$ 0.00 5.59 $\pm$ 0.00 5.67 $\pm$ 0.00
(a) Depend on Learning Rates
(b) Estimation on Server
(c) Adaptive Learning Rate
Figure 1: Effectiveness of Proposed methods
(a) Different Sample Sizes
(b) Different Learning Rates
(c) Different H
Figure 2: Dynamic contribution with Different Sample Sizes, Learning Rates, and $H$
Comparison to Centralized Setting. Table 1 shows that adapting performative gradients to the FL framework improves performance. There are 10 clients with $\gamma_{i}\in[\overline{\gamma}-\alpha,\overline{\gamma}+\alpha]$, where larger $\alpha$ indicates more heterogeneity. As client heterogeneity increases, the performance of both algorithms degrades. Under homogeneous distribution shifts ($\alpha=0$), both algorithms perform similarly. ProFL (ours) outperforms PG when $\alpha=0.25$ or $\alpha=0.5$, as the centralized approach struggles to handle variance shifts by estimating a single $\frac{df}{d\theta}$ on the server.
Effectiveness of our proposed methods. Fig. 1(a) shows that contamination can significantly degrade performance, but the proposed robust gradient can effectively reduce the impact of contaminated data when computing the gradients i.e., the performance with the robust gradient in our method is almost the same as the case when outliers do not exist at all. This case study, pricing with dynamic demands, involves 10 heterogeneous clients with different $D_{i}(\theta)$, each encounters an $\epsilon_{i}$ fraction of contaminated data from a fixed but unknown exogenous distribution $Q_{i}$. Although we present the results when $Q_{i}\sim\mathcal{N}(\mu_{o},\sigma^{2}_{o})$, we observed similar results when $Q_{i}$ follows other distributions and are different among clients.
Fig. 1(b) shows that with a limited sample size and many clients ($n_{i}=60,|\mathcal{V}|=100$), estimating $\widehat{\frac{df_{i}}{d\theta}}$ on the server side results in a lower loss value compared to local estimation and approaches the optimal value. Fig. 1(c) demonstrates the effectiveness of adaptive sample sizing. We tested two tolerances (0.05 for purple, 0.1 for red), with each client processing up to 1000 samples per iteration. The adaptive approach significantly reduces the sample size while achieving similar results within the error tolerance, leading to substantial computational savings. Both experiments are based on the case study pricing with dynamic contribution, involving a non-linear $f_{i}(\theta)$.
Impact of hyperparameters. We examine the effects of sample size $n_{i}$, learning rate $\eta$, and estimation window $H$. Fig. 2(a) shows that increasing $n_{i}$ improves stability and brings the algorithm closer to the optimal point. Fig. 2(b) indicates that a smaller $\eta$ slows convergence but moves $\theta$ nearer to the optimum. Fig. 2(c) demonstrates that a smaller $H$ results in estimates estimates of $\frac{df_{i}}{d\theta}$ based on closer but smaller groups of $\widehat{f_{i}}$ and $\theta_{i}$, resulting in convergence closer to the performative optimum, albeit at a slightly slower speed.
Table 2: Accuracy of Binary Classification on Synthetic and Realistic Datasets
DATASET POFL PROFL
Same Distribution 88.00 $\pm$ 2.69 99.23 $\pm$ 0.29
Different Distributions 62.44 $\pm$ 0.52 92.50 $\pm$ 0.29
Credit Dataset 94.73 $\pm$ 0.00 94.73 $\pm$ 0.00
Adult Dataset 54.74 $\pm$ 0.02 60.78 $\pm$ 0.81
Convergence to the performative optimal solution. Table 2 shows that ProFL outperforms PFL in the case study of binary strategic classification with 10 heterogeneous clients. Since PFL cannot handle contaminated data, we set $\epsilon_{i}=0$ for these experiments. The first two rows show that ProFL is more stable and significantly more accurate than PFL on Gaussian synthetic data. PFL converges to a PS point rather than a PO point. On the Give Me Some Credit dataset (third row), both algorithms converge to the same point, but ProFL has a faster convergence rate (see Fig. 7(c) in the Appendix). With the Adult dataset, ProFL achieves $60.78\%$ accuracy, outperforming PFL’s $54.74\%$. These results highlight ProFL’s practical advantages in both convergence speed and accuracy. More results are in Appendix D.1.
Appendix A Algorithms and Details
A.1 Pseudocode of the PoFL
Algorithm 2 presents the pseudocode of the PoFL algorithm discussed in the main paper.
Algorithm 2 PoFL
1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$.
2:  for $t=0$ to $T-1$ do
3:     If $t\mod R=0$, server sends $\overline{\theta}^{t}$ to a selected set $\mathcal{I}_{t}\subseteq\mathcal{V}$ of clients and client $i\in\mathcal{I}_{t}$ sets the local model $\theta^{t}_{i}=\overline{\theta}^{t}$; otherwise $\mathcal{I}_{t}=\mathcal{I}_{t-1}.$
4:     for $i\in\mathcal{I}_{t}$ in parallel do
5:        Deploy $\theta^{t}_{i}$ and local environment changes and draw $n_{i}$ samples $\mathcal{S}_{i}:=(z_{j})^{n_{i}}_{j=1}\mathop{\sim}\limits^{iid}P_{i}(\theta^{t}_{i})$.
6:        if $t\leq H$ then
7:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
8:           Update $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})\right)$.
9:        else
10:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
11:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
12:           $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\left(\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})+\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})\right)\right)$.
13:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
14:  Output: $\overline{\theta}^{T}$
Algorithm 3 Robust Gradient
1:  Input: Set $\mathcal{S}_{i}$ of samples $(z_{j})_{j=1}^{n_{i}}$, thresholds $C,J\in(0,1)$, model $\theta^{t}_{i}\in\Theta$
2:  Let ^$\widehat{\nabla}\leftarrow\frac{1}{|\mathcal{S}_{i}|}\sum_{z_{j}\in\mathcal{S}_{i}}\nabla\ell(z_{j};\theta^{t}_{i})$.
3:  Let $\left[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right]_{z_{j}\in\mathcal{S}_{i}}$ be the $|\mathcal{S}_{i}|\times d$ matrix of centered gradients.
4:  Apply SVD to this matrix and find top right singular vector $v$.
5:  Compute the outlier score $\tau_{j}$ of $\nabla\ell(z_{j};\theta^{t}_{i})$ by Eq. (9).
6:  Divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments. Let $\phi_{k}$ be the number of $\tau_{j}$ in the $k$-th segment.
7:  Find the smallest $k\in\{1,\cdots,B\}$ with $\phi_{k}<C\cdot|\mathcal{S}_{i}|$. Set $\phi$ as the lower bound of the $k$-th segment.
8:  Find the set $\mathcal{S}_{i}^{\prime}\leftarrow\{z_{j}\in\mathcal{S}_{i}:\tau_{j}<\phi\}$.
9:  Compare the average gradients $\widehat{\nabla}$ of $\mathcal{S}_{i}$ and $\widehat{\nabla}^{\prime}$ of $\mathcal{S}^{\prime}_{i}$. If ^^^$\frac{\|\widehat{\nabla}-\widehat{\nabla}^{\prime}\|}{\|\widehat{\nabla}\|}<J$ or $|\mathcal{S}^{\prime}_{i}|\leq\frac{n_{i}}{2}$, return $\mathcal{S}^{\prime}_{i}$ and $\widehat{\nabla}^{\prime}$; otherwise, set $\mathcal{S}_{i}\leftarrow\mathcal{S}^{\prime}_{i}$ and repeat from line 2.
A.2 Outlier removal mechanism for robust gradient
To estimate $\nabla\mathcal{L}_{i,1}$, we should average $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $D_{i}({\theta}_{i}^{t})$. Because the samples collected by client $i$ in round $t$ are drawn from $P_{i}(\theta_{i}^{t})$, they are noisy, with an unknown $\epsilon_{i}$ fraction of outliers from an unknown distribution $Q_{i}$. ProFL needs to identify contaminated data and eliminate the corresponding contaminated gradients. We propose a mechanism that can remove these contaminated gradients and estimate gradients that are robust to outliers. Importantly, ProFL is independent of the specific outlier identification and removal mechanism used and can function with any such mechanism. In our paper, we assume that contaminated data can follow an arbitrary distribution. In practice, additional information about the contamination process may be available based on the application or context, allowing us to design more efficient (and less expensive) outlier identification mechanisms leveraging this knowledge.
As shown in Algorithm 3, our mechanism takes noisy samples $\mathcal{S}_{i}$ and current local model $\theta_{i}^{t}$ as inputs. The mechanism first computes the gradients $\nabla\ell(z_{j};\theta_{i}^{t})$ of all samples and iteratively identifies and removes the contaminated gradients.
Since both the distribution $Q_{i}$ and fraction $\epsilon_{i}$ of contaminated data are unknown, we consider gradients contaminated if they exhibit two crucial properties: 1) they have large effects on the learned model and can disturb the training process (i.e., gradients have large magnitudes and systematically point in a specific direction); 2) they are located farther away from the average in the vector space (i.e., “long tail” data). To detect gradients that satisfy these properties, we adapt the approach in Shan et al., (2023), which leverages singular value decomposition (SVD). Instead of preselecting the threshold for the sum of all outlier scores $\tau_{j}$, which is difficult to determine without any prior knowledge of the data, our threshold is based on the two crucial properties, making it much easier to predict the values of $C$ and $J$. Specifically, we construct a matrix $[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}]_{z_{j}\in\mathcal{S}_{i}}$ using the centered gradients and find the top right singular vector $v$, where $\widehat{\nabla}$ represents the average of all gradients (lines 2-4). The centered gradients that are closer to $v$ are more likely to be outliers and we assign each gradient an outlier score $\tau_{j}$ as follows:
$\tau_{j}={\left(\left(\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right)^{T}v\right)}^{2}$ (9)
Given the outlier scores $\{\tau_{j}\}$, the gradients with scores above a threshold $\phi$ are considered contaminated and are discarded. Unlike Shan et al., (2023) that uses a fixed pre-defined threshold, our method finds $\phi$ automatically based on the outlier score distribution. Since we assume the outliers are in the “long tail” of the distribution, we divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments and compute the relative frequency distribution of scores. $\phi$ is set as the smallest segment with the relative frequency below a threshold (lines 6-7). For the remaining gradients with $\tau_{j}<\phi$, we compute their average $\widehat{\nabla}^{\prime}$ (lines 8-9).
Appendix B Additional Related Work
B.1 Performative Prediction in Centralized Settings
Performative prediction (Perdomo et al.,, 2020) was first introduced in 2020 corresponding to the optimization framework to deal with endogenous data distribution shifts where the model deployed to the environment will affect the subsequent data, resulting in the collected data changing as the deployed model changes. Common applications of performative prediction include strategic classification(Hardt et al.,, 2016; Xie and Zhang,, 2024) and predictive policing (Ensign et al.,, 2018). Perdomo et al., (2020) first proposes an iterative optimization procedure named Repeated Risk Minimizationto find a performative stable point and also bounded the distance between the performative stable point and the performative optimal point. Mendler-Dünner et al., (2020) designed the first algorithm to find the performative stable point under the online learning setting. Miller et al., (2021) first solved the problem of performative prediction by directly optimizing performance risk to find the performative optimal point, but the scope of the distribution map is restricted. Mendler-Dünner et al., (2022) proved the convergence of greedy deployment and lazy deployment after each random update under the assumption of smoothness and strong convexity and gave the convergence rate. Izzo et al., (2021) also designed an algorithm performative gradient descent for finding the performative optimal solution. Zhao, (2022) relaxes the assumption of strong convexity to the weakly-strongly convex case and proves the convergence of the deployment. Somerstep et al., (2024); Bracale et al., (2024) focused on learning the distribution map in performative prediction and proposed a framework named Reverse Causal Performative Prediction. In addition, performative prediction is also related to reinforcement learning(Zheng et al.,, 2022) and bandit problems(Chen et al.,, 2024).
B.2 Performative Prediction in Distributed System
To the best of our knowledge, Jin et al., (2024) is the only work discussing performativity under the distributed setting, where they generalized Fedavg to P-Fedavg and proved the uniqueness of the performative stable solution found by the algorithm. They also quantified the distance between the performative stable solution and the performative optimal solution. While the work is a breakthrough as the first to take model-dependent distribution shifts into account under the FL setting, it is only a strict generalization of Perdomo et al., (2020) without the ability to find the performative optimal point and deal with the contaminated data.
B.3 Federated Learning under Dynamic Data
Previous works (Tan et al.,, 2023; Li et al., 2020a, ) pointed out that the training paradigm of FedAvg may violate the i.i.d. assumptions and cause some feature classes to be over-represented while others to be under-represented. Li et al., 2020b ; Li et al., (2021) steer the local models towards a global model by adding a regularization term to guarantee convergence when the data distributions among different clients are non-IID. Li et al., 2020b proposes FedProx as a solution. Li et al., (2021) aims at obtaining good local models on clients rather than a global consensus model. Meanwhile, another line of work focused on dealing with the statistical heterogeneity by clustering (Briggs et al.,, 2020; Ghosh et al.,, 2020; Sattler et al.,, 2020) by aggregating clients with similar distribution into the same cluster. While most previous works assumed the “dynamic” is among different clients, another line of literature focuses on FL in a dynamic environment where various distribution shifts occur. Some works aim to deal with time-varying contribution rates of clients with local heterogeneity (Sattler et al.,, 2021; Park et al.,, 2021; Wang and Ji,, 2022; Zhu et al.,, 2021).
B.4 Huber’s $\epsilon$-contamination Model
Proposed by Huber, (1992), Huber’s epsilon contamination model serves as a framework for robust statistical analysis with outliers and contaminations in the dataset. This model assumes that besides the data distribution we aim to learn, a small proportion of samples can come from an arbitrary distribution. This contamination framework is crucial for creating machine learning algorithms robust to outliers, thereby ensuring more reliable analysis in practical scenarios where data imperfections are common. Huber’s model has been fundamental in the field of robust statistics, influencing a wide range of applications and subsequent research. Specifically, denote $D$ as the true data distribution and $Q$ as a random distribution. $P=(1-\epsilon)D+Q$ is the distribution with some corruption $Q$ and the portion of corruption data is $\epsilon$. The local dataset of client $i$ is $P_{i}=(1-\epsilon_{i})D_{i}(\theta_{i})+\epsilon_{i}Q_{i}$.
Appendix C Additional Proofs not Detailed in the Main Paper
C.1 Proof of Theorem 5.10
Proof.
With Assumption 5.1 of $\gamma_{i}$-sensitivity and Assumption 5.2 of $L$-smoothness, we obtain
$\|\nabla\ell(z;\mathbf{\theta})-\nabla\ell(z^{\prime};\mathbf{\theta^{\prime}})\|\leq L(\|\mathbf{\theta}-\mathbf{\theta^{\prime}}\|+\|z-z^{\prime}\|)\leq L(1+\gamma_{i})\|\mathbf{\theta}-\mathbf{\theta^{\prime}}\|$
Then we have
$\mathcal{L}(\overline{\theta}^{t+1})\leq\mathcal{L}(\overline{\theta}^{t})+\langle\nabla\mathcal{L}(\overline{\theta}^{t}),\overline{\theta}^{t+1}-\overline{\theta}^{t}\rangle+\frac{L\sum^{N}_{i=1}\alpha_{i}(1+\gamma_{i})}{2}{\|\overline{\theta}^{t+1}-\overline{\theta}^{t}\|}^{2},$
where $\overline{\theta}^{t+1}-\overline{\theta}^{t}=-g^{t}\eta$ and $g^{t}=\sum_{i=1}^{N}\alpha_{i}g^{t}_{i}$.
By taking expectation on both sides we obtain that
$\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})]$
$\leq\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})]+\mathbb{E}\left[\langle\nabla\mathcal{L}(\overline{\theta}^{t}),\overline{\theta}^{t+1}-\overline{\theta}^{t}\rangle\right]+\frac{L\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})}{2}\mathbb{E}\left[\|\overline{\theta}^{t+1}-\overline{\theta}^{t}\|^{2}\right]$
$=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})]\underbrace{-\mathbb{E}\left[\langle\nabla\mathcal{L}(\overline{\theta}^{t}),g^{t}\rangle\right]}_{T_{1}}\eta+\frac{L\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})}{2}\underbrace{\mathbb{E}\left[\|g^{t}\|^{2}\right]}_{T_{2}}\eta^{2}.$ (10)
Let’s start to find the upper bound of $T_{2}$. With $R$ steps local updates, by using Jensen’s inequality, we obtain that
$\mathbb{E}\left[\|g^{t}\|^{2}\right]$
$=\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}g^{t}_{i}\right\|^{2}\right]$
$\leq\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\left(g^{t}_{i}-\nabla\mathcal{L}_{i}(\theta^{t}_{i})+\nabla\mathcal{L}_{i}(\theta^{t}_{i})\right)\right\|^{2}\right]$
$\leq 2\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\left(g^{t}_{i}-\nabla\mathcal{L}_{i}(\theta^{t}_{i})\right)\right\|^{2}\right]+2\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}(\theta^{t}_{i})\right\|^{2}\right]$
$\leq 2\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\left(g^{t}_{i}-\nabla\mathcal{L}_{i}(\theta^{t}_{i})\right)\right\|^{2}\right]+2G^{2}.$
Then we start to find the upper bound of $T_{1}$.
$-\mathbb{E}\left[\langle\nabla\mathcal{L}(\overline{\theta}^{t}),g^{t}\rangle\right]$
$=-\frac{1}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\overline{\theta}^{t})\right\|^{2}\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+\frac{1}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\overline{\theta}^{t})-g^{t}\right\|^{2}\right]$
$=-\frac{1}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\overline{\theta}^{t})\right\|^{2}\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+\frac{1}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\overline{\theta}^{t})-\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})+\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$
$\leq-\rho\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\left(\nabla\mathcal{L}_{i}(\overline{\theta}^{t})-\nabla\mathcal{L}_{i}({\theta}^{t}_{i})\right)\right\|^{2}\right]+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$
$\leq-\rho\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+L^{2}\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$
$\leq-\rho\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+L^{2}(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right].$
The first inequality arises from the application of Jensen’s inequality to two terms. The second inequality arises from Assumption 5.2. The last inequality arises form Lemma C.1.
Finally, bring $T_{1}$ and $T_{2}$ into Eq. (10) and subtract $\mathcal{L}(\overline{\theta}^{\text{PO}})$ from both sides, we can get that
$\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]$
$\leq(1-\rho\eta)\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]+\frac{L^{2}(R-1)^{2}G^{2}\eta^{3}}{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+LG^{2}\eta^{2}(1+\overline{\gamma})$
$+L\eta^{2}(1+\overline{\gamma})\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]+\mathbb{E}\left[|\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}-\frac{1}{2}\left\|g^{t}\right\|^{2}\right]\eta,$
where $\overline{\gamma}=\sum^{N}_{i=1}\alpha_{i}\gamma_{i}$ and $\sigma^{2}_{\gamma}=\sum^{N}_{i=1}\alpha_{i}(\overline{\gamma}-\gamma_{i})^{2}$.
Denote $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$ as $a_{t}$. Then we can obtain that
$a_{t+1}$ $\leq(1-\rho\eta)a_{t}+\frac{L^{2}(R-1)^{2}G^{2}\eta^{3}}{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$
$\quad+LG^{2}\eta^{2}(1+\overline{\gamma})+L\eta^{2}(1+\overline{\gamma})\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$
$\quad+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}-\frac{1}{2}\left\|g^{t}\right\|^{2}\right]\eta$
$=(1-\rho\eta)a_{t}+\underbrace{\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}-\frac{1}{2}\left\|g^{t}\right\|^{2}\right]}_{T_{4}}\eta$
$\quad+\underbrace{\frac{1}{2}\left(L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+2LG^{2}(1+\overline{\gamma})+2L(1+\overline{\gamma})\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]\right)}_{T_{3}}\eta^{2}.$ (11)
Denote the empirical result of the performative gradient of client $i$ in global iteration $t$ and local iteration $r$ as ^$\hat{\nabla}\mathcal{L}_{i}(\theta_{i}^{t})=\hat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\hat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$. According to Lemma C.2, we have
$\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$ $\leq\sum^{N}_{i=1}\alpha_{i}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}\right\|^{2}\right]$
$\leq(1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(LW_{1}\left(D,Q\right)_{\max}+\overline{\boldsymbol{\omega_{Q}}}+\overline{\boldsymbol{\omega_{D}}}\right)^{2}.$
where $\overline{\epsilon}=\sum^{N}_{i=1}\alpha_{i}\epsilon_{i},\sigma^{2}_{\epsilon}=\sum^{N}_{i=1}\alpha_{i}(\overline{\epsilon}-\epsilon_{i})^{2},$ and $\overline{\boldsymbol{\omega_{Q}}},\overline{\boldsymbol{\omega_{D}}}$ are the maximum of $\boldsymbol{\omega_{Q}}$ and $\boldsymbol{\omega_{D}}$ of all clients and all iterations. We can find the upper bound of $T_{3}$.
$T_{3}$ $\leq\frac{1}{2}\left(L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+2LG^{2}(1+\overline{\gamma})+2L(1+\overline{\gamma})\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]\right)$
$\leq\frac{1}{2}\left(L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)+L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(LW_{1}\left(D,Q\right)_{\max}+\overline{\boldsymbol{\omega_{Q}}}+\overline{\boldsymbol{\omega_{D}}}\right)^{2}+G^{2}\right).$
Let
$C=$ $\left(\frac{L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)}{2}+L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(LW_{1}\left(D,Q\right)_{\max}+\overline{\boldsymbol{\omega_{Q}}}+\overline{\boldsymbol{\omega_{D}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
$+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(LW_{1}\left(D,Q\right)_{\max}+\overline{\boldsymbol{\omega_{Q}}}+\overline{\boldsymbol{\omega_{D}}}\right)^{2}\right)\eta.$
We will obtain
$a_{t+1}\leq(1-\rho\eta)a_{t}+C$
and
$a_{t}\leq(1-\rho\eta)^{t}a_{0}+\frac{1-(1-\rho\eta)^{t}}{1-(1-\rho\eta)}C=(1-\rho\eta)^{t}a_{0}+\frac{1-(1-\rho\eta)^{t}}{\rho\eta}C\leq(1-\rho\eta)^{t}a_{0}+C,$
which means
$\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]\leq(1-\rho\eta)^{t}\mathbb{E}[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})]+C.$
C.2 Proof of Lemma 5.7
Proof.
$\frac{\widehat{df}}{d\theta}$ is the estimate of $\frac{df}{d\theta}$. To simplify the notation in this proof we denote ${f_{i}}_{z\sim D_{i}(\theta^{t}_{i})}(z)$ as ${f_{t}}$ and $\nabla{f_{i}}_{z\sim D_{i}(\theta^{t}_{i})}(z)$ as $\nabla{f_{t}}$. $\widehat{f}_{t}$ is the estimate of $f(\theta).$ Because $\widehat{f}_{t}$ is estimated by data sampled from $P_{i}(\theta^{t}_{i})=(1-\epsilon_{i})D_{i}(\theta^{t}_{i})+\epsilon_{i}Q_{i}$ and we estimate the mean of mixture Gaussian or the $p$ of Binomial distribution $X\sim(n,p)$ we have
$\widehat{f}_{r}$ $=\widehat{f_{i}}_{z\sim P_{i}(\theta^{t}_{i})}(z)$
$=(1-\epsilon_{i})\widehat{f_{i}}_{z\sim D_{i}(\theta^{t}_{i})}(z)+\epsilon_{i}\widehat{f_{i}}_{z\sim Q_{i}}(z)$
$=(1-\epsilon_{i})\left({f_{i}}_{z\sim D_{i}(\theta^{t}_{i})}(z)+\text{err}^{t}_{i}\right)+\epsilon_{i}\widehat{f_{i}}_{z\sim Q_{i}}(z),$
where $\text{err}^{t}_{i}$ is the error result in sampling of client $i$ on iteration $t$ and $[\text{err}^{t}_{i}]=0$ for all $t\in[0,T-1].$ Similarly we have denote $\theta^{t}_{i}$ as $\theta_{t}$ , $\theta^{t-1}_{i}$ as $\theta_{t-1}$ , $\text{err}^{t}_{i}$ as $\text{err}_{t}$ , and $\text{err}^{t-1}_{i}$ as $\text{err}_{t-1}.$
For each $1\leq j\leq d$ by an approximation of Taylor’s series (ignoring higher order terms), we obtain
$f_{t,j}-f_{t-1,j}=\nabla f_{t-1,j}^{\intercal}(\theta_{t}-\theta_{t-1})+\frac{1}{2}(\theta_{t}-\theta_{t-1})^{\intercal}\nabla^{2}f_{i}(\xi_{t-1,j})(\theta_{t}-\theta_{t-1}),$
where $\xi_{t-1,j}$ lies on the line segment joining $\theta_{t-1}$ and $\theta_{t}$.
Denote
$a_{t-k,j}=\frac{1}{2}(\theta_{t-k+1}-\theta_{t-k})^{\intercal}\nabla^{2}f_{i}(\xi_{t-k,j})(\theta_{t-k+1}-\theta_{t-k})~{}~{}\text{ and }~{}~{}a_{k}=\begin{bmatrix}a_{t-k,1}\\ \vdots\\ a_{t-k,d}\end{bmatrix}$
Then we have
$f_{t}-f_{t-1}=\frac{df}{d\theta}(\theta_{t}-\theta_{t-1})+a_{1}.$
By using the update rule of Algorithm 1 we obtain that
$\widehat{f}_{t}-\widehat{f}_{t-1}$ $=(1-\epsilon_{i})\left({f}_{t}-{f}_{t-1}\right)+\epsilon_{i}\left({f}_{z\sim Q_{i}}(z)-{f}_{z\sim Q_{i}}(z)\right)+(1-\epsilon_{i})(\text{err}_{t}-\text{err}_{t-1})$
$=(1-\epsilon_{i})\begin{bmatrix}f_{t,1}-f_{t-1,1}\\ \vdots\\ f_{t,d}-f_{t-1,d}\end{bmatrix}+(\text{err}_{t}-\text{err}_{t-1}).$
Then we can write $\Delta_{1}f$ in terms of $\frac{df}{d\theta}$.
$\Delta_{1}f=\widehat{f}_{t}-\widehat{f}_{t-1}=(1-\epsilon_{i})\left(\frac{df}{d\theta}(\theta_{t}-\theta_{t-1})+a_{1}+(\text{err}_{t}-\text{err}_{t-1})\right).$
Similarly, we can obtain
$\Delta_{k}f$ $=\widehat{f}_{t}-\widehat{f}_{t-k}=\sum^{k}_{q=1}\left(\widehat{f}_{t-q+1}-\widehat{f}_{t-q}\right)$
$=(1-\epsilon_{i})\frac{df}{d\theta}(\theta_{t}-\theta_{t-k})+(1-\epsilon_{i})\frac{df}{d\theta}\sum^{k}_{q=1}a_{q}+(1-\epsilon_{i})(\text{err}_{t}-\text{err}_{t-k}),$
where $1\leq k\leq H.$
Now we can get
[…]
Define matrices
[…]
Then we can find the difference between $\frac{\widehat{df}}{d\theta}$ and $\frac{df}{d\theta}:$
$\frac{\widehat{df}}{d\theta}-\frac{df}{d\theta}=\Delta f(\Delta\theta)^{\dagger}-\frac{df}{d\theta}=-\epsilon_{i}\frac{df}{d\theta}+(1-\epsilon_{i})A(\Delta\theta)^{\dagger}+(1-\epsilon_{i})E(\Delta\theta)^{\dagger}$
and we can obtain that
$\mathbb{E}\left[\left\|\frac{\widehat{df}}{d\theta}-\frac{df}{d\theta}\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|-\epsilon_{i}\frac{df}{d\theta}+(1-\epsilon_{i})A(\Delta\theta)^{\dagger}+(1-\epsilon_{i})E(\Delta\theta)^{\dagger}\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|-\epsilon_{i}\frac{df}{d\theta}+(1-\epsilon_{i})A(\Delta\theta)^{\dagger}\right\|^{2}\right]+(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|E(\Delta\theta)^{\dagger}\right\|^{2}\right]$ (12)
$\leq 2\epsilon_{i}^{2}\mathbb{E}\left[\left\|\frac{df}{d\theta}\right\|^{2}\right]+2(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|A(\Delta\theta)^{\dagger}\right\|^{2}\right]+(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|E(\Delta\theta)^{\dagger}\right\|^{2}\right]$ (13)
$\leq 2\epsilon_{i}^{2}\mathbb{E}\left[\left\|\frac{df}{d\theta}\right\|^{2}\right]+(1-\epsilon_{i})^{2}\left(2\mathbb{E}\left[\left\|A\right\|^{2}_{F}\right]+\mathbb{E}\left[\left\|E\right\|^{2}_{F}\right]\right)\left\|\Delta\theta^{\dagger}\right\|^{2}$ (14)
$\leq 2\epsilon_{i}^{2}\mathbb{E}\left[\left\|\frac{df}{d\theta}\right\|^{2}\right]+(1-\epsilon_{i})^{2}{2}\mathbb{E}\left[\sum_{k=1}^{H}\left\|\theta_{t}-\theta_{t-k}\right\|^{4}\left(\sum_{j=1}^{d}\left\|\nabla^{2}f_{i}(\xi_{t-k,j})\right\|^{2}\right)\right]\left\|\Delta\theta^{\dagger}\right\|^{2}$
$+(1-\epsilon_{i})^{2}\mathbb{E}\left[\sum_{k=1}^{H}\left\|\text{err}_{t}-\text{err}_{t-k}\right\|^{2}\right]\left\|\Delta\theta^{\dagger}\right\|^{2}$
$\leq 2\epsilon_{i}^{2}\left\|\frac{df}{d\theta}\right\|^{2}+\left(\frac{(1-\epsilon_{i})^{2}M^{2}\eta^{4}G^{4}H^{6}}{2}+(1-\epsilon_{i})^{2}\mathbb{E}\left[\sum_{k=1}^{H}\left\|\text{err}_{t}-\text{err}_{t-k}\right\|^{2}\right]\right)\left\|\Delta\theta^{\dagger}\right\|^{2}$
$\leq 2\epsilon_{i}^{2}\left\|\frac{df}{d\theta}\right\|^{2}+\left(\frac{(1-\epsilon_{i})^{2}M^{2}\eta^{4}G^{4}H^{6}}{2}+(1-\epsilon_{i})^{2}\mathbb{E}\left[\sum_{k=1}^{H}\left(\left\|\text{err}_{t}\right\|^{2}+\left\|\text{err}_{t-k}\right\|^{2}\right)\right]\right)\left\|\Delta\theta^{\dagger}\right\|^{2}$ (15)
$\leq 2\epsilon_{i}^{2}F^{2}+(1-\epsilon_{i})^{2}\left(\frac{M^{2}\eta^{4}G^{4}H^{6}}{2}+\frac{2Hd\left\|\text{err}\right\|^{2}_{i}}{\eta^{2}}\right)\lambda^{-2}_{i,\min}.$ (16)
Where .$\left\|.\right\|_{F}$ denotes the Frobenius norm. (12) and (15) arise from the property that $\mathbb{E}\left[\text{err}_{t}\right]=0$ for all $t\in[0,T]$ and $\text{err}_{t}$ is independent to $a_{k}$ and $\frac{df}{d\theta}$. (13) arises from Jensen’s inequality of two terms. (14) arises form the properties of matrix norm that if $A$ and $B$ are two matrix $\left\|AB\right\|\leq\left\|A\right\|\left\|B\right\|$ and $\left\|A\right\|\leq\left\|A\right\|_{F}.$ (16) arises form $\left\|\Delta\theta^{\dagger}\right\|\leq\left\|\Delta\theta\right\|^{-1}\leq\lambda^{-1}_{i,min}$, where $\lambda_{i,min}$ is the smallest singular values of $\Delta\theta$ of client $i$ for all $t\in[0,T]$ and $\left\|\text{err}\right\|^{2}_{i}$ is the maximum of $\left\|\text{err}_{t}\right\|^{2}$ for all $t\in[0,T]$. Because $\left\|\text{err}_{t}\right\|^{2}$ depends on the data distribution, the estimation methods, the number of samples $n_{i}$, as well as the probability $\varphi_{i}$ of $\mathbf{Pr}(\left\|\text{err}_{t}\right\|\leq 1-\varphi_{i})$, we use $f_{\varphi}(\varphi_{i};n_{i})$ to represent this term.
Finally we obtrain
$\mathbb{E}\left[\left\|\frac{\widehat{df}}{d\theta}-\frac{df}{d\theta}\right\|^{2}\right]\leq 2\epsilon_{i}^{2}F^{2}+(1-\epsilon_{i})^{2}\left(\frac{M^{2}\eta^{4}G^{4}H^{6}}{2}+\frac{2Hdf_{\varphi}(\varphi_{i};n_{i})}{\eta^{2}}\right)\lambda^{-2}_{i,\min}.$
C.3 Proof of Lemma 5.9
Proof.
First we start to find the upper bound of $\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}z\sim D_{i}(\theta_{i}^{t})}(z,\theta_{i}^{t})\right\|^{2}\right].$
By Eq. 3 we obtain that
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}z\sim D_{i}(\theta_{i}^{t})}(z,\theta_{i}^{t})\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|\ell(Z_{i};\theta_{i}^{t}){\frac{df_{i,t}}{d\theta}}^{\intercal}\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}-\widehat{\ell}(Z_{i};\theta_{i}^{t}){\frac{\widehat{df_{i,t}}}{d\theta}}^{\intercal}\frac{\partial\log p\left(Z_{i};\widehat{f}_{i}(\theta_{i}^{t})\right)}{\partial\widehat{f}_{i}(\theta_{i}^{t})}\right\|^{2}\right].$
When the local client’s sample size $n_{i}\rightarrow\infty$, we have
$\mathbb{E}\left[\left\|\ell(Z_{i};\theta)-\widehat{\ell}(Z_{i};\theta)\right\|^{2}\right]\rightarrow 0$
$\mathbb{E}\left[\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}-\frac{\partial\log p\left(Z_{i};\widehat{f}_{i}(\theta)\right)}{\partial\widehat{f}_{i}(\theta)}\right\|^{2}\right]\rightarrow 0$
Thereby, $\mathbb{E}\left[\left\|\ell(Z_{i};\theta_{i}^{t}){\frac{df_{i,t}}{d\theta}}^{\intercal}\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}-\widehat{\ell}(Z_{i};\theta_{i}^{t}){\frac{\widehat{df_{i,t}}}{d\theta}}^{\intercal}\frac{\partial\log p\left(Z_{i};\widehat{f}_{i}(\theta_{i}^{t})\right)}{\partial\widehat{f}_{i}(\theta_{i}^{t})}\right\|^{2}\right]$ mainly includes two parts.
The first part is the following:
$\mathbb{E}\left[\left\|\ell(Z_{i};\theta_{i}^{t})\right\|^{2}\left\|\frac{df_{i,t}}{d\theta}-\frac{\widehat{df_{i,t}}}{d\theta}\right\|^{2}\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}\right]$
$\leq\ell^{2}_{\max}\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}_{\max}\mathbb{E}\left[\left\|\frac{df_{i,t}}{d\theta}-\frac{\widehat{df_{i,t}}}{d\theta}\right\|^{2}\right]$ (17)
$=\mathcal{O}\left(\ell^{2}_{\max}d\left(\boldsymbol{\omega_{F}}\right)\right).$ (18)
Because $\Theta$ is a closed and bounded set and a continuous function has a maximum value on $\Theta$, there is an upper bound $\ell^{2}_{\max}$ of $\left\|\ell(Z_{i};\theta)\right\|^{2}$ and an upper bound $\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}_{\max}$ of $\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}$. (18) arises from the upper bound of $\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}$ is $\mathcal{O}(d)$ and it is related to the covariance $\Sigma$ of the data for the Gaussian distribution.
The second part will decrease to zero as $n_{i}\rightarrow 0.$ The upper bound of this part is $\mathcal{O}\left(\ell^{2}_{\max}dF^{2}f_{\varphi}(\varphi_{i};n_{i})\right)$ by using Lemma 12 in Izzo et al., (2021). Then we can obtain
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}z\sim D_{i}(\theta_{i}^{t})}(z,\theta_{i}^{t})\right\|^{2}\right]\leq\mathcal{O}\left(\ell^{2}_{\max}d\left(\left(\boldsymbol{\omega_{F}}\right)+F^{2}f_{\varphi}(\varphi_{i};n_{i})\right)\right).$
Next we find the upper bound of $\mathbb{E}\left[\left\|{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right].$
$\mathbb{E}\left[\left\|{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|{\ell}_{z\sim Q_{i}}(Z_{i};\theta_{i}^{t}){\frac{\widehat{df_{i,t}}}{d\theta}}^{\intercal}\frac{\partial\log p\left(Z_{i};\widehat{f}_{i}(\theta_{i}^{t})\right)}{\partial\widehat{f}_{i}(\theta_{i}^{t})}\right\|^{2}\right]$ (19)
$\leq\mathcal{O}\left(\ell^{2}_{\max}d\mathbb{E}\left[\left\|\frac{\widehat{df_{i,t}}}{d\theta}\right\|^{2}\right]\right)$ (20)
$\leq\mathcal{O}\left(2\ell^{2}_{\max}d\left(\mathbb{E}\left[\left\|\frac{{df_{i,t}}}{d\theta}\right\|^{2}\right]+\mathbb{E}\left[\left\|\frac{df_{i,t}}{d\theta}-\frac{\widehat{df_{i,t}}}{d\theta}\right\|^{2}\right]\right)\right)$ (21)
$=\mathcal{O}\left(2\ell^{2}_{\max}d\left(\left(\boldsymbol{\omega_{F}}\right)+F^{2}\right)\right).$ (22)
(19) arises from Eq. (3). (20) arises from Lemma 10 of Izzo et al., (2021) and Lemma 5.6. (21) arises from Jensen’s inequality of two terms. (22) arises from Lemma 5.7. ∎
C.4 Additional Lemmas and their Proofs
Lemma C.1.
Let $\overline{\gamma}=\sum^{N}_{i=1}\alpha_{i}\gamma_{i}$ and $\sigma^{2}_{\gamma}$ be the variance of $\gamma_{i}$ of all clients. Then $\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$ is upper bounded by $(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$.
Proof of Lemma C.1.
$\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$ $=\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\mathbb{E}\left[\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$
$=\eta^{2}\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\mathbb{E}\left[\left\|\sum_{i=1}^{N}\alpha_{i}\sum_{j=t-b}^{t}\nabla\mathcal{L}_{i}({\theta}_{i}^{j})-\sum_{j=t-b}^{t}\nabla\mathcal{L}_{i}({\theta}_{i}^{j})\right\|^{2}\right]$
$\leq\eta^{2}\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\mathbb{E}\left[\left\|\sum_{j=t-b}^{t}\nabla\mathcal{L}_{i}({\theta}_{i}^{t})\right\|^{2}\right]$
$\leq c^{2}G^{2}\eta^{2}\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}$
$\leq(R-1)^{2}G^{2}\eta^{2}\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}.$
$c=t\mod R$. The first inequality rises from $\mathbb{E}\left[\|\mathbb{E}[X]-X\|^{2}\right]\leq\mathbb{E}[X^{2}]$. The second inequality arises from Lemma 5.6 and Jensen’s inequality.
Denote $\sum^{N}_{i=1}\alpha_{i}\gamma_{i}$ as $\overline{\gamma}$ and variance of $\gamma_{i}$ as $\sigma^{2}_{\gamma}$. Because
$\sum^{N}_{i=1}\alpha_{i}(1+\gamma_{i})^{2}$ $=\mathbb{E}\left[(1+\gamma_{i})^{2}\right]$
$=\left(\mathbb{E}\left[1+\gamma_{i}\right]\right)^{2}+\mathbb{V}\left[1+\gamma_{i}\right]$
$=\left(\mathbb{E}\left[1+\gamma_{i}\right]\right)^{2}+\mathbb{V}\left[\gamma_{i}\right]=(1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}$
we will finally obtain
$\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]\leq(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right).$
Lemma C.2.
If there is contaminated data with distribution $Q_{i}$ and fraction $\epsilon_{i}$ , with Assumption 5.2, we can obtain that
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}\right\|^{2}\right]\leq(1-\epsilon_{i})(\boldsymbol{\omega_{D}})^{2}+\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})+(\boldsymbol{\omega_{D}})\right)^{2}.$
The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemmas 5.8 and 5.9. $W_{1}\left(D,Q\right)_{\max}$ is the largest $W_{1}$ distance between $D_{i}(\theta)$ and $Q_{i}$ for all clients $i\in\mathcal{V}$ among all iteration $t\in[0,T-1]$.
Proof of Lemma C.2.
From Eq. (3) we can obtain
$\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}=$ $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$
$=$ $(1-\epsilon_{i})\left(\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right)$
$+\epsilon_{i}\left(\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i,1}}(Q_{i},\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right)$
$+\epsilon_{i}\left({\nabla}\mathcal{L}_{{i,1}z\sim Q_{i}}(z,\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(Q_{i},\theta_{i}^{t})+{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right)$
$=$ $(1-\epsilon_{i})\left(\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right)$
$+\epsilon_{i}\left(\nabla\mathcal{L}_{i}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i}z\sim Q_{i}}(z,\theta_{i}^{t})\right)-\epsilon_{i}\left(\widehat{\nabla}\mathcal{L}_{{i,2}z\sim Q_{i}}(z,\theta_{i}^{t})\right).$
The last equality arises form ${\nabla}\mathcal{L}_{{i,1}}(Q_{i},\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}z\sim}(Q_{i},\theta_{i}^{t})=0\text{ and }{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})=0$. Because $\mathbb{E}\left[\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})]\right]=0$ and Assumption 5.2 of $L$-smoothness.
The proof of the Lemma 12 of Izzo et al., (2021) shows that
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]=\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right).$
Finally, we can obtain that
$\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}\right\|^{2}\right]$
$=(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]+(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]$
$+(1-\epsilon_{i})^{2}\epsilon_{i}\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t}),\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\rangle\right]$
$+\epsilon_{i}^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]+\epsilon_{i}^{2}\mathbb{E}\left[\left\|\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]$
$+2(1-\epsilon_{i})\epsilon_{i}\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t}),\nabla\mathcal{L}_{i}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i}}(Q_{i},\theta_{i}^{t})\right\rangle\right]$
$-2(1-\epsilon_{i})\epsilon_{i}\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t}),\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\rangle\right]$
$-\epsilon^{2}_{i}\mathbb{E}\left[\left\langle\nabla\mathcal{L}_{i}(\theta_{i}^{t})-{\nabla}\mathcal{L}_{{i}}(Q_{i},\theta_{i}^{t}),\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\rangle\right]$
$\leq(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]+(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]$
$+(1-\epsilon_{i})^{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|\right]$
$+\epsilon_{i}^{2}L^{2}\left\|D_{i}(\theta_{i}^{t})-Q_{i}\right\|^{2}+\epsilon_{i}^{2}\mathbb{E}\left[\left\|\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|^{2}\right]$
$+2(1-\epsilon_{i})\epsilon_{i}L\left\|D_{i}(\theta_{i}^{t})-Q_{i}\right\|\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|\right]$
$+2(1-\epsilon_{i})\epsilon_{i}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|\left\|\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|\right]$
$+\epsilon^{2}_{i}L\left\|D_{i}(\theta_{i}^{t})-Q_{i}\right\|\mathbb{E}\left[\left\|\widehat{\nabla}\mathcal{L}_{{i,2}}(Q_{i},\theta_{i}^{t})\right\|\right]$
$\leq(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}+(1-\epsilon_{i})^{2}\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right)(\boldsymbol{\omega_{D}})+\epsilon_{i}^{2}\left(L^{2}W^{2}_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})^{2}+W_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{Q}})\right)$
$+2(1-\epsilon_{i})\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{D}})+(\boldsymbol{\omega_{D}})(\boldsymbol{\omega_{Q}})\right)$
The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemma 5.9. $W_{1}\left(D,Q\right)_{\max}$ is the largest $W_{1}$ distance between $D_{i}(\theta)$ and $Q_{i}$ for all clients $i\in\mathcal{V}$ among all iteration $t\in[0,T-1]$. To simplify the upper bound first we use the fact that
$(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}+(1-\epsilon_{i})^{2}\mathcal{O}(f_{\varphi}(\varphi_{i};n_{i}))(\boldsymbol{\omega_{D}})\approx(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}$ (23)
because the estimation error of $\nabla\mathcal{L}_{i,1}$ is much smaller than $\nabla\mathcal{L}_{i,2}$ by using Lemmas 5.8 and 5.9. Apply the fact that $\epsilon_{i}\in[0,1]$ and we can obtain
$(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}+(1-\epsilon_{i})^{2}\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right)(\boldsymbol{\omega_{D}})+\epsilon_{i}^{2}\left(L^{2}W^{2}_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})^{2}+W_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{Q}})\right)$
$+2(1-\epsilon_{i})\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{D}})+(\boldsymbol{\omega_{D}})(\boldsymbol{\omega_{Q}})\right)$
$\leq(1-\epsilon_{i})(\boldsymbol{\omega_{D}})^{2}+\epsilon_{i}\left(L^{2}W^{2}_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})^{2}+W_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{Q}})+LW_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{D}})+(\boldsymbol{\omega_{D}})(\boldsymbol{\omega_{Q}})\right)$
$\leq(1-\epsilon_{i})(\boldsymbol{\omega_{D}})^{2}+\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})+(\boldsymbol{\omega_{D}})\right)^{2}.$
Lemma C.3 (Sensitivity of contribution Dynamic).
For any client $i\in\mathcal{V}$, if Assumption 5.1and 5.3 hold, then with $D_{i}(\theta)=\nu_{i,1}(\theta)D_{i,1}(\theta)+\left(1-\nu_{i,1}(\theta)\right)D_{i,2}(\theta)$ for all $\theta,\theta^{\prime}\in\Theta$, we have:
$W_{1}(D_{i}(\theta),D_{i}(\theta^{\prime}))\leq(\gamma_{i}+W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))F)\|\theta-\theta^{\prime}\|.$
Proof of Lemma C.3.
In the proof, we consider two groups with fractions $\nu_{i,1}$ and $\nu_{i,2}=1-\nu_{i,1}$. When $\nu_{i,1}(\theta)\geq\nu_{i,1}(\theta^{\prime})$, we obtain:
$W_{1}(D_{i}(\theta),D_{i}(\theta^{\prime}))$ $\leq\nu_{i,1}(\theta^{\prime})W_{1}(D_{i,1}(\theta),D_{i,1}(\theta^{\prime}))+\left(1-\nu_{i,1}(\theta)\right)W_{1}(D_{i,2}(\theta),D_{i,2}(\theta^{\prime}))$
$\quad+(\nu_{i,1}(\theta)-\nu_{i,1}(\theta^{\prime}))\left(W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))+W_{1}\left(D_{i,2}(\theta),D_{i,2}(\theta^{\prime})\right)\right)$
$\leq\nu_{i,1}(\theta^{\prime})\gamma_{i}\|\theta-\theta^{\prime}\|+(1-\nu_{i,1}(\theta^{\prime}))\gamma_{i}\|\theta-\theta^{\prime}\|$
$\quad+(\nu_{i,1}(\theta)-\nu_{i,1}(\theta^{\prime}))W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))$
$=\gamma_{i}\|\theta-\theta^{\prime}\|+(\nu_{i,1}(\theta)-\nu_{i,1}(\theta^{\prime}))W_{1}(D_{i,1}(\theta),D_{i,2}(\theta)).$
The first inequality arises from the triangle inequality for the Wasserstein-1 distance. The second inequality arises from Assumption 5.1. Similarly, when $\nu_{i,1}(\theta)\leq\nu_{i,1}(\theta^{\prime})$, we obtain:
$W_{1}(D_{i}(\theta),D_{i}(\theta^{\prime}))$ $\leq\gamma_{i}\|\theta-\theta^{\prime}\|+(\nu_{i,1}(\theta^{\prime})-\nu_{i,1}(\theta))W_{1}(D_{i,1}(\theta),D_{i,2}(\theta)).$
The estimate function $f_{i}(\theta)=\nu_{i,1}$ here. By combining these results and Lemma 5.6 that $\left\|\frac{df_{i}(\theta)}{d\theta}\right\|$ has the upper bound $F$, we finally obtain:
$W_{1}(D_{i}(\theta),D_{i}(\theta^{\prime}))$ $\leq\gamma_{i}\|\theta-\theta^{\prime}\|+\|\nu_{i,1}(\theta^{\prime})-\nu_{i,1}(\theta)\|\,W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))$
$\leq\left(\gamma_{i}+W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))F\right)\|\theta-\theta^{\prime}\|.$
Appendix D Additional Experimental Results
D.1 Additional Experimental Results of Convergence to Performative Optimal Solution
Fig. 3 shows the results of case study house pricing regression, where an FL system is trained with 10 heterogeneous clients. Although PFL can stabilize the training process when data changes based on the model, it converges to an undesirable stable solution. In contrast, our algorithm converges to a stable solution with a much lower loss. In Fig. 3(a), the results show that both algorithms exhibit faster convergence rates with larger enrollment fraction. However, significant oscillations occur in PFL when the enrollment fraction is low, while such oscillations are much weaker in ProFL.
Fig. 3(b) shows the impact of varying degrees of heterogeneity, where $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$ with $\alpha\in\{0.1,0.3,1\}$. The results show that, despite the varying levels of heterogeneity, our algorithm maintains the capacity to converge to ${\theta}^{\text{PO}}$.
(a) Enrollment Fraction
(b) Different Heterogeneity
Figure 3: Linear Regression Results with Different Enrollment Fractions and Heterogeneity
(a) Sample Size
(b) Local Rounds $R$
Figure 4: Linear Regression Results with Different Sample Size and Local Rounds
In Fig. 4(a), the results show that both algorithms exhibit faster convergence rates with larger sample sizes. With a larger sample size, the results of PFL are closer to the performative stable point and exhibit fewer oscillations. However, PFL cannot converge to the performative optimal point, no matter how many samples are used. Fig. 4(b) demonstrates that a larger value of $R$ corresponds to faster convergence, reducing the communication time required in the FL system to reach convergence.
Fig. 5(a) shows the results of the case study pricing with dynamic contribution with 10 heterogeneous clients. The results indicate that, unlike PFL, which converges to $\theta^{\text{PS}}$, our algorithm converges to a solution, which is close to the performative optimal point. Fig. 5(b) shows the results of the case study regression with dynamic contribution with 10 heterogeneous clients. ProFL has faster convergence speed and less oscillation, although there is a gap between the results of both algorithms and the performative optimal point. The gap is a result of the non-linear $f_{i}(.)$, which has been discussed in more detail in the main paper.
(a) Pricing
(b) Regression
Figure 5: Contribution Dynamics
D.2 Performance when ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
In this section, we introduce a special case where $\theta^{\text{PO}}=\theta^{\text{PS}}$. This special case can arise in both strategic and dynamic contribution settings. We present the results in Fig. 6, demonstrating that when $\theta^{\text{PO}}=\theta^{\text{PS}}$, both algorithms converge to the same point.
(a) Strategic Dynamic
(b) Contribution Dynamic
Figure 6: Dynamics with ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
D.2.1 Performance without Model-dependent Shifts.
We also evaluate the algorithm when the data distribution does not change based on the model. In this case study, we use the Adult dataset for a binary classification task. Specifically, there are 10 homogeneous clients with $\gamma_{i,0}=\gamma_{i,1}=0$. We consider four features from the Adult dataset. When the data remains constant as the model is updated, both PFL and our algorithm reduce to FedAvg, and they converge to the same solution. This is illustrated in Fig. 6(a).
D.2.2 Performance when $\theta^{\text{PO}}=\theta^{\text{PS}}$ with Dynamic Contribution
Consider case study pricing with dynamic contribution with $K=2$ groups with different fixed distributions but a different one-dimensional fraction where $f_{i,1}(\theta)=0.5+0.5\theta$ and $f_{i,2}(\theta)=1-f_{i,1}(\theta)$. Under this special case, the performative stable and optimal points are identical, i.e., $\theta^{\text{PO}}=\theta^{\text{PS}}$. We can calculate $\theta^{\text{PO}}=\frac{a_{1}+a_{2}}{2(a_{1}-a_{2})}$ and $\theta^{\text{PS}}=\frac{a_{1}+a_{2}}{a_{1}-a_{2}}$. When $a_{1}+a_{2}=0$, $\theta^{\text{PO}}=\theta^{\text{PS}}$. In this case, we set $a_{1}=-a_{2}=0.5$. Fig. 6(b) illustrates the results for this specific case study, showing that both our algorithm and PFL converge to the same solution. However, our method exhibits faster convergence compared to PFL.
D.2.3 Figures of Table 2
The figures illustrating the experimental results from Table 2 in the main paper are shown in Fig. 7.
(a) Same Distributions
(b) Different Distributions
(c) Credit Dataset
(d) Adult Dataset
Figure 7: Binary Classification on Synthetic and Realistic Dataset
Appendix E Training details
The experimental setup and parameter selections follow the methodology established in Izzo et al., (2021); Jin et al., (2024), ensuring consistency and comparability with existing results.
E.1 Realistic Data
  • Adult data (Becker and Kohavi,, 1996) where the goal is to predict whether a person’s annual income exceeds \$50K based on 14 features (e.g., age, years of education).
  • Give Me Some Credit data (Credit Fusion,, 2011) which has 11 features (e.g., monthly income, debt ratio) and can be utilized to predict whether a person has experienced a 90-day past-due delinquency or worse.
E.2 Loss Functions and Ridge Penalty
We use ridge-regularized cross-entropy loss for all binary classification cases, and the ridge penalty is $0.01$. We use ridge-regularized squared loss for all regression cases, and the ridge penalty is $\frac{10}{3}$.
E.3 Random Seeds
In our experiments, which show the results of 10 trials, the random seeds used range from 0 to 9.
E.4 Parameters and Distributions
The experimental parameters are listed in Table 3, and the distribution $D_{i}(\cdot)$ for each experiment is detailed in Table 4. The corresponding case studies for the experiments are shown in Table 5.
Table 3: Parameters in the Experiments
Figure/Table $\eta$ $R$ H $n_{i}$ $\epsilon_{i}$
Table 1: PG 0.01 -- 1 500 0
Table 1: ProFL 0.01 5 1 500 0
Table 2: Same $D_{i}$ 0.03 10 500 500 0
Table 2: Different $D_{i}$ 0.03 500 1 500 0
Table 2: Credit 0.01 5 1 5812 0
Table 2: Adult 0.03 5 1 2368 0
Figure 1a 0.001 5 25 500 0, 0.1,0.2,0.4
Figure 1b 0.0001 3 10 60 0
Figure 1c 0.001 5 100 1000 0
Figure 2a 0.001 5 20 50,500,5000 0
Figure 2b 0.01,0.03,0.001 5 20 2500 0
Figure 2c 0.001 5 4,20,100 500 0
Figure 3a PFL:0.005 ProFL:0.001 5 0.2, 0.5, 1 500 0
Figure 3b PFL:0.005 ProFL:0.001 5 1 500 0
Figure 3a 0.03 5 1 2368 0
Figure 3b 0.03 5 1 2368 0
Figure 4a PFL:0.005 ProFL:0.001 5 1 5, 50, 500 0
Figure 4b PFL:0.005 ProFL:0.001 5,20,50 1 500 0
Figure 5a 0.03 5 1 500 0
Figure 5b 0.001 5 1 500 0
Figure 6a 0.03 5 1 500 0
Figure 6b 0.03 5 1 500 0
Table 4: Data Distributions in the Experiments
Figure/Table Distribution
Table 1: PG and ProFL $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0,0.25,0.5\}$
Table 2: Same $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$ and $X_{i}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$.
Table 2: Different $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$, $X_{1}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$
and $X_{2}\sim\mathcal{N}(-0.8-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$
Table 2: Credit $\gamma_{i,1}=3$ and $\gamma_{i,0}=0.$ Give Me Some Credit datasetCredit Fusion, (2011)
Table 2: Adult $\gamma_{i,1}\in[2.8,3.2]$ and $\gamma_{i,0}\in[0,0.004].$ Adult dataset (Becker and Kohavi,, 1996)
Figure 1a $\mu_{0}\in[6,7]$, $\sigma=1$, $\gamma_{i}\in[1,3]$
Figure 1b,1c,2a-c $\gamma_{i}\in[0.09,0.11]$
Figure 3a $\gamma_{i}\in[1.5,1.8]$
Figure 3b $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0.1,0.3,1\}$
Figure 4a $\gamma_{i}\in[1.5,1.8]$
Figure 4b $\gamma_{i}\in[1.5,1.8]$
Figure 5a $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$ $f(\theta)=0.5+0.5\theta$
Figure 5b $y=f_{i}(\theta)x+3(1-f_{i}(\theta))x+\varepsilon,\varepsilon\sim\mathcal{N}(0,4)$
Figure 6a $\gamma_{i}=0.$ 10 homogeneous clients with Adult dataset
Figure 6b $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$
Table 5: Type of Case Study and Corresponding Figures or Tables
Figure/Table Type of Case Study
Table 1: house pricing regression
Table 2: binary strategic classification
Figure 1a pricing with dynamic demands
Figure 1b,1c,2a-c pricing with dynamic contribution
Figure 3a,3b,4a,4b house pricing regression
Figure 5a pricing with dynamic contribution
Figure 5b regression with dynamic contribution
Figure 6a binary strategic classification
Figure 6b pricing with dynamic contribution

### v2

@@
 Abstract
-Performative prediction (PP) is a framework that captures distribution shifts that occur during the training of machine learning models due to their deployment.
-As the trained model is used, its generated data could cause the model to evolve, leading to deviations from the original data distribution.
-The impact of such model-induced distribution shifts in the federated learning (FL) setup remains unexplored, despite being increasingly likely to transpire in real-life use cases.
-Although Jin et al., (2024) recently extended PP to FL in a straightforward manner, the resulting model only converges to a performative stable point, which may be far from optimal.
-The methods in Izzo et al., (2021); Miller et al., (2021) can find a performative optimal point in centralized settings, but they require the performative risk to be convex and the training data to be noiseless—assumptions often violated in realistic FL systems.
-This paper overcomes all of these shortcomings and proposes Performative robust optimal Federated Learning (ProFL), an algorithm that finds performative optimal points in FL from noisy and contaminated data.
+Performative prediction is a framework that captures distribution shifts that occur during the training of machine learning models due to their deployment.
+As the trained model is used, data generation causes the model to evolve, leading to deviations from the original data distribution.
+The impact of such model-induced distribution shifts in federated learning is increasingly likely to transpire in real-life use cases.
+A recently proposed approach extends performative prediction to federated learning with the resulting model converging to a performative stable point, which may be far from the performative optimal point.
+Earlier research in centralized settings has shown that the performative optimal point can be achieved under model-induced distribution shifts, but these approaches require the performative risk to be convex and the training data to be noiseless—assumptions often violated in realistic federated learning systems.
+This paper overcomes all of these shortcomings and proposes Performative Robust Optimal Federated Learning, an algorithm that finds performative optimal points in federated learning from noisy and contaminated data.
 We present the convergence analysis under the Polyak-Lojasiewicz condition, which applies to non-convex objectives.
-Extensive experiments on multiple datasets validate our proposed algorithms’ efficiency.
-Keywords Federated learning  $\cdot$ Performative prediction $\cdot$ Robust
+Extensive experiments on multiple datasets demonstrate the advantage of Robust Optimal Federated Learning over the state-of-the-art.
+Keywords Performative prediction, Federated learning, Robust optimization, Distribution shift
 1 Introduction
-Federated learning (FL) is a distributed learning paradigm that enables efficient learning from heterogeneous clients.
-Conventional FL (e.g., collaborative model training) often relies on the unrealistic assumption that each client’s data distribution remains static throughout the training process or between the training and testing phases.
-Recent work addresses this weakness by training models on time-evolving data (Guo et al.,, 2023; Ma et al.,, 2022) or training models on static data but deploying them on different data distributions (Nguyen et al.,, 2022; Jiang and Lin,, 2022).
-However, these works assume that the changes in data distribution are exogenous (i.e., are independent of the FL system).
-Examples of such exogenous distribution shifts include variations in data quality, lighting conditions, dynamic climate data, etc.
-In many real-world applications, data often undergoes performative endogenous shifts induced by the model’s deployment.
-For instance, content recommended by machine learning algorithms on digital platforms can reshape consumer preferences, steering them toward modes of consumption that are easier to predict and monetize (Dean and Morgenstern,, 2022).
-Similarly, navigation systems can influence traffic patterns, while individuals may adapt their behavior to game decision-making systems, such as those used for loan approvals or job applications (Hardt et al.,, 2016; Zhang et al.,, 2022).
-These types of shifts have been studied in centralized settings under the framework of performative prediction (PP), which seeks performative stable (Perdomo et al.,, 2020) or optimal (Izzo et al.,, 2021; Miller et al.,, 2021) solutions.
-Addressing endogenous shifts in federated learning (FL) is not a simple extension of these centralized efforts and remains a significant challenge.
-Failing to do so can lead to suboptimal performance or even non-convergence.
+Federated learning (FL) is a distributed learning paradigm that enables efficient training across decentralized clients while preserving data privacy.
+This privacy preservation is achieved by keeping the raw training data on each client and only exchanging model updates with the central server.
+Such a framework has attracted significant attention in privacy-sensitive applications such as the Internet of Things, healthcare, and financial services.
+Conventional FL (e.g., collaborative model training via FedAvg) typically assumes that each client’s data distribution remains static throughout the training process or between the training and deployment phases.
+However, this assumption often fails in real-world applications, where user behavior and data characteristics are continuously evolving.
+Although some previous work considers “dynamics” in FL, they largely focus on changes across clients (i.e., statistical heterogeneity) or time-varying participation rates rizk2020dynamic; park2021tackling; wang2022unified; zhu2021diurnal, rather than on distributional shifts within each client’s local data over time.
+Data shifts are common in FL deployment.
+For instance, periodic deployment and retraining often lead to a mismatch between the data used for training and the data encountered in the future.
+Recent work begins to address such settings by training on time-evolving data guo2021towards; ma2022continual, or by training on one distribution and deploying on another nguyen2022fedsr; jiang2022test.
+However, these works primarily consider exogenous shifts—distribution changes that occur independently of the model, such as variations in sensor quality, lighting, or environmental conditions.
+In contrast, many practical FL applications involve endogenous distribution shifts, where the deployed model itself induces changes in the data distribution.
+For example, recommendation systems shape user preferences dean2022, navigation apps influence traffic flow, and large language models affect language usageanderson2025modelmisalignmentlanguagechange.
+Users may also adapt their behavior to exploit automated decision systems, such as those in hiring or credit scoring Hardt2016a; zhang2022.
+These model-driven shifts fall under the framework of performative prediction (PP), which seeks performative stable perdomo2020performative or optimal pmlr-v139-izzo21a; miller2021outside solutions.
+Extending performative prediction to the federated setting is highly nontrivial.
+While centralized approaches can rely on access to a global dataset and full observability of distribution shifts, FL systems must contend with decentralized and partially observable dynamics.
+Moreover, the interaction between distributed optimization and performative feedback can lead to suboptimal convergence or even divergence if not properly addressed.
+Therefore, a principled treatment of performative shifts in FL is both necessary and fundamentally different from centralized solutions.
 Another challenge in practical FL systems, beyond model-induced endogenous shifts, is data contamination.
@@
 Unlike centralized systems, in FL, such data cannot be removed before training, as it remains with the clients.
-Developing FL systems that are robust to endogenous shifts and data contamination remains a significant challenge.
-The only work to date that addresses the endogenous shifts in FL is Jin et al., (2024).
-Their approach assumes noiseless client data and only finds a performative stable point, which may be suboptimal.
-While Izzo et al., (2021); Miller et al., (2021) achieve a performative optimal point, their methods are centralized, assume noiseless data, and require convex performative risk.
-As we will show, the method that directly extends Izzo et al., (2021) to federated settings is highly vulnerable to data contamination and performs poorly when distribution shifts are complex.
+Developing FL systems that are simultaneously robust to endogenous shifts and data contamination remains a significant challenge.
+1.1 Related Work
+Performative prediction perdomo2020performative was first introduced in 2020 as an optimization framework to deal with endogenous data distribution shifts, where the model deployed to the environment affected the subsequent data, resulting in the collected data changing as the deployed model changed.
+Common applications of performative prediction included strategic classification Hardt2016a; xie2024nonlinear and predictive policing ensign2018runaway.
+Early work focused primarily on finding performative stable (PS) points.
+For example, perdomo2020performative proposed Repeated Risk Minimization to find the PS point and bounded the distance between the PS point and the performative optimal (PO) point.
+mendler2022anticipating designed the first algorithm to find the performative stable point under the online learning setting.mendler2020stochastic proved convergence of greedy and lazy deployment under smoothness and strong convexity.
+Subsequent work moved toward directly identifying PO point.
+miller2021outside optimized performative risk directly, although the scope of the distribution map was restricted, and pmlr-v139-izzo21a designed an algorithm, performative gradient descent, for finding the performative optimal solution.
+Zhao2022OptimizingTP relaxed the assumption of strong convexity to the weakly-strongly convex case and proved the convergence of the deployment.
+somerstep2024learning; bracale2024learning focused on learning the distribution map in performative prediction and proposed a framework named reverse causal PP.
+In addition, PP was also related to reinforcement learning pmlr-v162-zheng22c and bandit problems chen2023performative.
+To the best of our knowledge, jin2023performative is the only work discussing performativity under the distributed setting, where they generalized Fedavg to Performative Federated Learning (PFL) and proved the uniqueness of the performative stable solution found by the algorithm.
+They also quantified the distance between the performative stable solution and the performative optimal solution.
+While the work is a breakthrough as the first to take model-dependent distribution shifts into account under the FL setting, it is only a strict generalization of perdomo2020performative without the ability to find the performative optimal point and deal with the contaminated data.
+1.2 Contributions
 In this paper, we solve all these challenges simultaneously and propose Performative robust optimal Federated Learning (ProFL), an FL algorithm that finds a performative optimal solution under endogenous distribution shifts with contaminated data.
 We provide a theoretical analysis of data contamination and the convergence of ProFL.
-Compared to the federated extension of Izzo et al., (2021), ProFL employs simple yet effective strategies and converges to the performative optimal solution for a larger class of (non-convex) objective functions.
-In Appendix B, we provide more related work and highlight differences.
-Summary of contributions. Our key contributions compared to prior works are:
-  1. 1.
-    Finding the optimal solution. Unlike Jin et al., (2024), which only guarantees convergence to the stable point, ProFL finds the optimal point.
-  2. 2.
+Compared to the federated extension of pmlr-v139-izzo21a, ProFL employs simple yet effective strategies and converges to the performative optimal solution for a larger class of (non-convex) objective functions.
+Our key contributions compared to prior works are:
+  • Performative the optimal solution. Unlike jin2023performative, which only guarantees convergence to the stable point, our ProFL finds the optimal point.
   • Convergence analysis for a larger class of objectives.
-Unlike prior works (Izzo et al.,, 2021; Jin et al.,, 2024) that show the convergence of the proposed algorithms for the convex performative risk, we show the convergence of ProFL under Polyak-Lojasiewicz (PL) condition, which is non-convex.
-  3. 3.
-    Robustness to data contamination.
-We analyze the impact of contaminated data on the FL system and show that ProFL effectively tackles complex distribution shifts and is more resilient to estimation error and data contamination.
-2 Problem Formulation
-Consider a set of clients $\mathcal{V}=\{1,\ldots,N\}$ collaboratively training an FL model $\theta\in\Theta\subset\mathbb{R}^{d}$ without sharing their local data.
-Here, $\Theta$ is a compact set (closed and bounded), and $d$ represents the model’s dimension.
-Let $\mathcal{Z}$ denote the sample space.
-Throughout training [^1] One can also interpret training as retraining or fine-tuning of models after deployment., each client interacts with the local model, causing its data to evolve based on the perceived model.
-Such model-induced data distributions, studied in PP (Perdomo et al.,, 2020), are characterized by a performative distribution map $D_{i}:\Theta\to\mathcal{P}(\mathcal{Z})$.
+Unlike prior works pmlr-v139-izzo21a; jin2023performative that show the convergence of the proposed algorithms for the convex performative risk, we show the convergence of our ProFL converges under non-convex Polyak-Lojasiewicz (PL) condition.
+  • Robustness to distribution shifts and data contamination.
+Our ProFL algorithm is designed for realistic FL settings with heterogeneous clients and limited observability.
+It integrates a principled outlier detection mechanism to enhance Robustness against client-side contamination and estimation errors.
+Despite operating without access to raw data, ProFL consistently achieves reliable convergence under dynamic and heterogeneous FL environments with contaminated data.
+2 System Model
+2.1 Performative Prediction Model
+2.1.1 Centralized Performative Prediction Model
+UserData $D(\theta^{t})$Model $\theta^{t+1}$PredictionsPredictions affect behavior, which shifts data distribution
+Figure 1: System model of performative prediction. The model $\theta$ influences the data distribution $D(\theta)$, forming a feedback loop that drives endogenous data shifts.
+Performative prediction addresses settings where the deployed model influences future data distributions (see Fig.
+1).
+Let $\theta\in\Theta\subseteq\mathbb{R}^{d}$ denote the model parameters, where $\Theta$ is closed and convex and $d$ is the dimension.
+The data distribution is $\mathcal{D}(\theta)$ which depends on $\theta$ and a distribution map $\mathcal{D}(\cdot)$.
+$z\in\mathcal{Z}$ denotes a sample from the sample space $\mathcal{Z}$.
+Denote the loss function with model $\theta$ and data sample $z$ as $\ell(z;\theta)$.
+PP considers two risk formulations.
+The performative risk (PR)is defined as
+$PR(\theta):=\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta)].$
+This is called performative risk because, instead of measuring the risk over a fixed distribution, it measures the expected loss of model $\theta$ on the data distribution $D(\theta)$ it induces.
+While the decoupled performative risk (DPR) is given by
+$DPR(\theta,\theta^{\prime}):=\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta^{\prime})].$
+Unlike PR, where $D_{i}(\theta)$ depends on $\theta$, the DPR decouples the two.
+Based on the definitions of DPR and PR, we define two key concepts: the PS and the PO points.
+The PS point $\theta^{\text{PS}}$ is a fixed point that captures an equilibrium state and is defined as:
+$\theta^{\text{PS}}:=\arg\min_{\theta}\mathbb{E}_{z\sim\mathcal{D}(\theta^{\text{PS}})}[\ell(z;\theta)].$ (1)
+The data distribution is induced by ${\theta}^{\text{PS}}$, while $\theta$ is optimized.
+As ${\theta}^{\text{PS}}$ is the fixed point of Eq.
+(1), it can be found by repeated risk minimization (RRM) or repeated gradient descent (RGD) perdomo2020performative.
+RRM recursively finds a model $\theta^{t+1}$ that minimizes DPR on the distribution $D(\theta^{t})$ induced by the previous model, whereas RGD recursively finds a model $\theta^{t+1}$ through gradient descent, and $\theta^{t+1}$ is not necessarily the minimizer of the DPR on $D(\theta^{t})$.
+The performative optimal (PO) point, $\theta^{\text{PO}}$, minimizes the performative risk (PR) and is defined as:
+$\theta^{\text{PO}}:=\arg\min_{\theta}\mathbb{E}_{z\sim\mathcal{D}(\theta)}[\ell(z;\theta)].$
+The key to finding the PO point is to update the model using the actual gradient of PR, $\nabla\mathcal{L}(\theta)$, also known as the performative gradient (PG).
+Assume the performative distribution $D(\theta)$ has a density function $p(z;f(\theta))$, where the parameter $f(\theta)$ is a differentiable function of $\theta$.
+Then, the performative risk can be written as
+$\mathcal{L}(\theta)=\int\ell(z;\theta)\,p(z;f(\theta))\,dz,$
+provided that $p$ and $f$ are continuously differentiable.
+[^1] An example of $D(\theta)$ is a mixture of Gaussians, $\sum_{k=1}^{K}\nu_{k}\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where the mean depends on the model pmlr-v139-izzo21a.
+This can approximate any smooth density to arbitrary precision.
+We now present the formal problem setting of federated performative prediction, where the objective is to find the global performative optimal point in the presence of heterogeneous data shifts and possible data contamination.
+2.1.2 Federated Performative Prediction Model
+Consider a set of clients $\mathcal{V}=\{1,\ldots,N\}$ collaboratively training an FL model $\theta$ without sharing their local data.
+Such model-induced data distributions, introduced in PP perdomo2020performative, are characterized by a performative distribution map $D_{i}:\Theta\to\mathcal{P}(\mathcal{Z})$.
 Specifically, deploying model $\theta$ at client $i$ results in data following distribution $D_{i}(\theta)$.
+The distribution map $D_{i}(\cdot)$ can vary across clients, resulting in data heterogeneity in the federated learning system.
 The goal is to train a global model that minimizes the total loss across all clients:
 ${{\theta}}^{\text{PO}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}(\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$ (2)
-where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\ell(Z_{i};\theta)]$ is the local performative risk of client $i$ for a given loss function $\ell(Z_{i};\theta)$ on local data $Z_{i}\sim D_{i}(\theta)$.
-This is called performative risk because, instead of measuring the risk over a fixed distribution, it measures the expected loss of model $\theta$ on the data distribution $D_{i}(\theta)$ it induces.
-$\mathcal{L}(\theta)$ is the performative risk (PR) of all clients.
+where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\ell(Z_{i};\theta)]$ is the PR of client $i$ for a given loss function $\ell(Z_{i};\theta)$ on local data $Z_{i}\sim D_{i}(\theta)$.
 Here, $\alpha_{i}\geq 0$ represents the fraction of client $i$’s data relative to the total, with $\sum_{i\in\mathcal{V}}\alpha_{i}=1$.
-The solution ${\theta}^{\text{PO}}$ that minimizes performative risk is called performative optimal (PO) point (Perdomo et al.,, 2020).
+In the FL setting, the PR over all clients is denoted by $\mathcal{L}(\theta)$, and the performative optimal point $\theta^{\text{PO}}$ is defined as its minimizer.
+Since the data distribution $D_{i}(\theta)$ depends on $\theta$, which is the variable being optimized, finding the PO point is often challenging.
+Thus, existing works perdomo2020performative; mendler2020stochastic; Zhao2022OptimizingTP have mostly focused on finding the PS point.
+In an FL system, the PS point ${{\theta}}^{\text{PS}}$ is defined as follows:
+${{\theta}}^{\text{PS}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}({{\theta}}^{\text{PS}};\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}({{\theta}}^{\text{PS}};\theta),$ (3)
+where $\mathcal{L}_{i}({\theta}^{\prime};\theta)=\mathbb{E}_{Z_{i}\sim D_{i}({\theta}^{\prime})}[\ell(Z_{i};\theta)]$ denotes client $i$’s DPR.
+2.2 Data Contamination Model
 In realistic FL systems, each client’s data distribution $D_{i}(\theta)$ is unknown, and the training data may be noisy, corrupted, or maliciously manipulated.
+In this work, we consider a general contamination model based on Huber’s $\epsilon$-contamination framework huber1992robust, which serves as a foundation for robust statistical analysis in the presence of outliers and data corruption.
+This model assumes that besides the data distribution we aim to learn, a small proportion of samples can come from an arbitrary distribution.
+This contamination framework is crucial for creating machine learning algorithms robust to outliers, thereby ensuring more reliable analysis in practical scenarios where data imperfections are common.
+Huber’s model has been fundamental in the field of robust statistics, influencing a wide range of applications and subsequent research.
 To account for this, we assume each client $i$ has a fraction $\epsilon_{i}\in[0,1]$ of data samples drawn from another fixed but arbitrary distribution $Q_{i}$.
@@
 We term $Q_{i}$ data “contamination”.
-Our goal is to (i) design FL algorithms that converge to the performative optimal ${\theta}^{\text{PO}}$; (ii) study the impact of data contamination $Q_{i}$ and develop robust solutions.
+2.3 Research Objectives and Challenges
+Our goal is to design FL algorithms that converge to the performative optimal ${\theta}^{\text{PO}}$ even with contaminated data $Z_{i}\sim P_{i}(\theta)$.
 While data contamination has been studied in FL with static data, its effect in performative settings is unexplored.
-A moment’s thought reveals that when client data depends on the FL system, contamination in one client may cause a cascading effect across the network, making it crucial to investigate how this disrupts FL convergence.
+When client data depends on the FL system, contamination in one client may cause a cascading effect across the network, making it crucial to investigate how this disrupts FL convergence.
 We will explore this in the paper.
-We next present the metrics in PP.
-Performative stable point.
-Since the data distribution $D_{i}(\theta)$ depends on $\theta$, which is the variable being optimized, finding the performative optimal point is often challenging.
-Thus, existing works (Perdomo et al.,, 2020; Mendler-Dünner et al.,, 2020; Zhao,, 2022) have mostly focused on finding the so-called performative stable (PS) point.
-In an FL system, the PS point ${{\theta}}^{\text{PS}}$ is defined as follows:
-${{\theta}}^{\text{PS}}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}({{\theta}}^{\text{PS}};\theta):=\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}({{\theta}}^{\text{PS}};\theta)$ (2)
-where $\mathcal{L}_{i}({{\theta}}^{{}^{\prime}};\theta)=\mathbb{E}_{Z_{i}\sim D_{i}({{\theta}}^{{}^{\prime}})}\left[\ell(Z_{i};\theta)\right]$ is called the decoupled performative risk (DPR) of client $i$.
-Unlike performative risk (PR), where the distribution $D_{i}(\theta)$ depends on $\theta$, the DPR decouples the two; i.e., the data distribution is induced by ${\theta}^{\text{PS}}$ while the variable being optimized is $\theta$.
-Since ${\theta}^{\text{PS}}$ is the fixed point of Eq.
-(2), it can be found via repeated risk minimization (RRM) or repeated gradient descent (RGD) (Perdomo et al.,, 2020).
-RRM recursively finds a model $\theta^{t+1}$ that minimizes DPR on the distribution $D(\theta^{t})$ induced by the previous model, whereas RGD recursively finds a model $\theta^{t+1}$ through gradient descent, and $\theta^{t+1}$ is not necessarily the minimizer of the DPR on $D(\theta^{t})$.
-Following the idea of RGD, Jin et al., (2024) proposed the performative federated learning (PFL) algorithm to find the PS point ${\theta}^{\text{PS}}$ in FL.
-In PFL, each client $i$ updates its local model to reduce local DPR based on the previous model’s induced distribution: $\theta_{i}^{t+1}\leftarrow\theta_{i}^{t}-\eta_{t}\nabla\ell(Z_{i}^{t+1};\theta_{i}^{t})$, where $Z_{i}^{t+1}\sim D_{i}(\theta_{i}^{t})$ is drawn from the uncontaminated distribution.
-These updates are repeated for several epochs before being sent to the server for global aggregation.
-Jin et al., (2024) showed that PFL converges to ${\theta}^{\text{PS}}$ under clean data.
-Although the PS point is easier to find than the PO point, prior works (Izzo et al.,, 2021; Miller et al.,, 2021) show that the PS point may be far from optimal or may not even exist.
-Thus, algorithms like PFL, which target the PS point, are not ideal.
-[…]
+Remark 2.1.
+FL finds the solution by alternating between global model aggregation at the central server and local model updates at the clients.
+In our proposed algorithms, we focus on the local updates, which work with different aggregation rules imposed by the server, such as differing weighted averages.
+3 Algorithm Design
+3.1 Motivation: From Stable to Optimal in Federated PP
+Most existing works in the FL or decentralized setting perdomo2020performative; mendler2020stochastic; Zhao2022OptimizingTP focus on finding the PS point, which does not minimize the true PR.
+This paper aims to find the PO point—the model that minimizes the PR under its own induced distribution.
+This goal is significantly more difficult to achieve in FL than in centralized settings due to client heterogeneity, limited data and computational resources, and data contamination, which further distort the model-dependent distributions.
+To the best of our knowledge, no FL algorithm has been shown to converge to the PO point, even under clean data and convex loss assumptions.
+This paper fills that gap by proposing an FL algorithm that provably converges to the PO point, accommodates non-convex (PL) conditions, and is robust to data contamination.
+3.2 The Proposed Algorithm
+In this section, we propose several simple yet effective strategies that collectively lead to a robust federated performative prediction algorithm, called Performative robust optimal Federated Learning (ProFL).
+The pseudocode is given in Algorithm 1.
+Unlike jin2023performative that focus on PS point, ProFL aims to achieve convergence to the global PO point in federated PP.
+Achieving the PO point requires estimating the PG, which captures both the loss and the feedback from model-induced distribution shifts.
+We begin by describing how the PG is estimated in the federated setting and then introduce our strategies for adapting this process to the challenges of FL, including client heterogeneity, limited data and computational resources, and data contamination.
+3.2.1 Federated Estimation of the Performative Gradient
+The key to finding the PO point is to update the model using the performative gradient (PG), which captures both the direct loss gradient and the effect of model-induced distributional shifts.
+While pmlr-v139-izzo21a proposed a method to estimate the PG, their analysis is limited to a centralized setting with a specific distribution map $D(\theta)$, where $f(\theta)$ is assumed to be a linear function of $\theta$, and the performative risk (PR) is convex.
+However, such assumptions are restrictive and often fail to hold in practice, especially under data contamination.
+In FL, the challenge becomes even greater.
+Unlike the centralized setting, which computes a single performative gradient $\nabla\mathcal{L}(\theta)$ on the server, FL requires each client to update its model using a local PG: $\nabla\mathcal{L}_{i}(\theta)$.
+Assume the local performative distribution $D_{i}(\theta)$ has density $p(z;f_{i}(\theta))$ with parameter $f_{i}(\theta)$ being a function of $\theta$, then $\mathcal{L}_{i}(\theta)=\int\ell(z;\theta)p(z;f_{i}(\theta))dz$ when functions $p$ and $f_{i}$ are continuously differentiable [^2] An example of $D_{i}(\theta)$ is a mixture of Gaussians, $\sum^{K}_{k=1}\nu_{k}\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where the mean depends on the model pmlr-v139-izzo21a.
+This can approximate any smooth density distribution to arbitrary precision.
+.
+PG can be derived
 $\nabla\mathcal{L}_{i}(\theta)=\nabla\mathcal{L}_{i,1}(\theta)+\nabla\mathcal{L}_{i,2}(\theta)$ (4)
@@
 $\nabla\mathcal{L}_{i,1}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\mathbf{\theta})}[\nabla\ell(Z_{i};\mathbf{\theta})]$
-and
-$\nabla\mathcal{L}_{i,2}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}\left[\ell(Z_{i};\theta)\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}\right].$
-Note that PFL proposed by Jin et al., (2024) only updates the model with $\nabla\mathcal{L}_{i,1}(\theta)$ while completely neglecting $\nabla\mathcal{L}_{i,2}(\theta)$.
-This error can accumulate, and eventually, PFL can converge to a PS point far from the optimal.
-Thus, computing the PG for a more accurate estimate is crucial, but estimating $\nabla\mathcal{L}_{i,2}(\theta)$ is challenging.
-Although Izzo et al., (2021) proposed a method to estimate the PG, its performance is only evaluated in a centralized setting under a specific performative distribution map $D_{i}(\theta)$ (i.e., when $f_{i}(\theta)$ is a linear function of $\theta$) without considering data contamination.
-Moreover, it assumes that PR is convex, which is often unrealistic.
-In summary, an FL algorithm that yields the PO point, even under a clean distribution and with rather limiting assumptions, has not been developed previously.
-This paper fills that gap by proposing an FL algorithm that provably converges to the PO point, even with non-convex PR, and handles data contamination.
-Remark 2.1.
-FL finds the solution by alternating between global model aggregation at the central server and local model updates at the clients.
-In our proposed algorithms, we focus on the local updates, which work with different aggregation rules imposed by the server, such as differing weighted averages.
-3 A Naive Extension from CS to FL
-Consider first directly extending Izzo et al., (2021) to the FL setting.
-We shall call this naive algortihm Performative optimal Federated Learning (PoFL).
-The pseudocode is provided in Appendix A.
-In this algorithm, each active client $i$ in set $\mathcal{I}_{t}\subseteq\mathcal{V}$, after synchronization with the current global model $\overline{\theta}^{t}$, uses its local data to update the model from $\theta_{i}^{t}$ to $\theta_{i}^{t+R}$.
-During these updates, the data distribution changes, and consequently, client $i$ would use samples from the contaminated distribution $P_{i}(\theta_{i}^{t})$ to update $\theta_{i}^{t}$.
+$\nabla\mathcal{L}_{i,2}(\theta)=\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}\left[\ell(Z_{i};\theta)\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p\left(Z_{i};f_{i}(\theta)\right)}{\partial f_{i}(\theta)}\right]$}
+Note that PFL jin2023performative only updates the model using $\nabla\mathcal{L}_{i,1}(\theta)$ and neglects $\nabla\mathcal{L}_{i,2}(\theta)$.
+As a result, estimation error accumulates, and the algorithm may converge to a PS point that is far from PO point.
+Therefore, accurately estimating the PG is critical; however, estimating $\nabla\mathcal{L}_{i,2}(\theta)$ remains a significant challenge in FL, particularly under data contamination.
+In ProFL, each active client $i$ in set $\mathcal{I}_{t}\subseteq\mathcal{V}$, after synchronization with the current global model $\overline{\theta}^{t}$, uses its local data to update the model from $\theta_{i}^{t}$ to $\theta_{i}^{t+R}$.
+During these updates, the data distribution evolves with the model, and consequently, client $i$ uses a sample set $\mathcal{S}_{i}$ drawn from the contaminated distribution $P_{i}(\theta_{i}^{t})$ to update its local model $\theta_{i}^{t}$.
 To find the optimal ${\theta}^{\text{PO}}=\arg\min_{\theta}\sum_{i\in\mathcal{V}}\alpha_{i}\mathcal{L}_{i}(\theta)$, each client $i$ should estimate its local performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$ to update the local model $\theta_{i}^{t}$, i.e.,
@@
 (4), $\nabla\mathcal{L}_{i}(\theta_{i}^{t})=\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})+\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$ consists of two terms.
-PoFL estimates $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$ by simply averaging $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$.
-We denote the empirical risk on data drawn from distribution $Q_{i}$ or $D_{i}(\mathbf{\theta}_{i}^{t})$ as ^$\widehat{\nabla}\mathcal{L}_{i,1}(Q_{i},\theta_{i}^{t})$ and ^$\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$, respectively.
-Then the empirical risk on $P_{i}(\mathbf{\theta}_{i}^{t})$ is ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t}).$ Let ^$\widehat{\mathcal{L}}_{i,1}(\theta_{i}^{t})$ denotes the average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$.
-We denote ^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$, ^$\widehat{\nabla}\mathcal{L}_{i,2}(Q_{i},\theta_{i}^{t})$, and ^$\widehat{\nabla}\mathcal{L}_{i,2}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$ similarly.
-^$\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ is estimated by a weighted average of $\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $P_{i}(\mathbf{\theta}_{i}^{t})$ with weight vector
-$w(z_{j};\theta_{i}^{t}):=\left.\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p(z_{j};f_{i}(\theta))}{\partial f_{i}(\theta)}\right|_{\theta=\mathbf{\theta}_{i}^{t}}\in\mathbb{R}^{d}$ (5)
-Because the distribution map $D_{i}(\theta)=p(z;f_{i}(\theta))$ and $f_{i}(\theta)$ are unknown, we need to estimate $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}.$ Following Izzo et al., (2021), we adopt finite difference approximation $\left.\frac{df_{i}(\theta)}{d\theta}\right|_{\theta=\theta^{t}_{i}}\approx\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
+To estimate $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$, ProFL applies a robust gradient strategy that filters the contaminated sample set.
+Specifically, a refined subset $\mathcal{S}^{\prime}_{i}$ is first extracted from the original sample set $\mathcal{S}_{i}\sim P_{i}(\theta_{i}^{t})$ via
+$\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}(\mathcal{S}_{i};\theta^{t}_{i};J_{1};J_{2}),$
+where $J_{1}$ and $J_{2}$ are hyperparameters controlling the selection criteria.
+The final estimate $\widehat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})$ is obtained by averaging $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples $z_{j}\in\mathcal{S}^{\prime}_{i}$, where $\mathcal{S}^{\prime}_{i}$ is the cleaned sample set selected via the robust gradient strategy (detailed in Section 3.3).
+For reference, we denote the empirical gradient estimate on samples drawn from distributions $Q_{i}$ and $D_{i}(\theta_{i}^{t})$ as $\widehat{\nabla}\mathcal{L}_{i,1}(Q_{i},\theta_{i}^{t})$ and $\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$, respectively.
+Similarly, the empirical risk over cleaned samples is denoted as $\widehat{\mathcal{L}}_{i,1}(\theta_{i}^{t})$, computed by averaging $\ell(z_{j};\theta_{i}^{t})$ over $z_{j}\in\mathcal{S}^{\prime}_{i}$.
+We also define $\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ and its counterparts $\widehat{\nabla}\mathcal{L}_{i,2}(Q_{i},\theta_{i}^{t})$ and $\widehat{\nabla}\mathcal{L}_{i,2}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})$ analogously.
+Specifically, $\widehat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$ is computed as a weighted average of $\ell(z_{j};\theta_{i}^{t})$ over the cleaned sample set $\mathcal{S}^{\prime}_{i}$, with weights determined by sensitivity to model-induced distributional parameters.
+The weight vector is defined as
+$w(z_{j};\theta_{i}^{t}):=\left.\left(\frac{df_{i}(\theta)}{d\theta}\right)^{T}\frac{\partial\log p(z_{j};f_{i}(\theta))}{\partial f_{i}(\theta)}\right|_{\theta=\mathbf{\theta}_{i}^{t}}\in\mathbb{R}^{d}$}, (6)
+which encodes how the density $p(z_{j};f_{i}(\theta))$ changes with respect to the model parameter $\theta$ through the mapping $f_{i}(\theta)$.
+Since the performative distribution $D_{i}(\theta)=p(z;f_{i}(\theta))$ and $f_{i}(\theta)$ are both unknown in practice, we must estimate both $f_{i}(\theta)$ and its Jacobian $\frac{df_{i}(\theta)}{d\theta}$ from the observed data.
+Following pmlr-v139-izzo21a, we adopt finite difference approximation $\left.\frac{df_{i}(\theta)}{d\theta}\right|_{\theta=\theta^{t}_{i}}\approx\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
 Specifically, at iteration $t$, each client $i$ collects the $H$ most recent models (excluding $\theta^{t}_{i}$) and forms a matrix $\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}$, then
-$\Delta\theta_{i}$ $=\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}-\theta^{t}_{i}\textbf{1}^{T}_{H};$ (6)
-$\Delta f_{i}$ $=\begin{bmatrix}f_{i}(\theta^{t-H}_{i})&\cdots&f_{i}(\theta^{t-1}_{i})\end{bmatrix}-f_{i}(\theta^{t}_{i})\textbf{1}^{T}_{H},$
+$\Delta\theta_{i}=\begin{bmatrix}\theta^{t-H}_{i}&\cdots&\theta^{t-1}_{i}\end{bmatrix}-\theta^{t}_{i}\textbf{1}^{T}_{H};\quad\Delta f_{i}$ $=\begin{bmatrix}f_{i}(\theta^{t-H}_{i})&\cdots&f_{i}(\theta^{t-1}_{i})\end{bmatrix}-f_{i}(\theta^{t}_{i})\textbf{1}^{T}_{H},$
 Where $\textbf{1}_{H}$ is an $H$-dimensional all-ones vector, both $\Delta f_{i}$ and $\Delta\theta_{i}$ are $d\times H$ matrices, and $(\Delta\theta_{i})^{\dagger}$ is the pseudo-inverse of $\Delta\theta_{i}$.
 For each iteration of each client, we know $\theta_{i}^{t}$, but do not know the exact value of $f_{i}(\theta_{i}^{t})$.
-Denote the estimator as $\widehat{f}_{i}(\mathcal{S}_{i})$, meaning we use the dataset $\mathcal{S}_{i}$ and the function $\widehat{f}_{i}(.)$ to estimate the value of $f_{i}(\theta_{i}^{t})$.
+Denote the estimator as $\widehat{f}_{i}(\mathcal{S}^{\prime}_{i})$, meaning we use the dataset $\mathcal{S}^{\prime}_{i}$ and the function $\widehat{f}_{i}(.)$ to estimate the value of $f_{i}(\theta_{i}^{t})$.
 This estimator is then used to compute an estimate of $\frac{df_{i}}{d\theta}$.
 The estimator of $\frac{df_{i}}{d\theta}$ is denoted as $\frac{\widehat{df_{i}}}{d\theta}$ and can be computed as $\frac{\widehat{df_{i}}}{d\theta}\approx\widehat{f}_{i}(\Delta\theta_{i})^{\dagger}$.
-Finally, $\frac{\widehat{df_{i}}}{d\theta}$ will be used to calculate ^$\widehat{\nabla}\mathcal{L}_{i,2}$.
-By combining this with ^$\widehat{\nabla}\mathcal{L}_{i,1}$, we obtain the performative gradient, allowing each client to update the local model $\theta_{i}^{t}$ using Eq.
+Finally, $\frac{\widehat{df_{i}}}{d\theta}$ will be used to calculate $\widehat{\nabla}\mathcal{L}_{i,2}$.
+By combining this with $\widehat{\nabla}\mathcal{L}_{i,1}$, we obtain the performative gradient, allowing each client to update the local model $\theta_{i}^{t}$ using Eq.
 (5).
-Concerns with POFL.
-Although Izzo et al., (2021) showed that the above method for estimating $f_{i}(\theta)$ and $\frac{df_{i}(\theta)}{d\theta}$ was effective in centralized settings, it was evaluated under specific scenarios where $f_{i}(\theta)$ is linear in $\theta$ and data samples perfectly represent $D_{i}(\theta)$ without contamination.
-In practical FL settings with model-induced distribution shift, this method will perform poorly due to the following reasons:
-  1. 1.
-    Non-linearity of $f_{i}(\theta)$.
-When $f_{i}(\theta)$ is non-linear, ignoring the higher-order terms in the Taylor expansion of $\frac{\widehat{df_{i}}}{d\theta}$ may be biased.
-While Izzo et al., (2021) demonstrated that this method works well for linear $f_{i}(\theta)$, the impact of non-linearity on performance remains unclear.
-  2. 2.
-    Amplified sampling error.
-Since $\frac{df_{i}(\theta)}{d\theta}$ is estimated from limited samples, the sampling error of the estimation $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$ is inherited when computing PG, $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$.
-As shown in Lemmas 5.7, 5.8 and  5.9 in Section 5, the sampling error in estimating $\frac{df_{i}(\theta)}{d\theta}$ is amplified when computing $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$, making the performative gradient and algorithm more sensitive to sampling errors.
-  3. 3.
-    Contaminated data.
-Due to contamination $Q_{i}$, the data used for estimating $\frac{df(\theta)}{d\theta}$ and $f_{i}(\theta)$ does not follow the actual $D_{i}(\theta)$.
-Even a small error in a single model update for one client can cascade, affecting subsequent data and the entire network.
-4 The Proposed Algorithm: PROFL
-The previous section summarized the drawbacks of a straightforward extension of the centralized solution to the federated setting.
-To address all these weaknesses, we propose several simple yet effective strategies that lead to a robust new FL algorithm.
-We shall call it Performative robust optimal Federated Learning (ProFL).
-The pseudocode is provided in Algorithm 1.
-Below are its key features.
 Algorithm 1 ProFL
-1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$, error bound $\Phi_{i}$ for all $i\in\mathcal{V}$.
-2:  for $t=0$ to $T-1$ do
+1:Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$, error bound $\Phi_{i}$ for all $i\in\mathcal{V}$, threshold $J_{1},J_{2}\in(0,1)$, number of clusters $C$.
+2:for $t=0$ to T1T-1 do
 3:  If $t\mod R=0$, server sends $\overline{\theta}^{t}$ to a selected set $\mathcal{I}_{t}\subseteq\mathcal{V}$ of clients and client $i\in\mathcal{I}_{t}$ sets the local model $\theta^{t}_{i}=\overline{\theta}^{t}$; otherwise $\mathcal{I}_{t}=\mathcal{I}_{t-1}.$
@@
 6:   if $t\leq H$ then
-7:           ^$\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right).$
+7:    $\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i};J_{1};J_{2}\right).$
 8:    Compute estimator $\widehat{f}\left(\mathcal{S^{\prime}}_{i}\right)$ of $f^{t}_{i}$.
@@
 10:   else
-11:           ^$\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right)$
+11:     $\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i});\mathcal{S}^{\prime}_{i}\leftarrow\textsc{Robust Gradient}\left(\mathcal{S}_{i};\theta^{t}_{i}\right)$
 12:    Compute estimator $\widehat{f}\left(\mathcal{S^{\prime}}_{i}\right)$ of $f^{t}_{i}$.
-13:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
+13:    Compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$ by $\Delta\theta_{i}$ and $\Delta f_{i}$.
 14:    $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\left(\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})+\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})\right)\right)$.
-15:           Input={^$=\{\widehat{\mathcal{L}}_{i,1}(\theta^{t}_{i});H;d;\eta;\varphi_{i};\Delta f_{i}(\Delta\theta_{i})^{\dagger};\Phi_{i}\}$
+[…]
 16:    $n_{i}\leftarrow\textsc{Adaptive Sample Size}\left(\texttt{Input}\right)$.
-17:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
-18:  Output: $\overline{\theta}^{T}$
-  1. 1.
-    Handling complex distribution shifts with non-linear $f_{i}(\theta)$.
-ProFL employs the following strategies: (i) reducing the learning rate $\eta$ and the estimation window $H$; (ii) estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server.
+17:    Lines 18–20 are necessary only for small $n_{i}$.
+18:    Send $\widehat{\frac{df_{i}}{d\theta}}$ to the server.
+19:    $\widehat{\frac{df_{c}}{d\theta}}\leftarrow$ ServerAggregation $(\widehat{\frac{df_{i}}{d\theta}},C)$
+20:    Compute $\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})$ using the aggregated $\widehat{\frac{df_{c}}{d\theta}}$.
+21:  If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
+22:Output: $\overline{\theta}^{T}$
+23:Server Aggregation:
+24: Use $k$-means clustering to group clients into $C$ clusters.
+25: Compute the aggregated gradient estimate $\widehat{\frac{df_{c}}{d\theta}}$ by averaging $\widehat{\frac{df_{i}}{d\theta}}$ for all $i$ in cluster $c$, for each cluster $c\in\{1,\ldots,C\}$.
+26: Send $\widehat{\frac{df_{c}}{d\theta}}$ back to the clients belonging to cluster $c$ for all $c\in\{1,\ldots,C\}$.
+3.3 Details of Robust Strategies
+Although pmlr-v139-izzo21a demonstrated that the method for estimating $f_{i}(\theta)$ and its Jacobian $\frac{df_{i}(\theta)}{d\theta}$ is effective in centralized settings, their analysis relied on that $f_{i}(\theta)$ is linear in $\theta$ and the data perfectly reflects the clean distribution $D_{i}(\theta)$.
+In federated learning, however, model-induced distribution shifts are often nonlinear and local data may be limited or contaminated, which significantly undermines the accuracy of the PG estimation used in pmlr-v139-izzo21a.
+To overcome these challenges, ProFL incorporates a set of carefully designed robustness strategies, described in detail below.
+Non-linearity of $f_{i}(\theta)$.
+When $f_{i}(\theta)$ is non-linear, ignoring the higher-order terms in the Taylor expansion of $\frac{\widehat{df_{i}}}{d\theta}$ may be biased.
+While pmlr-v139-izzo21a demonstrated that this method works well for linear $f_{i}(\theta)$, the impact of non-linearity on performance remains unclear.
+To mitigate the error introduced by non-linearity, ProFL employs the following strategies: (i) reducing the learning rate $\eta$ and the estimation window $H$; (ii) estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server.
 Reducing $\eta$ and $H$ helps reduce the estimation error resulting from non-linear $f_{i}(\theta)$ and stabilizes the algorithm.
@@
 Estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server mitigates the negative impact of sampling error.
-  2. 2.
 Balancing the impact of sampling error and computational costs.
-ProFL adaptively selects the sample size $n_{i}$ based on the error bounds of the PG and local loss.
+In practice, $\frac{df_{i}(\theta)}{d\theta}$ must be estimated from finite samples, which introduces sampling error into the computation of the performative gradient $\nabla\mathcal{L}_{i}(\theta_{i}^{t})$.
+As shown in Lemmas 4.7, 4.8 and  4.9 in Section 4, the sampling error in estimating $\frac{df_{i}(\theta)}{d\theta}$ is amplified when computing $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$, making the performative gradient and algorithm more sensitive to sampling errors.
+To mitigate this, ProFL adaptively selects the sample size $n_{i}$ based on the error bounds of the PG and local loss.
 While a larger $n_{i}$ generally leads to a more accurate estimate, it may not always be feasible in FL due to clients’ limited computational capabilities.
-By using predefined error bounds and relevant computed values—such as model variance, loss function, gradient information, and learning rates—we can determine the lower bound for each client’s $n_{i}$, balancing sampling error and computational costs.
+Therefore, ProFL balances estimation accuracy and resource constraints by computing a theoretical lower bound on $n_{i}$ using quantities such as model variance, loss curvature, learning rate, and empirical gradient variability.
 As these values change during updates, $n_{i}$ should be selected adaptively.
-More details are provided in Section 5.3.
-  3. 3.
+More details of the analysis are provided in Section 4.3.
 Handling data contamination.
-ProFL mitigates contamination $Q_{i}$ by reducing $\epsilon_{i}$ through local outlier identification and removal.
-Since the samples $\mathcal{S}_{i}\sim P_{i}(\theta_{i}^{t})$are noisy, and the raw data is not visible to the central server, contamination must be identified and cleaned locally before estimating the performative gradient $\nabla\mathcal{L}_{i}=\nabla\mathcal{L}_{i,1}+\nabla\mathcal{L}_{i,2}$.
-Given the contaminated dataset $\mathcal{S}_{i}$, each client in ProFL uses algorithm Robust gradient to identify “clean” samples $\mathcal{S}^{\prime}_{i}=\{z_{j}\in\mathcal{S}_{i}:z_{j}\sim D_{i}(\theta_{i}^{t})\}$ and compute robust gradient $\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})$.
-$\mathcal{S}^{\prime}_{i}$ are then used to estimate $\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})$.
-The outlier identification and removal mechanism reduces $\epsilon_{i}$ (the fraction of contaminated data) and can be selected based on prior knowledge of $\epsilon_{i},Q_{i}$ (if there are any).
-In Appendix A, we provide an example of Robust Gradient for a general case with unknown $\epsilon_{i}$ and arbitrary $Q_{i}$.
-It uses singular value decomposition (SVD) to detect gradients that meet two key properties: 1) having a significant impact on the model and potentially disrupting training (i.e., gradients with large magnitudes that systematically point in a specific direction); and 2) being far from the average in the vector space (i.e., “long tail” data).
-We iteratively remove gradients whose distance from the average exceeds a threshold.
-More details on the Robust Gradient mechanism and its comparison to prior work are provided in Appendix  A.
-In section 5, we analyze both algorithms’ convergence and the impact of the factors above on performance.
-Compared to PFL Jin et al., (2024) and methods in Izzo et al., (2021); Miller et al., (2021), we relax assumptions and show convergence under weaker conditions.
+Due to contamination $Q_{i}$, the data used to estimating $\frac{df(\theta)}{d\theta}$ and $f_{i}(\theta)$ does not follow the actual $D_{i}(\theta)$.
+Even a small update error from one client can cascade and impact future data and the entire network.
+In FL, because data must remain local for privacy reasons, contamination must also be detected and removed locally rather than at the central server.
+Accordingly, ProFL reduce $\epsilon_{i}$ via local outlier identification and removal.
+To estimate $\nabla\mathcal{L}_{i,1}$, we should average $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $D_{i}({\theta}_{i}^{t})$; however, the samples actually collected by client $i$ are drawn from $P_{i}(\theta_{i}^{t})$and thus include an unknown fraction $\epsilon_{i}$ of outliers from an unknown distribution $Q_{i}$.
+ProFL needs to identify contaminated data and eliminate the corresponding contaminated gradients.
+We propose a mechanism that can remove these contaminated gradients and estimate gradients that are robust to outliers.
+Importantly, ProFL is independent of the specific outlier identification and removal mechanism used and can function with any such mechanism.
+In our paper, we assume that contaminated data can follow an arbitrary distribution.
+In practice, additional information about the contamination process may be available based on the application or context, allowing us to design more efficient and less expensive outlier identification mechanisms leveraging this knowledge.
+Algorithm 2 Robust Gradient
+1:Input: Set $\mathcal{S}_{i}$ of samples $(z_{j})_{j=1}^{n_{i}}$, thresholds $J_{1},J_{2}\in(0,1)$, model $\theta^{t}_{i}\in\Theta$
+2: Let $\widehat{\nabla}\leftarrow\frac{1}{|\mathcal{S}_{i}|}\sum_{z_{j}\in\mathcal{S}_{i}}\nabla\ell(z_{j};\theta^{t}_{i})$.
+3: Let $\left[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right]_{z_{j}\in\mathcal{S}_{i}}$ be the $|\mathcal{S}_{i}|\times d$ matrix.
+4: Apply SVD to this matrix and find top right singular vector $v$.
+5: Compute $\tau_{j}=\left(\left(\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right)^{T}v\right)^{2}$ of $\nabla\ell(z_{j};\theta^{t}_{i})$.
+6: Divide the interval $0,\max\tau_{j}]$ into $B$ equal-length segments. Let $\phi_{k}$ be the number of $\tau_{j}$ in the $k$-th segment.
+7: Find the smallest $k\in\{1,\cdots,B\}$ with $\phi_{k}<J_{1}\cdot|\mathcal{S}_{i}|$. Set $\phi$ as the lower bound of the $k$-th segment.
+8: Find the set $\mathcal{S}_{i}^{\prime}\leftarrow\{z_{j}\in\mathcal{S}_{i}:\tau_{j}<\phi\}$.
+9: Compare the average gradients $\widehat{\nabla}$ of $\mathcal{S}_{i}$ and $\widehat{\nabla}^{\prime}$ of $\mathcal{S}^{\prime}_{i}$.
+If $\frac{\|\widehat{\nabla}-\widehat{\nabla}^{\prime}\|}{\|\widehat{\nabla}\|}<J_{2}$ or $|\mathcal{S}^{\prime}_{i}|\leq\frac{n_{i}}{2}$, return $\mathcal{S}^{\prime}_{i}$ and $\widehat{\nabla}^{\prime}$; otherwise, set $\mathcal{S}_{i}\leftarrow\mathcal{S}^{\prime}_{i}$ and repeat from line 2.
+As shown in Algorithm 2, our mechanism takes noisy samples $\mathcal{S}_{i}$ and current local model $\theta_{i}^{t}$ as inputs.
+The mechanism first computes the gradients $\nabla\ell(z_{j};\theta_{i}^{t})$ of all samples and iteratively identifies and removes the contaminated gradients.
+Since both the distribution $Q_{i}$ and fraction $\epsilon_{i}$ of contaminated data are unknown, we consider gradients contaminated if they exhibit two crucial properties: 1) they have large effects on the learned model and can disturb the training process (i.e., gradients have large magnitudes and systematically point in a specific direction); 2) they are located farther away from the average in the vector space (i.e., “long tail” data).
+To detect gradients that satisfy these properties, we adapt the approach in pmlr-v97-diakonikolas19a, which leverages singular value decomposition (SVD).
+Specifically, we construct a matrix $\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}]_{z_{j}\in\mathcal{S}_{i}}$ using the centered gradients and find the top right singular vector $v$, where $\widehat{\nabla}$ represents the average of all gradients (lines 2-4).
+The centered gradients that are closer to $v$ are more likely to be outliers and we assign each gradient an outlier score $\tau_{j}$ as follows:
+$\tau_{j}={\left(\left(\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right)^{T}v\right)}^{2}$ (7)
+Given the outlier scores $\{\tau_{j}\}$, the gradients with scores above a threshold $\phi$ are considered contaminated and are discarded.
+Key Innovations of Robust Gradient in ProFL.
+Unlike the method in pmlr-v97-diakonikolas19a, which requires a manually preselected threshold and resampling after each removal, ProFL determines the outlier threshold $\phi$ automatically from the empirical distribution of outlier scores $\tau_{j}$ and use the same dataset to get the roustabout gradient because it is difficult to determine without any prior knowledge of the data and sampled new data after each time of removal in impractical in PP.
+Since we assume the outliers lie in the “long tail,” we divide the range $0,\max\tau_{j}]$ into $B$ equal-length segments and compute the relative frequency of each.
+The threshold $\phi$ is chosen as the smallest segment whose relative frequency falls below a predefined level.
+This method exploits the score concentration and tail sparsity properties, making the filtering process predictable and stable.
+The final robust gradient estimate $\widehat{\nabla}^{\prime}$ is computed by averaging gradients with scores below $\phi$, ensuring resilience to noise and contamination without requiring prior knowledge.
 4 Performance Analysis
-To facilitate analysis, we consider scenarios where $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}(\theta)$ represents a mixture of $K$ group distributions.
-$\nu_{i,k}(\theta)$ is the proportion of samples from group $k$ and $\sum_{k=1}^{K}\nu_{i,k}(\theta)=1$.
-Let $f_{i,k}(.)$ denote the function for each group’s distribution.
-[…]
+To facilitate analysis, we study two special cases: (i) Contribution dynamics, where only $\nu_{i,k}(\theta)$ changes while the group distribution remains fixed, i.e., $D_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}(\theta)D_{i,k}$.
+In this case, $f_{i,k}(\theta)=\nu_{i,k}(\theta)$, and $\widehat{f}_{i,k}$ estimates the sample proportion from group $k$.
+(ii) Distribution dynamics, where only the distribution $D_{i,k}(\theta)$ changes while the contribution from each group remains fixed, i.e., $D_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}D_{i,k}(\theta)$.
+We consider a mixture of Gaussians with $D_{i,k}(\theta)=\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where $f_{i,k}(\theta)$ is the mean of group $k$ for client $i$, and the overall mean for client $i$ is $f_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}f_{i,k}(\theta)$.
+$\widehat{f}_{i}(\mathcal{S}_{i})$ represents the empirical mean from data samples $\mathcal{S}_{i}$.
+Our theoretical results apply to both cases.
+Before presenting them, we introduce the technical assumptions.
+For simplicity, $\|.\|_{2}$ is denoted as $\|.\|$.
+Proofs are in Appendix.
 Assumption 4.1.
@@
 $f_{i}(\theta)$ is twice continuously differentiable for all $\theta\in\Theta,$ i.e., the first and second derivatives $\frac{df_{i}(\theta)}{d\theta}$ and $\frac{d^{2}f_{i}(\theta)}{d\theta^{2}}$ exist and are continuous.
-Assumption 5.4 (PL Condition Polyak, (1964)).
+Assumption 4.4 (PL Condition polyak1964some).
 The local performative risk $\mathcal{L}_{i}(\theta)$ of client $i$ satisfies Polyak- Lojasiewicz (PL) condition, i.e., for all $\theta\in\Theta$, the following holds for some $\rho>0$:
@@
 Under Assumption 4.3, let $\lambda_{i,\min}$ denote the minimal singular value of $\Delta\theta_{i}$ defined in Eq.
-(6) during all iterations.
+(3.2.1) during all iterations.
 With probability at least $1-\varphi_{i}$, the estimation error of $\frac{\widehat{df_{i}}}{d\theta}$ is bounded by $\boldsymbol{\omega_{F}}$ specified as follows:
@@
 Lemma 4.9 (Estimation error of $\nabla\mathcal{L}_{i,2}$).
-Under Assumptions 5.1, 5.2 and 5.3, the following holds with probability at least $1-\varphi_{i}$:
+Under Assumptions 4.1, 4.2 and 4.3, the following holds with probability at least $1-\varphi_{i}$:
 $\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]\leq(\boldsymbol{\omega_{D}})^{2}:=\mathcal{O}\left(\ell^{2}_{\max}d\left(\boldsymbol{\omega_{F}}+\frac{f_{\varphi}(\varphi_{i};n_{i})F^{2}}{n_{i}}\right)\right).$
@@
 Additionally, the notations $\boldsymbol{\omega_{D}}$ and $\boldsymbol{\omega_{Q}}$ represent upper bounds.
-From Lemmas 5.8 and 5.9, we find that since the estimation error of performative gradient is primarily dominated by the estimation error of $\nabla\mathcal{L}_{i,2}$, we have:
+From Lemmas 4.8 and 4.9, we find that since the estimation error of performative gradient is primarily dominated by the estimation error of $\nabla\mathcal{L}_{i,2}$, we have:
 $\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{i,1}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]+\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,2}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,2}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]\leq(\boldsymbol{\omega_{D}})^{2}.$ (9)
 4.2 Convergence analysis
-Given Lemmas 5.7, 5.8, and 5.9, we are now ready to analyze the convergence.
+Given Lemmas 4.7, 4.8, and 4.9, we are now ready to analyze the convergence.
 Denote $\overline{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}\gamma_{i}$, $\sigma^{2}_{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}(\overline{\gamma}-\gamma_{i})^{2}$, $\overline{\epsilon}=\sum_{i\in\mathcal{V}}\alpha_{i}\epsilon_{i}$, and $\varphi=\sum_{i\in\mathcal{V}}\varphi_{i}.$ Throughout the training process, let $W_{1}(D,Q)_{\max}$ represent the maximum Wasserstein-1 distance $W_{1}\left(D(\theta_{i}^{t}),Q_{i}\right)$ between actual data distribution and contamination $\forall i\in\mathcal{V}$.
@@
 Theorem 4.10 (Convergence rate).
-Let $\{\overline{\theta}^{t}\}_{t\geq 0}$ be a sequence of global models generated by ProFL (and PoFL). Under Assumptions 5.1, 5.2, 5.3, and 5.4, with probability at least $1-T\varphi$, we have:
-$\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{T}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]\leq$ $\left(1-\rho\eta\right)^{T}\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{0}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}-\frac{g^{2}_{\min}}{2}\right)\eta$
-$\quad+\left(L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
-$\quad+\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\eta^{3}.$
+Let $\{\overline{\theta}^{t}\}_{t\geq 0}$ be a sequence of global models generated by ProFL (and PoFL). Under Assumptions 4.1, 4.2, 4.3, and 4.4, with probability at least $1-T\varphi$, we have:
+$\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{T}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]\leq$
+$~\left(1-\rho\eta\right)^{T}\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{0}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]$
+$+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}-\frac{g^{2}_{\min}}{2}\right)\eta$
+$+\left(L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
+$+\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\eta^{3}.$
 where $\overline{\boldsymbol{\omega_{D,Q}}}=\overline{\boldsymbol{\omega_{D}}}+LW_{1}(D,Q)_{\max}+\overline{\boldsymbol{\omega_{Q}}}.$
-Theorem 5.10, together with Lemmas 5.7, 5.8 and 5.9 in Section 5.1, highlight the impact of sampling error, data contamination, client heterogeneity, and the non-linearity of $f_{i}(\theta)$ on convergence, as discussed below.
+Theorem 4.10, together with Lemmas 4.7, 4.8 and 4.9 in Section 4.1, highlight the impact of sampling error, data contamination, client heterogeneity, and the non-linearity of $f_{i}(\theta)$ on convergence, as discussed below.
 When there is no data contamination and $\epsilon_{i}=0$, $\boldsymbol{\omega_{F}}$ is reduced to $\left(\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}+\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}\right)\lambda_{i,\min}^{-2}$, where the first term is due to the non-linearity of $f_{i}(\theta)$ and becomes zero if $f_{i}(\theta)$ is linear with $M=0$; the second term, related to sampling error, approaches zero as the sample size $n_{i}\to\infty$.
@@
 The term $\overline{\epsilon}(\overline{\boldsymbol{\omega_{D}}}+LW_{1}(D,Q)_{\max}+\overline{\boldsymbol{\omega_{Q}}})^{2}$ reflects the impact of contaminated data, increasing as the fraction of contamination $\epsilon_{i}$ and/or the distance between actual data distribution $D_{i}(\theta)$ and contamination $Q_{i}$ increase.
-The term $L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$ captures the effect of client heterogeneity.
+The term $L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$ captures client heterogeneity.
 Recall that $R$ is the number of local updates.
@@
 If $\overline{\epsilon}=0$ (without contamination), the bound in Theorem 4.10 is reduced to the following:
-$\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq$ $(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right]+\left(\overline{\boldsymbol{\omega_{D}}}^{2}-\frac{g^{2}_{\min}}{2}\right)\eta+L(1+\overline{\gamma})\left(\overline{\boldsymbol{\omega_{D}}}^{2}+G^{2}\right)\eta^{2}$
-$\quad+\left(\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)\eta^{3}.$
+$\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right]$
+$+\left(\overline{\boldsymbol{\omega_{D}}}^{2}-\frac{g^{2}_{\min}}{2}\right)\eta+L(1+\overline{\gamma})\left(\overline{\boldsymbol{\omega_{D}}}^{2}+G^{2}\right)\eta^{2}$
+$+\left(\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)\eta^{3}.$
 Because $g^{2}_{\min}\geq 0,$ we consider two cases:
-  1. 1.
-    If $g^{2}_{\min}>0$, we can always adjust learning rate $\eta$, estimation window $H$, number of local updates $R$, and sample size $n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
+1) If $g^{2}_{\min}>0$, we can always adjust learning rate $\eta$, estimation window $H$, number of local updates $R$, and sample size $n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
 $\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right],$}
 i.e., $\overline{\theta}^{t}$ converges to ${\theta}^{\text{PO}}$ at linear convergence rate.
-  2. 2.
-    If $g^{2}_{\min}=0$, this means there is at least one iteration $t$ that has $\|g^{t}\|=0$ and $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$.
+2) If $g^{2}_{\min}=0$, this means there is at least one iteration $t$ that has $\|g^{t}\|=0$ and $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$.
 Except for the iterations where $\|g^{t}\|=0$, all the other iterations have positive $\|g^{t}\|>0$.
@@
 $\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T-b}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right],$}
-    where $b$ is the number of iterations with $\|g^{t}\|=0$.
-All above theoretical results apply to both ProFL and PoFL. The robust strategies in ProFL reduce the upper bound in Theorem 5.10.
+where $b$ is the number of iterations with $\|g^{t}\|=0$. All above theoretical results apply to both ProFL and PoFL. The robust strategies in ProFL reduce the upper bound in Theorem 4.10.
 4.3 Discussion
-Section 3 introduced simple yet effective strategies in ProFL to enhance PoFL. We now explain why these strategies are effective.
-[…]
+Handling complex distribution shifts with non-linear $f_{i}(\theta)$.
+Based on Lemma 4.7, the non-linear function $f_{i}(\theta)$ primarily impacts the estimation error through the term $\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}$, which decreases as $\eta$ andHH are reduced.
+However, to ensure the matrix $\Delta\theta_{i}$ in Eq.
+(3.2.1) is nonsingular, H1H-1 must be larger than the dimensionality of $\theta$.
+In cases where $H$ cannot be decreased, reducing $\eta$ is essential to attain a smaller $\boldsymbol{\omega_{F}}$.
+Moreover, the upper bound in Theorem 4.10 takes the form $A\eta+B\eta^{2}+C\eta^{3}$, where $B,C,\eta>0$.
+Thus, a smaller $\eta$ results in a smaller upper bound.
+However, reducing $\eta$ will also increase the error term $\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}$ in Eq.
+(8).
+Nevertheless, if the sample size $n_{i}$ is sufficiently large and $\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}\gg\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}$, reducing the learning rate is still effective, as we verify in Fig.
+3(b).
+In cases where local samples are limited, estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server side could be considered, e.g., by clustering clients with similar performative distribution map $D_{i}$ and perform global aggregation to mitigate estimation error.
 Balancing Sampling Error and Computational Costs.
@@
 5 Experiments
-Experimental Setup.
-We evaluate our algorithm in five case studies: 1) pricing with dynamic demands, 2) pricing with dynamic contribution, 3) binary strategic classification, 4) house pricing regression, and 5) regression with dynamic contribution.
-Pricing with Dynamic Demands.
-Consider a company that aims to find the proper prices for various goods.
-Let $\theta\in\mathbb{R}^{d}$ be the prices associated with the set of $d$ goods, and $Z_{i}\in\mathbb{R}^{d}$ be the retailer $i$’s demands for these goods.
-We assume $Z_{i}\sim\mathcal{N}(\mu(\theta),\sigma^{2})$ is model dependent Gaussian distribution, i.e., mean demand for each good decreases linearly as the price increases.
-$\mu(\theta)=\mu_{0}-\gamma_{i}\theta$, here $\gamma_{i}\geq 0$ measures the price sensitivity and varies across different retailers.
-The goal is to maximizes total revenue over all retailers $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\theta^{T}Z_{i}]$.
-Pricing with Dynamic Contribution.
-Consider a similar setting as above where a company aims to find prices for various goods that maximize the total revenue.
-Suppose the retailers it faces have consumers from $K$ groups (e.g., old/young) and these consumers have fixed demands but they have options to purchase the goods from other companies.
-Here we assume the demands of retailer $i$’s consumers from group $k\in[K]$ have fixed distribution $D_{i,k}=\mathcal{N}(\mu_{i,k},\sigma_{i}^{2})$, but the fraction of groups change dynamically based on prices $\theta$.
-Let $r_{i,k}(\theta)=B-\mathbb{E}_{Z_{i}\sim D_{i,k}}[\theta^{T}Z_{i}]$ be the expected remaining budget of group $k$ after purchasing the goods (with initial budget $B$) and assume the fraction of group $k$ is $\nu_{i,k}(\theta)=\frac{r_{i,k}(\theta)}{\sum_{k=1}^{K}r_{i,k}(\theta)}$ (i.e., groups with more remaining budgets have more incentives to stay in the system).
-Then the retailer $i$’s total demands would follow the mixture of distributions $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}$.
-Binary Strategic Classification.
-Consider an FL system is used to make binary decisions about human agents (e.g., loan application, hiring, college admission), where client $i$ with data $Z_{i}=(X_{i},Y_{i})$ may manipulate their features strategically to increasing their chances of receiving favorable outcomes at the lowest costs.
-Suppose the clients are assigned positive decisions whenever $21/\left(1+\exp(-X^{T}\theta)\right)\geq 1/$ 2 and the cost it takes to manipulate feature from $X_{i}$ to $X_{i}^{\prime}$ is $\frac{1}{2\gamma_{i}}||X^{\prime}-X||$, then strategic clients would react to the decision model $\theta$ by manipulating their features to $\arg\max_{X_{i}^{\prime}}\left(-\langle X^{\prime}_{i},\theta\rangle-\frac{1}{2\gamma_{i}}\|X_{i}^{\prime}-X_{i}\|_{2}^{2}\right)=X_{i}-\gamma_{i}\theta$.
-For the initial data, we acquire it from a synthetic data where $X|Y=y\sim\mathcal{N}(\mu_{y},\sigma_{y}^{2})$ for $y\in\{0,1\}$ and two realistic Adult and Give Me Some Credit data.
-The details of these data are in Appendix E.1.
-The goal is to minimizes prediction loss over all clients.
-House Pricing Regression.
-Consider a regression task that aims to find the proper listing prices for houses based on features such as size, age, number of bedrooms.
-Suppose the features of houses in district $i$ follow $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$.
-Given predictive model $\theta$, let the listing price be $h(X_{i})=\theta^{T}X_{i}$.
-Since the houses with higher listing prices tend to lower the demand, the actual selling price $Y$ depends on $\theta$ and we assume $Y_{i}=(a_{i}-\gamma_{i}\theta)^{T}X_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{n})$ is Gaussian noise.
-The goal is to minimizes prediction error
-Regression with dynamic contribution.
-Consider an example of retail inventory management, where the regression model $\theta$ predicts product sales considering features like seasonal trends and advertising.
-Distributor $i$ has retailers from two groups $k\in\{1,2\}$, and each retailer’s data distribution is fixed.
-The fraction of retailers from each group is $\nu_{i,k}(\theta)=\frac{\ell_{i,-k}(\theta)+c}{\sum_{k^{\prime}\in\{1,2\}}\left(\ell_{i,k^{\prime}}(\theta)+c\right)}$ where $-k=\{1,2\}\setminus k$ because the retailers are more likely to stop purchasing if they experience higher prediction error $\ell_{i,k}(\theta)$.
-The goal is to find $\theta$ that minimizes expected prediction error $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{(X_{i},Y_{i})\sim D_{i}(\theta)}\left[(\theta^{T}X_{i}-Y_{i})^{2}\right]$, where $Y_{i}$ is a weighted combination of fixed distributions.
-For binary classification, we use both synthetic and real data (Adult and Give Me Some Credit datasets), while the other cases use synthetic data.
-All dynamics are synthetic.
-The parameter choices is in Table 3 of Appendix E.1.
-Information on the real datasets is in Appendix E.4.
-Each method is run 10 times under the same setup, and averages and variance are reported.
+5.1 Experimental Setup.
+Case Studies.
+We evaluate our algorithm on five case studies covering both types of dynamics introduced in Section 4: (i) Contribution dynamics, where group contributions change dynamically; and (ii) Distributional dynamics, where group distributions change dynamically.
+We adopt the same mixture-model formulation $D_{i}(\theta)$ as in the theoretical analysis.
+Specifically, we study five scenarios covering both dynamics: (1) Dynamic Pricing with Demand Shift—prices influence mean demands across retailers; (2) Dynamic Pricing with Group Contributions—group shares of demand adjust based on remaining budgets; (3) Binary Strategic Classification—clients strategically manipulate features to maximize favorable outcomes; (4) House Pricing Regression—listing and selling prices interact with the model’s predictions; and (5) Regression with Dynamic Contributions—group fractions of retailers adapt to model prediction errors.
+Scenarios (2) and (5) correspond to Contribution dynamics, while the remaining scenarios correspond to Distributional dynamics.
+Client Configuration and Heterogeneity.
+For each experimental setting, every method is independently run 10 times with random seeds 0–9.
+We report the average and variance across these runs.
+Table 1 and Figures 23 (case study: pricing with dynamic demands) use 10 heterogeneous clients, except the $\alpha=0$ case in Table 1, which uses homogeneous clients, and Figure 2(b), which uses 100 clients.
+Figures 4 (binary strategic classification), 6 (house pricing regression), 5(a) (pricing with dynamic contribution), and 5(b) (regression with dynamic contribution) all use 10 heterogeneous clients.
 Baselines.
-To date, only Jin et al., (2024) has addressed model-dependent distribution shifts in FL.
+To date, only jin2023performative has addressed model-dependent distribution shifts in FL.
 Thus, we compare our algorithm with this Performative Federated Learning (PFL).
 Unlike our algorithm, theirs converges to a performative stable solution ${\theta}^{\text{PS}}$ under stricter assumptions.
-We also compare the Performative Gradient (PG) algorithm from Izzo et al., (2021), in the centralized setting.
-Table 1: Performative Loss of PG and ProFL
+We also compare with the Performative Gradient (PG) algorithm from pmlr-v139-izzo21a in the centralized setting, as well as a straightforward extension of PG to the FL setting without our robust strategies, which we refer to as PoFL (Performative Optimal Federated Learning).
+Dataset.
+Binary classification uses both synthetic and real data (Adultmisc_adult_2 and Give Me Some CreditGiveMeSomeCredit datasets), while the other cases use synthetic data.
+All dynamics are synthetic.
+The supplementary material provides complete descriptions of the case study settings, data-generation models, training configurations, and all parameter values used in the experiments.
+5.2 Numerical Results
+Comparison to Centralized Setting.
+Table 1 shows that adapting PG to the FL framework improves performance.
+$\gamma_{i}\in[\overline{\gamma}-\alpha,\overline{\gamma}+\alpha]$.
+Larger $\alpha$ indicates more heterogeneity.
+As client heterogeneity increases, the performance of both algorithms degrades.
+Under homogeneous distribution shifts ($\alpha=0$), both algorithms perform similarly.
+ProFL outperforms PG at $\alpha=\{0.25,0.5\}$, since the centralized method estimates only one $\frac{df}{d\theta}$ on the server and fails under variance shifts.
 $\alpha$ 0 0.25 0.5
@@
 ProFL 5.56 $\pm$ 0.00 5.59 $\pm$ 0.00 5.67 $\pm$ 0.00
+Table 1: Performative risk of ProFL (federated) and PG (centralized) under varying client heterogeneity ($\alpha$). ProFL outperforms PG as heterogeneity increases.
 (a) Depend on Learning Rates
@@
 (c) Adaptive Learning Rate
-Figure 1: Effectiveness of Proposed methods
+Figure 2: (a) Robust gradient mitigates the impact of data contamination.
+(b) Estimation of $\frac{df_{i}}{d\theta}$ on the server outperforms local estimation in the large-client, limited-sample setting ($|\mathcal{V}|=100$, $n_{i}=60$).
+(c) Adaptive sampling (tolerances 0.05/0.1) reduces sample usage while maintaining accuracy.
 (a) Different Sample Sizes
@@
 (c) Different H
-Figure 2: Dynamic contribution with Different Sample Sizes, Learning Rates, and $H$
-Comparison to Centralized Setting.
-Table 1 shows that adapting performative gradients to the FL framework improves performance.
-There are 10 clients with $\gamma_{i}\in[\overline{\gamma}-\alpha,\overline{\gamma}+\alpha]$, where larger $\alpha$ indicates more heterogeneity.
-As client heterogeneity increases, the performance of both algorithms degrades.
-Under homogeneous distribution shifts ($\alpha=0$), both algorithms perform similarly.
-ProFL (ours) outperforms PG when $\alpha=0.25$ or $\alpha=0.5$, as the centralized approach struggles to handle variance shifts by estimating a single $\frac{df}{d\theta}$ on the server.
+Figure 3: (a) Increasing sample size $n_{i}$ improves stability and convergence to the PO point.
+(b) A smaller learning rate $\eta$ slows convergence but yields solutions closer to the PO point.
+(c) A smaller estimation window $H$ produces more local estimates of $\frac{df_{i}}{d\theta}$, reducing final error at the cost of slower convergence.
 Effectiveness of our proposed methods.
 Fig.
-1(a) shows that contamination can significantly degrade performance, but the proposed robust gradient can effectively reduce the impact of contaminated data when computing the gradients i.e., the performance with the robust gradient in our method is almost the same as the case when outliers do not exist at all.
-This case study, pricing with dynamic demands, involves 10 heterogeneous clients with different $D_{i}(\theta)$, each encounters an $\epsilon_{i}$ fraction of contaminated data from a fixed but unknown exogenous distribution $Q_{i}$.
+2(a) shows that contamination can significantly degrade performance, but the proposed robust gradient can effectively reduce the impact of contaminated data when computing the gradients i.e., the performance with the robust gradient in our method is almost the same as the case when outliers do not exist at all.
+Each client encounters an $\epsilon_{i}$ fraction of contaminated data from a fixed but unknown exogenous distribution $Q_{i}$.
+We select $B=50$, $J_{1}=0.01$, and $J_{2}=0.05$ for the robust gradient method.
 Although we present the results when $Q_{i}\sim\mathcal{N}(\mu_{o},\sigma^{2}_{o})$, we observed similar results when $Q_{i}$ follows other distributions and are different among clients.
 Fig.
-1(b) shows that with a limited sample size and many clients ($n_{i}=60,|\mathcal{V}|=100$), estimating $\widehat{\frac{df_{i}}{d\theta}}$ on the server side results in a lower loss value compared to local estimation and approaches the optimal value.
+2(b) shows that with a limited sample size and many clients ($n_{i}=60,|\mathcal{V}|=100$), estimating $\widehat{\frac{df_{i}}{d\theta}}$ on the server side results in a lower loss value compared to local estimation and approaches the optimal value.
 Fig.
-1(c) demonstrates the effectiveness of adaptive sample sizing.
+2(c) demonstrates the effectiveness of adaptive sample sizing.
 We tested two tolerances (0.05 for purple, 0.1 for red), with each client processing up to 1000 samples per iteration.
 The adaptive approach significantly reduces the sample size while achieving similar results within the error tolerance, leading to substantial computational savings.
-Both experiments are based on the case study pricing with dynamic contribution, involving a non-linear $f_{i}(\theta)$.
+Both experiments have non-linear $f_{i}(\theta)$.
 Impact of hyperparameters.
@@
 Fig.
-2(a) shows that increasing $n_{i}$ improves stability and brings the algorithm closer to the optimal point.
+3(a) shows that increasing $n_{i}$ improves stability and brings the algorithm closer to the optimal point.
 Fig.
-2(b) indicates that a smaller $\eta$ slows convergence but moves $\theta$ nearer to the optimum.
+3(b) indicates that a smaller $\eta$ slows convergence but moves $\theta$ nearer to the optimum.
 Fig.
-2(c) demonstrates that a smaller $H$ results in estimates estimates of $\frac{df_{i}}{d\theta}$ based on closer but smaller groups of $\widehat{f_{i}}$ and $\theta_{i}$, resulting in convergence closer to the performative optimum, albeit at a slightly slower speed.
-Table 2: Accuracy of Binary Classification on Synthetic and Realistic Datasets
-DATASET POFL PROFL
-Same Distribution 88.00 $\pm$ 2.69 99.23 $\pm$ 0.29
-Different Distributions 62.44 $\pm$ 0.52 92.50 $\pm$ 0.29
-Credit Dataset 94.73 $\pm$ 0.00 94.73 $\pm$ 0.00
-Adult Dataset 54.74 $\pm$ 0.02 60.78 $\pm$ 0.81
+3(c) demonstrates that a smaller $H$ results in estimates estimates of $\frac{df_{i}}{d\theta}$ based on closer but smaller groups of $\widehat{f_{i}}$ and $\theta_{i}$, resulting in convergence closer to the performative optimum, albeit at a slightly slower speed.
+(a) Same Distributions
+(b) Different Distributions
+(c) Give Me Some Credit
+(d) Adult
+Figure 4: Convergence comparison between ProFL and PFL on both synthetic and real-world datasets. ProFL consistently achieves higher accuracy with fewer communication rounds.
 Convergence to the performative optimal solution.
-Table 2 shows that ProFL outperforms PFL in the case study of binary strategic classification with 10 heterogeneous clients.
-Since PFL cannot handle contaminated data, we set $\epsilon_{i}=0$ for these experiments.
-The first two rows show that ProFL is more stable and significantly more accurate than PFL on Gaussian synthetic data.
-PFL converges to a PS point rather than a PO point.
-On the Give Me Some Credit dataset (third row), both algorithms converge to the same point, but ProFL has a faster convergence rate (see Fig.
-7(c) in the Appendix).
-With the Adult dataset, ProFL achieves $60.78\%$ accuracy, outperforming PFL’s $54.74\%$.
-These results highlight ProFL’s practical advantages in both convergence speed and accuracy.
-More results are in Appendix D.1.
-Appendix A Algorithms and Details
-A.1 Pseudocode of the PoFL
-Algorithm 2 presents the pseudocode of the PoFL algorithm discussed in the main paper.
-Algorithm 2 PoFL
-1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$.
-2:  for $t=0$ to $T-1$ do
-3:     If $t\mod R=0$, server sends $\overline{\theta}^{t}$ to a selected set $\mathcal{I}_{t}\subseteq\mathcal{V}$ of clients and client $i\in\mathcal{I}_{t}$ sets the local model $\theta^{t}_{i}=\overline{\theta}^{t}$; otherwise $\mathcal{I}_{t}=\mathcal{I}_{t-1}.$
-4:     for $i\in\mathcal{I}_{t}$ in parallel do
-5:        Deploy $\theta^{t}_{i}$ and local environment changes and draw $n_{i}$ samples $\mathcal{S}_{i}:=(z_{j})^{n_{i}}_{j=1}\mathop{\sim}\limits^{iid}P_{i}(\theta^{t}_{i})$.
-6:        if $t\leq H$ then
-7:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
-8:           Update $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})\right)$.
-9:        else
-10:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
-11:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
-12:           $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\left(\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})+\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})\right)\right)$.
-13:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
-14:  Output: $\overline{\theta}^{T}$
-Algorithm 3 Robust Gradient
-1:  Input: Set $\mathcal{S}_{i}$ of samples $(z_{j})_{j=1}^{n_{i}}$, thresholds $C,J\in(0,1)$, model $\theta^{t}_{i}\in\Theta$
-2:  Let ^$\widehat{\nabla}\leftarrow\frac{1}{|\mathcal{S}_{i}|}\sum_{z_{j}\in\mathcal{S}_{i}}\nabla\ell(z_{j};\theta^{t}_{i})$.
-3:  Let $\left[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right]_{z_{j}\in\mathcal{S}_{i}}$ be the $|\mathcal{S}_{i}|\times d$ matrix of centered gradients.
-4:  Apply SVD to this matrix and find top right singular vector $v$.
-5:  Compute the outlier score $\tau_{j}$ of $\nabla\ell(z_{j};\theta^{t}_{i})$ by Eq. (9).
-6:  Divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments. Let $\phi_{k}$ be the number of $\tau_{j}$ in the $k$-th segment.
-7:  Find the smallest $k\in\{1,\cdots,B\}$ with $\phi_{k}<C\cdot|\mathcal{S}_{i}|$. Set $\phi$ as the lower bound of the $k$-th segment.
-8:  Find the set $\mathcal{S}_{i}^{\prime}\leftarrow\{z_{j}\in\mathcal{S}_{i}:\tau_{j}<\phi\}$.
-9:  Compare the average gradients $\widehat{\nabla}$ of $\mathcal{S}_{i}$ and $\widehat{\nabla}^{\prime}$ of $\mathcal{S}^{\prime}_{i}$.
-If ^^^$\frac{\|\widehat{\nabla}-\widehat{\nabla}^{\prime}\|}{\|\widehat{\nabla}\|}<J$ or $|\mathcal{S}^{\prime}_{i}|\leq\frac{n_{i}}{2}$, return $\mathcal{S}^{\prime}_{i}$ and $\widehat{\nabla}^{\prime}$; otherwise, set $\mathcal{S}_{i}\leftarrow\mathcal{S}^{\prime}_{i}$ and repeat from line 2.
-A.2 Outlier removal mechanism for robust gradient
-To estimate $\nabla\mathcal{L}_{i,1}$, we should average $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $D_{i}({\theta}_{i}^{t})$.
-Because the samples collected by client $i$ in round $t$ are drawn from $P_{i}(\theta_{i}^{t})$, they are noisy, with an unknown $\epsilon_{i}$ fraction of outliers from an unknown distribution $Q_{i}$.
-ProFL needs to identify contaminated data and eliminate the corresponding contaminated gradients.
-We propose a mechanism that can remove these contaminated gradients and estimate gradients that are robust to outliers.
-Importantly, ProFL is independent of the specific outlier identification and removal mechanism used and can function with any such mechanism.
-In our paper, we assume that contaminated data can follow an arbitrary distribution.
-In practice, additional information about the contamination process may be available based on the application or context, allowing us to design more efficient (and less expensive) outlier identification mechanisms leveraging this knowledge.
-As shown in Algorithm 3, our mechanism takes noisy samples $\mathcal{S}_{i}$ and current local model $\theta_{i}^{t}$ as inputs.
-The mechanism first computes the gradients $\nabla\ell(z_{j};\theta_{i}^{t})$ of all samples and iteratively identifies and removes the contaminated gradients.
-Since both the distribution $Q_{i}$ and fraction $\epsilon_{i}$ of contaminated data are unknown, we consider gradients contaminated if they exhibit two crucial properties: 1) they have large effects on the learned model and can disturb the training process (i.e., gradients have large magnitudes and systematically point in a specific direction); 2) they are located farther away from the average in the vector space (i.e., “long tail” data).
-To detect gradients that satisfy these properties, we adapt the approach in Shan et al., (2023), which leverages singular value decomposition (SVD).
-Instead of preselecting the threshold for the sum of all outlier scores $\tau_{j}$, which is difficult to determine without any prior knowledge of the data, our threshold is based on the two crucial properties, making it much easier to predict the values of $C$ and $J$.
-Specifically, we construct a matrix $[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}]_{z_{j}\in\mathcal{S}_{i}}$ using the centered gradients and find the top right singular vector $v$, where $\widehat{\nabla}$ represents the average of all gradients (lines 2-4).
-The centered gradients that are closer to $v$ are more likely to be outliers and we assign each gradient an outlier score $\tau_{j}$ as follows:
-$\tau_{j}={\left(\left(\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right)^{T}v\right)}^{2}$ (9)
-Given the outlier scores $\{\tau_{j}\}$, the gradients with scores above a threshold $\phi$ are considered contaminated and are discarded.
-Unlike Shan et al., (2023) that uses a fixed pre-defined threshold, our method finds $\phi$ automatically based on the outlier score distribution.
-Since we assume the outliers are in the “long tail” of the distribution, we divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments and compute the relative frequency distribution of scores.
-$\phi$ is set as the smallest segment with the relative frequency below a threshold (lines 6-7).
-For the remaining gradients with $\tau_{j}<\phi$, we compute their average $\widehat{\nabla}^{\prime}$ (lines 8-9).
-Appendix B Additional Related Work
-B.1 Performative Prediction in Centralized Settings
-Performative prediction (Perdomo et al.,, 2020) was first introduced in 2020 corresponding to the optimization framework to deal with endogenous data distribution shifts where the model deployed to the environment will affect the subsequent data, resulting in the collected data changing as the deployed model changes.
-Common applications of performative prediction include strategic classification(Hardt et al.,, 2016; Xie and Zhang,, 2024) and predictive policing (Ensign et al.,, 2018).
-Perdomo et al., (2020) first proposes an iterative optimization procedure named Repeated Risk Minimizationto find a performative stable point and also bounded the distance between the performative stable point and the performative optimal point.
-Mendler-Dünner et al., (2020) designed the first algorithm to find the performative stable point under the online learning setting.
-Miller et al., (2021) first solved the problem of performative prediction by directly optimizing performance risk to find the performative optimal point, but the scope of the distribution map is restricted.
-Mendler-Dünner et al., (2022) proved the convergence of greedy deployment and lazy deployment after each random update under the assumption of smoothness and strong convexity and gave the convergence rate.
-Izzo et al., (2021) also designed an algorithm performative gradient descent for finding the performative optimal solution.
-Zhao, (2022) relaxes the assumption of strong convexity to the weakly-strongly convex case and proves the convergence of the deployment.
-Somerstep et al., (2024); Bracale et al., (2024) focused on learning the distribution map in performative prediction and proposed a framework named Reverse Causal Performative Prediction.
-In addition, performative prediction is also related to reinforcement learning(Zheng et al.,, 2022) and bandit problems(Chen et al.,, 2024).
-B.2 Performative Prediction in Distributed System
-To the best of our knowledge, Jin et al., (2024) is the only work discussing performativity under the distributed setting, where they generalized Fedavg to P-Fedavg and proved the uniqueness of the performative stable solution found by the algorithm.
-They also quantified the distance between the performative stable solution and the performative optimal solution.
-While the work is a breakthrough as the first to take model-dependent distribution shifts into account under the FL setting, it is only a strict generalization of Perdomo et al., (2020) without the ability to find the performative optimal point and deal with the contaminated data.
-B.3 Federated Learning under Dynamic Data
-Previous works (Tan et al.,, 2023; Li et al., 2020a, ) pointed out that the training paradigm of FedAvg may violate the i.i.d.
-assumptions and cause some feature classes to be over-represented while others to be under-represented.
-Li et al., 2020b ; Li et al., (2021) steer the local models towards a global model by adding a regularization term to guarantee convergence when the data distributions among different clients are non-IID.
-Li et al., 2020b proposes FedProx as a solution.
-Li et al., (2021) aims at obtaining good local models on clients rather than a global consensus model.
-Meanwhile, another line of work focused on dealing with the statistical heterogeneity by clustering (Briggs et al.,, 2020; Ghosh et al.,, 2020; Sattler et al.,, 2020) by aggregating clients with similar distribution into the same cluster.
-While most previous works assumed the “dynamic” is among different clients, another line of literature focuses on FL in a dynamic environment where various distribution shifts occur.
-Some works aim to deal with time-varying contribution rates of clients with local heterogeneity (Sattler et al.,, 2021; Park et al.,, 2021; Wang and Ji,, 2022; Zhu et al.,, 2021).
-B.4 Huber’s $\epsilon$-contamination Model
-Proposed by Huber, (1992), Huber’s epsilon contamination model serves as a framework for robust statistical analysis with outliers and contaminations in the dataset.
-This model assumes that besides the data distribution we aim to learn, a small proportion of samples can come from an arbitrary distribution.
-This contamination framework is crucial for creating machine learning algorithms robust to outliers, thereby ensuring more reliable analysis in practical scenarios where data imperfections are common.
-Huber’s model has been fundamental in the field of robust statistics, influencing a wide range of applications and subsequent research.
-Specifically, denote $D$ as the true data distribution and $Q$ as a random distribution.
-$P=(1-\epsilon)D+Q$ is the distribution with some corruption $Q$ and the portion of corruption data is $\epsilon$.
-The local dataset of client $i$ is $P_{i}=(1-\epsilon_{i})D_{i}(\theta_{i})+\epsilon_{i}Q_{i}$.
-Appendix C Additional Proofs not Detailed in the Main Paper
-C.1 Proof of Theorem 5.10
+PFL converges to the PS point $\theta^{\text{PS}}$, which can be suboptimal, whereas ProFL consistently converges closer to the PO point $\theta^{\text{PO}}$.
+We conduct several experiments across diverse case studies to support this claim.
+On Gaussian synthetic data, ProFL demonstrates significantly better accuracy and stability than PFL (Figs.
+4(a)4(b)), which converges to $\theta^{\text{PS}}$.
+On real datasets (Figs.
+4(c)4(d)), ProFL achieves faster convergence and higher final accuracy (e.g., $60.78\%$ vs.
+$54.74\%$ on Adult).
+Since PFL is sensitive to outliers, we set $\epsilon_{i}=0$ for fair comparison.
+In the non-linear $f_{i}(\cdot)$ cases (Fig.
+5(a) and 5(b)), ProFL again converges faster and reaches a point much closer to $\theta^{\text{PO}}$, while PFL remains at the suboptimal $\theta^{\text{PS}}$.
+Additionally, in Fig.
+5(b), ProFL demonstrates more stable convergence despite a persistent gap to $\theta^{\text{PO}}$ caused by the non-linearity of $f_{i}(\cdot)$.
+Strategies to mitigate this gap are discussed in Subsection 4.3.
+(a) Pricing
+(b) Regression
+Figure 5: Convergence for pricing with dynamic contribution and regression with dynamic contribution.
+ProFL approaches the performative optimum more closely than PFL, with faster convergence and reduced oscillation, despite a gap due to non-linear $f_{i}(\cdot)$.
+(a) Different Heterogeneity
+(b) Sample Size
+Figure 6: Impact of (a) heterogeneity and (b) sample size on convergence. Only ProFL reaches the PO point.
+In Fig.
+6 only ProFL reaches the optimal solution.
+Across all settings, PFL either stagnates at $\theta^{\text{PS}}$ or exhibits oscillations under complex dynamics.
+Fig.
+6(a) illustrates that ProFL maintains convergence to $\theta^{\text{PO}}$ across increasing heterogeneity ($\alpha\in\{0.1,\,0.3,\,1\}$), and Fig.
+6(b) shows that increasing sample size improves the stability of both algorithms, although only ProFL ultimately reaches the optimal solution.
+6 Conclusion and Future Work
+In this work, we proposed ProFL, the first approach to find the PO point in a FL setting.
+We also analyzed the impact of data contamination, addressing key challenges commonly observed in FL: limited local data sample sizes, constrained local computation, and the potentially high communication cost associated with multiple rounds of model exchange.
+We provided a convergence analysis under the PL condition, which is weaker than the assumptions required by PFL.
+However, several limitations remain.
+Our experiments rely on synthetic distribution shifts due to the lack of real-world datasets with dynamic performative feedback.
+Future work includes exploring real datasets for more realistic evaluation and relaxing the PL condition to cover broader models such as neural networks, especially large language models.
+Appendix A Additional Proofs not Detailed in the Main Paper
+A.1 Proof of Theorem 4.10
 Proof.
@@
 $\leq-\rho\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}(\overline{\theta}^{\text{PO}})\right]-\frac{1}{2}\mathbb{E}\left[\left\|g^{t}\right\|^{2}\right]+L^{2}(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right].$
-The first inequality arises from the application of Jensen’s inequality to two terms. The second inequality arises from Assumption 5.2. The last inequality arises form Lemma C.1.
+The first inequality arises from the application of Jensen’s inequality to two terms. The second inequality arises from Assumption 4.2. The last inequality arises form Lemma A.1.
 Finally, bring $T_{1}$ and $T_{2}$ into Eq. (10) and subtract $\mathcal{L}(\overline{\theta}^{\text{PO}})$ from both sides, we can get that
@@
 $\quad+\underbrace{\frac{1}{2}\left(L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)+2LG^{2}(1+\overline{\gamma})+2L(1+\overline{\gamma})\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]\right)}_{T_{3}}\eta^{2}.$ (11)
-Denote the empirical result of the performative gradient of client $i$ in global iteration $t$ and local iteration $r$ as ^$\hat{\nabla}\mathcal{L}_{i}(\theta_{i}^{t})=\hat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\hat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$.
-According to Lemma C.2, we have
+Denote the empirical result of the performative gradient of client $i$ in global iteration $t$ and local iteration $r$ as $\hat{\nabla}\mathcal{L}_{i}(\theta_{i}^{t})=\hat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\hat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$.
+According to Lemma A.2, we have
 $\mathbb{E}\left[\left\|\sum^{N}_{i=1}\alpha_{i}\nabla\mathcal{L}_{i}({\theta}^{t}_{i})-g^{t}\right\|^{2}\right]$ $\leq\sum^{N}_{i=1}\alpha_{i}\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}\right\|^{2}\right]$
@@
 $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]\leq(1-\rho\eta)^{t}\mathbb{E}[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})]+C.$
-C.2 Proof of Lemma 5.7
+A.2 Proof of Lemma 4.7
 Proof.
@@
 $=(1-\epsilon_{i})\left({f_{i}}_{z\sim D_{i}(\theta^{t}_{i})}(z)+\text{err}^{t}_{i}\right)+\epsilon_{i}\widehat{f_{i}}_{z\sim Q_{i}}(z),$
-where $\text{err}^{t}_{i}$ is the error result in sampling of client $i$ on iteration $t$ and $[\text{err}^{t}_{i}]=0$ for all $t\in[0,T-1].$ Similarly we have denote $\theta^{t}_{i}$ as $\theta_{t}$ , $\theta^{t-1}_{i}$ as $\theta_{t-1}$ , $\text{err}^{t}_{i}$ as $\text{err}_{t}$ , and $\text{err}^{t-1}_{i}$ as $\text{err}_{t-1}.$
+where $\text{err}^{t}_{i}$ is the error result in sampling of client $i$ on iteration $t$ and $\text{err}^{t}_{i}]=0$ for all $t\in[0,T-1].$ Similarly we have denote $\theta^{t}_{i}$ as $\theta_{t}$ , $\theta^{t-1}_{i}$ as $\theta_{t-1}$ , $\text{err}^{t}_{i}$ as $\text{err}_{t}$ , and $\text{err}^{t-1}_{i}$ as $\text{err}_{t-1}.$
 For each $1\leq j\leq d$ by an approximation of Taylor’s series (ignoring higher order terms), we obtain
@@
 $\leq 2\epsilon_{i}^{2}F^{2}+(1-\epsilon_{i})^{2}\left(\frac{M^{2}\eta^{4}G^{4}H^{6}}{2}+\frac{2Hd\left\|\text{err}\right\|^{2}_{i}}{\eta^{2}}\right)\lambda^{-2}_{i,\min}.$ (16)
-Where .$\left\|.\right\|_{F}$ denotes the Frobenius norm.
+Where $\left\|.\right\|_{F}$ denotes the Frobenius norm.
 (12) and (15) arise from the property that $\mathbb{E}\left[\text{err}_{t}\right]=0$ for all $t\in[0,T]$ and $\text{err}_{t}$ is independent to $a_{k}$ and $\frac{df}{d\theta}$.
@@
 $\mathbb{E}\left[\left\|\frac{\widehat{df}}{d\theta}-\frac{df}{d\theta}\right\|^{2}\right]\leq 2\epsilon_{i}^{2}F^{2}+(1-\epsilon_{i})^{2}\left(\frac{M^{2}\eta^{4}G^{4}H^{6}}{2}+\frac{2Hdf_{\varphi}(\varphi_{i};n_{i})}{\eta^{2}}\right)\lambda^{-2}_{i,\min}.$
-C.3 Proof of Lemma 5.9
+A.3 Proof of Lemma 4.9
 Proof.
@@
 (18) arises from the upper bound of $\left\|\frac{\partial\log p\left(Z_{i};f_{i}(\theta_{i}^{t})\right)}{\partial f_{i}(\theta_{i}^{t})}\right\|^{2}$ is $\mathcal{O}(d)$ and it is related to the covariance $\Sigma$ of the data for the Gaussian distribution.
-The second part will decrease to zero as $n_{i}\rightarrow 0.$ The upper bound of this part is $\mathcal{O}\left(\ell^{2}_{\max}dF^{2}f_{\varphi}(\varphi_{i};n_{i})\right)$ by using Lemma 12 in Izzo et al., (2021).
+The second part will decrease to zero as $n_{i}\rightarrow 0.$ The upper bound of this part is $\mathcal{O}\left(\ell^{2}_{\max}dF^{2}f_{\varphi}(\varphi_{i};n_{i})\right)$ by using Lemma 12 in pmlr-v139-izzo21a.
 Then we can obtain
@@
 $=\mathcal{O}\left(2\ell^{2}_{\max}d\left(\left(\boldsymbol{\omega_{F}}\right)+F^{2}\right)\right).$ (22)
-(19) arises from Eq. (3). (20) arises from Lemma 10 of Izzo et al., (2021) and Lemma 5.6. (21) arises from Jensen’s inequality of two terms. (22) arises from Lemma 5.7. ∎
-C.4 Additional Lemmas and their Proofs
-Lemma C.1.
+(19) arises from Eq. (4). (20) arises from Lemma 10 of pmlr-v139-izzo21a and Lemma 4.6. (21) arises from Jensen’s inequality of two terms. (22) arises from Lemma 4.7. ∎
+A.4 Additional Lemmas and their Proofs
+Lemma A.1.
 Let $\overline{\gamma}=\sum^{N}_{i=1}\alpha_{i}\gamma_{i}$ and $\sigma^{2}_{\gamma}$ be the variance of $\gamma_{i}$ of all clients.
 Then $\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$ is upper bounded by $(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$.
-Proof of Lemma C.1.
+Proof of Lemma A.1.
 $\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$ $=\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\mathbb{E}\left[\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]$
@@
 $\mathbb{E}\left[\sum_{i=1}^{N}\alpha_{i}(1+\gamma_{i})^{2}\left\|\overline{\theta}^{t}-{\theta}^{t}_{i}\right\|^{2}\right]\leq(R-1)^{2}G^{2}\eta^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right).$
-Lemma C.2.
+Lemma A.2.
 If there is contaminated data with distribution $Q_{i}$ and fraction $\epsilon_{i}$ , with Assumption 4.2, we can obtain that
 $\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i}(\theta_{i}^{t})-g_{i}^{t}\right\|^{2}\right]\leq(1-\epsilon_{i})(\boldsymbol{\omega_{D}})^{2}+\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})+(\boldsymbol{\omega_{D}})\right)^{2}.$
-The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemmas 5.8 and 5.9.
+The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemmas 4.8 and 4.9.
 $W_{1}\left(D,Q\right)_{\max}$ is the largest $W_{1}$ distance between $D_{i}(\theta)$ and $Q_{i}$ for all clients $i\in\mathcal{V}$ among all iteration $t\in[0,T-1]$.
-Proof of Lemma C.2.
+Proof of Lemma A.2.
 From Eq. (4) we can obtain
@@
 Because $\mathbb{E}\left[\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})]\right]=0$ and Assumption 4.2 of $L$-smoothness.
-The proof of the Lemma 12 of Izzo et al., (2021) shows that
+The proof of the Lemma 12 of pmlr-v139-izzo21a shows that
 $\mathbb{E}\left[\left\|\nabla\mathcal{L}_{i,1}(\theta_{i}^{t})-\widehat{\nabla}\mathcal{L}_{{i,1}}(D_{i}(\theta_{i}^{t}),\theta_{i}^{t})\right\|^{2}\right]=\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right).$
@@
 $(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}+(1-\epsilon_{i})^{2}\mathcal{O}(f_{\varphi}(\varphi_{i};n_{i}))(\boldsymbol{\omega_{D}})\approx(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}$ (23)
-because the estimation error of $\nabla\mathcal{L}_{i,1}$ is much smaller than $\nabla\mathcal{L}_{i,2}$ by using Lemmas 5.8 and 5.9. Apply the fact that $\epsilon_{i}\in[0,1]$ and we can obtain
+because the estimation error of $\nabla\mathcal{L}_{i,1}$ is much smaller than $\nabla\mathcal{L}_{i,2}$ by using Lemmas 4.8 and 4.9. Apply the fact that $\epsilon_{i}\in[0,1]$ and we can obtain
 $(1-\epsilon_{i})^{2}(\boldsymbol{\omega_{D}})^{2}+(1-\epsilon_{i})^{2}\mathcal{O}\left(f_{\varphi}(\varphi_{i};n_{i})\right)(\boldsymbol{\omega_{D}})+\epsilon_{i}^{2}\left(L^{2}W^{2}_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})^{2}+W_{1}\left(D,Q\right)_{\max}(\boldsymbol{\omega_{Q}})\right)$
@@
 $\leq(1-\epsilon_{i})(\boldsymbol{\omega_{D}})^{2}+\epsilon_{i}\left(LW_{1}\left(D,Q\right)_{\max}+(\boldsymbol{\omega_{Q}})+(\boldsymbol{\omega_{D}})\right)^{2}.$
-Lemma C.3 (Sensitivity of contribution Dynamic).
-For any client $i\in\mathcal{V}$, if Assumption 5.1and 5.3 hold, then with $D_{i}(\theta)=\nu_{i,1}(\theta)D_{i,1}(\theta)+\left(1-\nu_{i,1}(\theta)\right)D_{i,2}(\theta)$ for all $\theta,\theta^{\prime}\in\Theta$, we have:
+Lemma A.3 (Sensitivity of contribution Dynamic).
+For any client $i\in\mathcal{V}$, if Assumption 4.1and 4.3 hold, then with $D_{i}(\theta)=\nu_{i,1}(\theta)D_{i,1}(\theta)+\left(1-\nu_{i,1}(\theta)\right)D_{i,2}(\theta)$ for all $\theta,\theta^{\prime}\in\Theta$, we have:
 $W_{1}(D_{i}(\theta),D_{i}(\theta^{\prime}))\leq(\gamma_{i}+W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))F)\|\theta-\theta^{\prime}\|.$
-Proof of Lemma C.3.
+Proof of Lemma A.3.
 In the proof, we consider two groups with fractions $\nu_{i,1}$ and $\nu_{i,2}=1-\nu_{i,1}$. When $\nu_{i,1}(\theta)\geq\nu_{i,1}(\theta^{\prime})$, we obtain:
@@
 $\leq\left(\gamma_{i}+W_{1}(D_{i,1}(\theta),D_{i,2}(\theta))F\right)\|\theta-\theta^{\prime}\|.$
-Appendix D Additional Experimental Results
-D.1 Additional Experimental Results of Convergence to Performative Optimal Solution
-Fig.
-3 shows the results of case study house pricing regression, where an FL system is trained with 10 heterogeneous clients.
-Although PFL can stabilize the training process when data changes based on the model, it converges to an undesirable stable solution.
-In contrast, our algorithm converges to a stable solution with a much lower loss.
-In Fig.
-3(a), the results show that both algorithms exhibit faster convergence rates with larger enrollment fraction.
-However, significant oscillations occur in PFL when the enrollment fraction is low, while such oscillations are much weaker in ProFL.
-Fig.
-3(b) shows the impact of varying degrees of heterogeneity, where $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$ with $\alpha\in\{0.1,0.3,1\}$.
-The results show that, despite the varying levels of heterogeneity, our algorithm maintains the capacity to converge to ${\theta}^{\text{PO}}$.
-(a) Enrollment Fraction
-(b) Different Heterogeneity
-Figure 3: Linear Regression Results with Different Enrollment Fractions and Heterogeneity
-(a) Sample Size
-(b) Local Rounds $R$
-Figure 4: Linear Regression Results with Different Sample Size and Local Rounds
-In Fig.
-4(a), the results show that both algorithms exhibit faster convergence rates with larger sample sizes.
-With a larger sample size, the results of PFL are closer to the performative stable point and exhibit fewer oscillations.
-However, PFL cannot converge to the performative optimal point, no matter how many samples are used.
-Fig.
-4(b) demonstrates that a larger value of $R$ corresponds to faster convergence, reducing the communication time required in the FL system to reach convergence.
-Fig.
-5(a) shows the results of the case study pricing with dynamic contribution with 10 heterogeneous clients.
-The results indicate that, unlike PFL, which converges to $\theta^{\text{PS}}$, our algorithm converges to a solution, which is close to the performative optimal point.
-Fig.
-5(b) shows the results of the case study regression with dynamic contribution with 10 heterogeneous clients.
-ProFL has faster convergence speed and less oscillation, although there is a gap between the results of both algorithms and the performative optimal point.
-The gap is a result of the non-linear $f_{i}(.)$, which has been discussed in more detail in the main paper.
-(a) Pricing
-(b) Regression
-Figure 5: Contribution Dynamics
-D.2 Performance when ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
-In this section, we introduce a special case where $\theta^{\text{PO}}=\theta^{\text{PS}}$.
-This special case can arise in both strategic and dynamic contribution settings.
-We present the results in Fig.
-6, demonstrating that when $\theta^{\text{PO}}=\theta^{\text{PS}}$, both algorithms converge to the same point.
-(a) Strategic Dynamic
-(b) Contribution Dynamic
-Figure 6: Dynamics with ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
-D.2.1 Performance without Model-dependent Shifts.
-We also evaluate the algorithm when the data distribution does not change based on the model.
-In this case study, we use the Adult dataset for a binary classification task.
-Specifically, there are 10 homogeneous clients with $\gamma_{i,0}=\gamma_{i,1}=0$.
-We consider four features from the Adult dataset.
-When the data remains constant as the model is updated, both PFL and our algorithm reduce to FedAvg, and they converge to the same solution.
-This is illustrated in Fig.
-6(a).
-D.2.2 Performance when $\theta^{\text{PO}}=\theta^{\text{PS}}$ with Dynamic Contribution
-Consider case study pricing with dynamic contribution with $K=2$ groups with different fixed distributions but a different one-dimensional fraction where $f_{i,1}(\theta)=0.5+0.5\theta$ and $f_{i,2}(\theta)=1-f_{i,1}(\theta)$.
-Under this special case, the performative stable and optimal points are identical, i.e., $\theta^{\text{PO}}=\theta^{\text{PS}}$.
-We can calculate $\theta^{\text{PO}}=\frac{a_{1}+a_{2}}{2(a_{1}-a_{2})}$ and $\theta^{\text{PS}}=\frac{a_{1}+a_{2}}{a_{1}-a_{2}}$.
-When $a_{1}+a_{2}=0$, $\theta^{\text{PO}}=\theta^{\text{PS}}$.
-In this case, we set $a_{1}=-a_{2}=0.5$.
-Fig.
-6(b) illustrates the results for this specific case study, showing that both our algorithm and PFL converge to the same solution.
-However, our method exhibits faster convergence compared to PFL.
-D.2.3 Figures of Table 2
-The figures illustrating the experimental results from Table 2 in the main paper are shown in Fig. 7.
-(a) Same Distributions
-(b) Different Distributions
-(c) Credit Dataset
-(d) Adult Dataset
-Figure 7: Binary Classification on Synthetic and Realistic Dataset
-Appendix E Training details
-The experimental setup and parameter selections follow the methodology established in Izzo et al., (2021); Jin et al., (2024), ensuring consistency and comparability with existing results.
-E.1 Realistic Data
-  • Adult data (Becker and Kohavi,, 1996) where the goal is to predict whether a person’s annual income exceeds \$50K based on 14 features (e.g., age, years of education).
-  • Give Me Some Credit data (Credit Fusion,, 2011) which has 11 features (e.g., monthly income, debt ratio) and can be utilized to predict whether a person has experienced a 90-day past-due delinquency or worse.
-E.2 Loss Functions and Ridge Penalty
+Appendix B Case Studies
+B.1 Pricing with Dynamic Demands
+Consider a company that aims to find the proper prices for various goods.
+Let $\theta\in\mathbb{R}^{d}$ be the prices associated with the set of $d$ goods, and $Z_{i}\in\mathbb{R}^{d}$ be the retailer $i$’s demands for these goods.
+We assume $Z_{i}\sim\mathcal{N}(\mu(\theta),\sigma^{2})$ is model dependent Gaussian distribution, i.e., mean demand for each good decreases linearly as the price increases.
+$\mu(\theta)=\mu_{0}-\gamma_{i}\theta$, here $\gamma_{i}\geq 0$ measures the price sensitivity and varies across different retailers.
+The goal is to maximizes total revenue over all retailers $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\theta^{T}Z_{i}]$.
+B.2 Pricing with Dynamic Contribution
+Consider a similar setting as above where a company aims to find prices for various goods that maximize the total revenue.
+Suppose the retailers it faces have consumers from $K$ groups (e.g., old/young) and these consumers have fixed demands but they have options to purchase the goods from other companies.
+Here we assume the demands of retailer $i$’s consumers from group $k\in[K]$ have fixed distribution $D_{i,k}=\mathcal{N}(\mu_{i,k},\sigma_{i}^{2})$, but the fraction of groups change dynamically based on prices $\theta$.
+Let $r_{i,k}(\theta)=B-\mathbb{E}_{Z_{i}\sim D_{i,k}}[\theta^{T}Z_{i}]$ be the expected remaining budget of group $k$ after purchasing the goods (with initial budget $B$) and assume the fraction of group $k$ is $\nu_{i,k}(\theta)=\frac{r_{i,k}(\theta)}{\sum_{k=1}^{K}r_{i,k}(\theta)}$ (i.e., groups with more remaining budgets have more incentives to stay in the system).
+Then the retailer $i$’s total demands would follow the mixture of distributions $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}$.
+B.3 Binary Strategic Classification
+[…]
+B.4 House Pricing Regression
+Consider a regression task that aims to find the proper listing prices for houses based on features such as size, age, number of bedrooms.
+Suppose the features of houses in district $i$ follow $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$.
+Given predictive model $\theta$, let the listing price be $h(X_{i})=\theta^{T}X_{i}$.
+Since the houses with higher listing prices tend to lower the demand, the actual selling price $Y$ depends on $\theta$ and we assume $Y_{i}=(a_{i}-\gamma_{i}\theta)^{T}X_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{n})$ is Gaussian noise.
+The goal is to minimizes prediction error
+B.5 Regression with dynamic contribution
+Consider an example of retail inventory management, where the regression model $\theta$ predicts product sales considering features like seasonal trends and advertising.
+Distributor $i$ has retailers from two groups $k\in\{1,2\}$, and each retailer’s data distribution is fixed.
+The fraction of retailers from each group is $\nu_{i,k}(\theta)=\frac{\ell_{i,-k}(\theta)+c}{\sum_{k^{\prime}\in\{1,2\}}\left(\ell_{i,k^{\prime}}(\theta)+c\right)}$ where $k=\{1,2\}\setminus k$ because the retailers are more likely to stop purchasing if they experience higher prediction error $\ell_{i,k}(\theta)$.
+The goal is to find $\theta$ that minimizes expected prediction error $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{(X_{i},Y_{i})\sim D_{i}(\theta)}\left[(\theta^{T}X_{i}-Y_{i})^{2}\right]$, where $Y_{i}$ is a weighted combination of fixed distributions.
+Appendix C Training details
+The experimental setup and parameter selections follow the methodology established in pmlr-v139-izzo21a; jin2023performative, ensuring consistency and comparability with existing results.
+C.1 Realistic Data
+  • Adult data misc_adult_2 where the goal is to predict whether a person’s annual income exceeds \$50K based on 14 features (e.g., age, years of education).
+  • Give Me Some Credit data GiveMeSomeCredit which has 11 features (e.g., monthly income, debt ratio) and can be utilized to predict whether a person has experienced a 90-day past-due delinquency or worse.
+C.2 Loss Functions and Ridge Penalty
 We use ridge-regularized cross-entropy loss for all binary classification cases, and the ridge penalty is $0.01$.
 We use ridge-regularized squared loss for all regression cases, and the ridge penalty is $\frac{10}{3}$.
-E.3 Random Seeds
-In our experiments, which show the results of 10 trials, the random seeds used range from 0 to 9.
-E.4 Parameters and Distributions
-The experimental parameters are listed in Table 3, and the distribution $D_{i}(\cdot)$ for each experiment is detailed in Table 4.
-The corresponding case studies for the experiments are shown in Table 5.
+C.3 Parameters and Distributions
+The experimental parameters are listed in Table 2.
 Table 2: Parameters in the Experiments
 Figure/Table $\eta$ $R$ H $n_{i}$ $\epsilon_{i}$
-Table 1: PG 0.01 -- 1 500 0
-Table 1: ProFL 0.01 5 1 500 0
-Table 2: Same $D_{i}$ 0.03 10 500 500 0
-Table 2: Different $D_{i}$ 0.03 500 1 500 0
-Table 2: Credit 0.01 5 1 5812 0
-Table 2: Adult 0.03 5 1 2368 0
+Table I: PG 0.01 - 1 500 0
+Table I: ProFL 0.01 5 1 500 0
 Figure 2a 0.001 5 25 500 0, 0.1,0.2,0.4
@@
 Figure 3c 0.001 5 4,20,100 500 0
+Figure 4a 0.03 10 500 500 0
+Figure 4b 0.03 500 1 500 0
+Figure 4c 0.01 5 1 5812 0
+Figure 4d 0.03 5 1 2368 0
+Figure 5a,5b PFL:0.005 ProFL:0.001 5 1 500 0
 Figure 6a PFL:0.005 ProFL:0.001 5 0.2, 0.5, 1 500 0
-Figure 3b PFL:0.005 ProFL:0.001 5 1 500 0
-Figure 3a 0.03 5 1 2368 0
-Figure 3b 0.03 5 1 2368 0
-Figure 4a PFL:0.005 ProFL:0.001 5 1 5, 50, 500 0
-Figure 4b PFL:0.005 ProFL:0.001 5,20,50 1 500 0
-Figure 5a 0.03 5 1 500 0
-Figure 5b 0.001 5 1 500 0
-Figure 6a 0.03 5 1 500 0
-Figure 6b 0.03 5 1 500 0
-Table 4: Data Distributions in the Experiments
-Figure/Table Distribution
-Table 1: PG and ProFL $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0,0.25,0.5\}$
-Table 2: Same $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$ and $X_{i}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
-$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$.
-Table 2: Different $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$, $X_{1}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$
-and $X_{2}\sim\mathcal{N}(-0.8-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
-$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$
-Table 2: Credit $\gamma_{i,1}=3$ and $\gamma_{i,0}=0.$ Give Me Some Credit datasetCredit Fusion, (2011)
-Table 2: Adult $\gamma_{i,1}\in[2.8,3.2]$ and $\gamma_{i,0}\in[0,0.004].$ Adult dataset (Becker and Kohavi,, 1996)
-Figure 1a $\mu_{0}\in[6,7]$, $\sigma=1$, $\gamma_{i}\in[1,3]$
-Figure 1b,1c,2a-c $\gamma_{i}\in[0.09,0.11]$
-Figure 3a $\gamma_{i}\in[1.5,1.8]$
-Figure 3b $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0.1,0.3,1\}$
-Figure 4a $\gamma_{i}\in[1.5,1.8]$
-Figure 4b $\gamma_{i}\in[1.5,1.8]$
-Figure 5a $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$ $f(\theta)=0.5+0.5\theta$
-Figure 5b $y=f_{i}(\theta)x+3(1-f_{i}(\theta))x+\varepsilon,\varepsilon\sim\mathcal{N}(0,4)$
-Figure 6a $\gamma_{i}=0.$ 10 homogeneous clients with Adult dataset
-Figure 6b $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$
-Table 5: Type of Case Study and Corresponding Figures or Tables
-Figure/Table Type of Case Study
-Table 1: house pricing regression
-Table 2: binary strategic classification
-Figure 1a pricing with dynamic demands
-Figure 1b,1c,2a-c pricing with dynamic contribution
-Figure 3a,3b,4a,4b house pricing regression
-Figure 5a pricing with dynamic contribution
-Figure 5b regression with dynamic contribution
-Figure 6a binary strategic classification
-Figure 6b pricing with dynamic contribution
+Figure 6b PFL:0.005 ProFL:0.001 5 1 5, 50, 500 0
+C.4 Experiments Compute Resources
+CPU: 11th Gen Intel(R) Core(TM) i7-11800H @ 2.30GHz 2.30 GHz.